### v1

# Implicit Bias of SignGD and Adam on Multiclass Separable Data

Abstract
In the optimization of overparameterized models, different gradient-based methods can achieve zero training error yet converge to distinctly different solutions inducing different generalization properties. While a decade of research on implicit optimization bias has illuminated this phenomenon in various settings, even the foundational case of linear classification with separable data still has important open questions. We resolve a fundamental gap by characterizing the implicit bias of both Adam and Sign Gradient Descent in multi-class cross-entropy minimization: we prove that their iterates converge to solutions that maximize the margin with respect to the classifier matrix’s max-norm and characterize the rate of convergence. We extend our results to general p-norm normalized steepest descent algorithms and to other multi-class losses.
Machine Learning, ICML
1 Introduction
Machine learning models are trained to minimize a surrogate loss function with the goal of finding model-weight configurations that generalize well. The loss function used for training is typically a proxy for the evaluation metric. The prototypical example is one-hot classification where the predominant choice, the cross-entropy (CE) loss, serves as a convex surrogate of the zero-one metric that measures correct class membership. Intuitively, a good surrogate loss is one where weight configurations minimizing the training loss also achieve strong generalization performance. However, modern machine learning models are overparameterized, leading to multiple weight configurations that achieve identical training loss but exhibit markedly different generalization properties (Zhang et al., 2017; Belkin et al., 2019).
This observation has motivated extensive research into the implicit bias (or implicit regularization) of gradient-based optimization. This theory investigates how optimizers select specific solutions from the infinite set of possible minimizers in overparameterized settings. The key insight is that gradient-based methods inherently prefer “simple” solutions according to optimizer-specific notions of simplicity. Understanding this preference requires analyzing not just loss convergence, but the geometric trajectory of parameter updates throughout training.
The prototypical example—which has rightfully earned its place as a “textbook” setting in the field—is the study of implicit optimization bias of gradient descent in linear classification. In this setting, embeddings of the training data are assumed fixed and only the classifier is learned, with overparameterization modeled through linear separability of the training data. (Soudry et al., 2018) demonstrated that while the GD parameters diverge in norm, they converge in direction to the maximum-margin classifier—the minimum $L_{2}$-norm solution that separates the data. This non-trivial result, despite (or perhaps because of) the simplicity of its setting, spurred numerous research directions extending to nonlinear models, alternative optimizers, different loss functions, and theoretical explanations of phenomena like benign overfitting (see Sec. 2).
Our paper contributes a fundamental result to this literature by establishing the implicit bias of Adam and its simpler variant, signed GD (SignGD), in the multiclass version of this textbook setting: minimization of CE loss on multiclass linearly separable data. Our work addresses several gaps between theory assumption and practice: the predominant focus on binary classification despite multiclass problems dominating real-world applications; the much theoretical emphasis on exponential loss despite practitioners’ overwhelming preference for cross-entropy loss; and the limited theoretical understanding of Adam’s implicit bias, which has only recently been studied in the binary case (Zhang et al., 2024) despite its widespread adoption in deep learning practice. Here, we provide a direct analysis of the multiclass setting by exploiting the nice properties of the softmax function, avoiding the traditional approach of reducing multiclass problems to binary ones (Wang & Scott, 2024).
Contributions. Our contributions are as follows: For multiclass separable data trained with CE loss, we show that the iterates of SignGD converge to a solution that maximizes the margin defined with respect to (w.r.t.) the matrix max-norm, with a rate $\mathcal{O}(\frac{1}{t^{1/2}})$. We generalize this result to normalized steepest descent algorithms w.r.t. any entry-wise matrix $p$-norm. This directly extends Nacson et al. (2019); Sun et al. (2023)’s results to multiclass classification and to the widely-used CE loss. To achieve this, we construct a proxy for the loss and show that it closely traces its value and gradient. We also show the same machinery applies to other multiclass losses such as the exponential loss (Mukherjee & Schapire, 2010) and the PairLogLoss (Wang et al., 2021b).
Under the same setting, we prove that the iterates of Adam also maximize the margin w.r.t. the matrix max-norm, with a rate $\mathcal{O}(\frac{1}{t^{1/3}})$. This matches the convergence rate previously established for binary classification (Zhang et al., 2024), and remarkably, is independent of the number of classes $k$. This class-independence is non-trivial since naive reduction approaches that map a $k$-class problem in $d$ dimensions to a binary problem in $kd$ dimensions inevitably introduce $k$-dependent factors. Our key insight is to decompose the CE proxy function class-wise, enabling us to bound Adam’s first and second gradient moments separately for each class.This decomposition reveals how each component of the proxy function governs the dynamics of the weight vectors associated with its corresponding class. Finally, we experimentally verify our theory predictions for the algorithms considered. Specific to Adam, we numerically demonstrate that the implicit bias results are consistent with or without the (small) stability constant in the multiclass setting, and the solutions found by SignGD and Adam favor the max-norm margin over the 2-norm margin.
2 Related Works
Starting with GD, the foundational result by (Soudry et al., 2018) showed that gradient descent optimization of logistic loss on linearly separable data converges in direction to the $L_{2}$ max-margin classifier at a rate $O(1/\log(t))$. Contemporaneous work by (Ji & Telgarsky, 2019) generalized this by removing the data separability requirement. (Ji et al., 2020) later connected these findings to earlier work on regularization paths of logistic loss minimization (Rosset et al., 2003), which enabled extensions to other loss functions (e.g., those with polynomial tail decay). More recently, (Wu et al., 2024b) extends these results to the large step size regime with the same $O(1/\log(t))$ rate. The relatively slow convergence rate to the max-margin classifier motivated investigation into adaptive step-sizes. (Nacson et al., 2019) showed that normalized gradient descent (NGD) with decaying step-size $\eta_{t}=1/\sqrt{t}$ achieves $L_{2}$-margin convergence at rate $O(1/\sqrt{t})$. This rate was improved to $O(1/t)$ by (Ji & Telgarsky, 2021) using constant step-sizes, and further to $O(1/t^{2})$ through a specific momentum formulation (Ji et al., 2021). Besides linear classifications, implicit bias of GD has been studied for least squares (Gunasekar et al., 2017, 2018), homogeneous (Lyu & Li, 2019; Ji & Telgarsky, 2020) and non-homogeneous neural networks (Wu et al., 2024a), and matrix factorization (Gunasekar et al., 2017); see (Vardi, 2023) for a survey.
[…]
On the other hand, the implicit bias of adaptive algorithms such as Adagrad (Duchi et al., 2011) or Adam (Kingma & Ba, 2014) is less explored compared to GD. (Qian & Qian, 2019) studied the implict bias of Adagrad and showed its directional convergence to a solution characterized by a quadratic minimization problem. (Wang et al., 2021a, 2022) demonstrated the normalized iterates of Adam (with non-negligible stability constant) converge to a KKT point of a $L_{2}$-margin maximization problem for homogeneous neural networks. More recently and most relevant to our work, (Zhang et al., 2024) studied the implicit bias of Adam without the stability constant on binary linearly separable data. They showed that unlike GD, the Adam iterates converge to a solution that maximizes the margin with respect to the $L_{\infty}$-norm. This study excluding the stability constant is practically-relevant given the magnitude of the constant is typically very small (default $81e-$ 8 in PyTorch (Paszke et al., 2019)). This setting is also the focus of another closely-related recent study of the implicit bias of AdamW (Xie & Li, 2024), where the authors again establish that convergence aligns with the $L_{\infty}$ (rather than $L_{2}$) geometry. Our work extends these latter studies to the multiclass setting (see Remark 6.5 for technical comparisons).
All the above mentioned works focus solely on binary classification. The noticeable gap in analysis of multi class classification in most existing literature is recently emphasized by (Ravi et al., 2024) who extend the implicit bias result of (Soudry et al., 2018) to multiclass classification for losses with exponential tails, including cross-entropy, multiclass exponential, and PairLogLoss. Their approach leverages a framework introduced by (Wang & Scott, 2024), mapping multiclass analysis to binary cases. Our work directly addresses their open questions regarding the implicit bias of alternative gradient-based methods in multiclass settings by analyzing methods with adaptive step-sizes. Thanks to the adaptive step-sizes, our rates of convergence to the margin improve to polynomial dependence on $t$. Furthermore, our technical approach differs: rather than mapping to binary analysis, we work directly with multiclass losses, exploiting properties of the softmax function to produce elegant proofs that apply to all three losses studied by (Ravi et al., 2024). Our class-wise decomposition is crucial for analyzing Adam with the same convergence rate as the binary case, avoiding any extra factors that depend on the number of classes.
3 Preliminaries
Notations
For any integer $k$, $[k]$ denotes $\{1,\ldots,k\}$. Matrices, vectors, and scalars are denoted by $\bm{A}$, $\bm{a}$, and $a$ respectively. For matrix $\bm{A}$, we denote its $(i,j)$-th entry as $\bm{A}[i,j]$, and for vector $\bm{a}$, its $i$-th entry as $\bm{a}[i]$. We consider entry-wise matrix $p$-norms defined as $\|\bm{A}\|_{p}=(\sum_{i,j}|{\bm{A}}[i,j]|^{p})^{1/p}$. Central to our results are: the infinity norm, denoted as ${\left\|\bm{A}\right\|_{\max}}=\max_{i,j}|\bm{A}[i,j]|$ and called the max-norm, and the entry-wise 1-norm, denoted as ${\left\|{\bm{A}}\right\|_{\rm{sum}}}=\sum_{i,j}|\bm{A}[i,j]|$. For any other entry-wise $p$-norm with $p>1$, we write $\|{\bm{A}}\|$ (dropping subscripts) for simplicity. We denote by $\|\bm{A}\|_{*}$ the dual-norm with respect to the standard matrix inner product $\langle\bm{A},\bm{B}\rangle=\operatorname{tr}(\bm{A}^{\top}\bm{B})$. The entry-wise 1-norm is dual to the max-norm. For vectors, the max-norm is equivalent to the infinity norm, denoted as $\|\bm{a}\|_{\infty}$, while we denote the $\ell_{1}$ norm as $\|\bm{a}\|_{1}$. Let indicator $\delta_{ij}$ such that $\delta_{ij}=1$ if and only if $i=j$. Denote $\mathbb{S}:\mathbb{R}^{k}\rightarrow\triangle^{k-1}$ the softmax map of $k$-dimensional vectors to the probability simplex $\triangle^{k-1}$ such that for $\bm{a}\in\mathbb{R}^{k}$:
$\mathbb{S}(\bm{a})=\big{[}\frac{\exp({\bm{a}[c]})}{\sum_{c\in[k]}\exp(\bm{a}[c])}\big{]}_{c=1}^{k}\in\triangle^{k-1}.$
Let $\mathbb{S}_{c}({\bm{v}})$ denote the $c$-th entry of $\mathbb{S}({\bm{v}})$. Let $\mathbb{S}^{\prime}(\bm{a})=\operatorname{diag}(\mathbb{S}(\bm{a}))-\mathbb{S}(\bm{a})\mathbb{S}(\bm{a})^{\top}$ denote the softmax gradient, with $\operatorname{diag}(\cdot)$ a diagonal matrix. Finally, let $\{\bm{e}_{c}\}_{c=1}^{k}$ be the standard basis vectors of $\mathbb{R}^{k}$.
Setup
Consider a multiclass classification problem with training data ${\bm{h}_{1},\ldots,\bm{h}_{n}}$ and labels ${y_{1},\ldots,y_{n}}$. Each datapoint $\bm{h}_{i}\in\mathbb{R}^{d}$ is a vector in a $d$-dimensional embedding space (denote data matrix ${\bm{H}}=[\bm{h}_{1},\ldots,\bm{h}_{n}]^{\top}\in\mathbb{R}^{n\times d}$), and each label $y_{i}\in[k]$ represents one of $k$ classes. We assume each class contains at least one datapoint. The classifier $f_{{\bm{W}}}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is a linear model with weight matrix ${\bm{W}}\in\mathbb{R}^{k\times d}$. The model outputs logits $\bm{\ell}_{i}=f_{{\bm{W}}}(\bm{h}_{i})={\bm{W}}\bm{h}_{i}$ for $i\in[n]$, which are passed through the softmax map to produce class probabilities $\hat{p}(c|\bm{h}_{i})=\mathbb{S}_{c}(\bm{\ell}_{i})$. We train using empirical risk minimization (ERM): $\mathcal{L}_{\text{ERM}}({\bm{W}}):=-\frac{1}{n}\sum_{i\in[n]}\ell\left({{\bm{W}}\bm{h}_{i}};y_{i}\right)\,,$ where the loss function $\ell$ takes as input the logits of a datapoint and its label. The predominant choice in classification is the CE loss
$\mathcal{L}({\bm{W}}):=-\frac{1}{n}\sum\nolimits_{i\in[n]}\log\big{(}\mathbb{S}_{{y_{i}}}({\bm{W}}\bm{h}_{i})\big{)}$ (1)
$=\frac{1}{n}\sum\nolimits_{i\in[n]}\log\big{(}1+\sum\nolimits_{c\neq y_{i}}\exp(-(\bm{e}_{y_{i}}-\bm{e}_{c})^{\top}{\bm{W}}\bm{h}_{i})\big{)}\,,$
We focus our discussions on the CE loss due to its ubiquity in practice. However, our results hold for other multiclass losses such as the exponential (Mukherjee & Schapire, 2010) and the PairLogLoss (Wang et al., 2021b) (see App. F).
[…]
$\gamma_{\|\cdot\|}:=\max_{\|{\bm{W}}\|\leq 1}\,\min_{\begin{subarray}{c}i\in[n]\\ c\neq{y_{i}}\end{subarray}}\,\left(\bm{e}_{y_{i}}-\bm{e}_{c}\right)^{\top}{\bm{W}}\bm{h}_{i}\,,$ (2)
[…]
Optimization Methods
To minimize the empirical loss $\mathcal{L}_{\text{ERM}}$, we use Adam without the stability constant, which performs the following coordinate-wise updates for iteration $t\geq 0$ and initialization ${\bm{W}}_{0}$ (Kingma & Ba, 2014):
$\mathbf{M}_{t}$ $=\beta_{1}\mathbf{M}_{t-1}+(1-\beta_{1})\nabla\mathcal{L}({\bm{W}}_{t})$ (3a)
$\mathbf{V}_{t}$ $=\beta_{2}\mathbf{V}_{t-1}+(1-\beta_{2})\nabla\mathcal{L}({\bm{W}}_{t})^{2}$ (3b)
${\bm{W}}_{t+1}$ $={\bm{W}}_{t}-\eta_{t}\frac{\mathbf{M}_{t}}{\sqrt{\mathbf{V}_{t}}},$ (3c)
where $\mathbf{M}_{t}$, $\mathbf{V}_{t}$ are first and second moment estimates of the gradient, with decay rates $\beta_{1}$ and $\beta_{2}$. The squaring $(\cdot)^{2}$ and dividing $\frac{\cdot}{\cdot}$ operations are applied entry-wise.
In the special case $\beta_{1}=\beta_{2}=0$, the updates simplify to:
${\bm{W}}_{t+1}$ $={\bm{W}}_{t}-\eta_{t}\texttt{sign}(\nabla\mathcal{L}({\bm{W}}_{t}))),$ (4)
[…]
${\bm{W}}_{t+1}={\bm{W}}_{t}-\eta_{t}\bm{\Delta}_{t},\qquad\text{where}$
$\bm{\Delta}_{t}:=\arg\max\nolimits_{\|\bm{\Delta}\|\leq 1}\langle\nabla\mathcal{L}({\bm{W}}_{t}),\bm{\Delta}\rangle\,.$ (5)
Note that this reduces to SignGD, Coordinate Descent (e.g., (Nutini et al., 2015)), or normalized gradient-descent (NGD) when the max-norm (i.e. $p=\infty$), the entry-wise $1$-norm, or the Frobenious Euclidean-norm (i.e. $p=2$) is used, respectively.
Assumptions
Establishing the implicit bias of the above mentioned gradient-based optimization algorithms, requires the following assumptions. First, we assume data are linearly separable, ensuring the margin $\gamma_{\|\cdot\|}$ is strictly positive, an assumption routinely used in previous works (Soudry et al., 2018; Ravi et al., 2024; Soudry et al., 2018; Gunasekar et al., 2018; Nacson et al., 2019; Wu et al., 2024b).
Assumption 3.1.
There exists ${\bm{W}}\in\mathbb{R}^{k\times d}$ such that $\min_{c\neq y_{i}}(\bm{e}_{y_{i}}-\bm{e}_{c})^{T}{\bm{W}}\bm{h}_{i}>0$ for all $i\in[n]$.
The second assumption ensures that all entries of the second moment buffer ${\bm{V}}_{t}$ of Adam are bounded away from 00 for all $t\geq 0$. Previously used by Zhang et al. (2024) in binary classification, this assumption is satisfied when the data distribution is continuous and non-degenerate. A similar assumption appears in (Xie & Li, 2024).
Assumption 3.2.
The Adam initialization satisfies $\nabla\mathcal{L}({\bm{W}}_{0})[c,j]^{2}\geq\omega$ for all $c\in[k]$ and $j\in[d]$.
In this work, we consider a decay learning rate schedule of the form $\eta_{t}=\Theta(\frac{1}{t^{a}})$, where $a\in(0,1]$. Such schedule has been studied in the convergence and implicit bias of various optimization algorithms (e.g., (Bottou et al., 2018; Nacson et al., 2019; Sun et al., 2023)) including Adam (Huang et al., 2021; Zhang et al., 2024; Xie & Li, 2024).
Assumption 3.3.
The learning rate schedule $\{\eta_{t}\}$ is decreasing with respect to $t$ and satisfies the following conditions: $\lim_{t\rightarrow\infty}\eta_{t}=0$ and $\sum_{t=0}^{\infty}\eta_{t}=\infty$.
Assumption 3.4 can be satisfied by the above learning rate for a sufficiently large $t$ shown in Zhang et al. (2024, Lemma C.1). It is used in our analysis of Adam.
Assumption 3.4.
The learning rate schedule satisfies the following: let $\beta\in(0,1)$ and $c_{1}>0$ be two constants, there exist time $t_{0}\in\mathbb{N}_{+}$ and constant $c_{2}=c_{2}(c_{1},\beta)>0$ such that $\sum_{s=0}^{t}\beta^{s}(e^{c_{1}\sum_{\tau=1}^{s}\eta_{s-\tau}}-1)\leq c_{2}\eta_{t}$ for all $t\geq t_{0}$.
Finally, we assume the $1$-norm of the data is bounded. Similar assumptions were used in (Ji & Telgarsky, 2019), (Nacson et al., 2019), (Wu et al., 2024b), and (Zhang et al., 2024).
Assumption 3.5.
There exists constant $B>0$ such that $\lVert\bm{h}_{i}\rVert_{1}\leq B$ for all $i\in[n]$.
4 ${\mathcal{G}}({\bm{W}})$ A proxy to $\mathcal{L}({\bm{W}})$
Analyzing margin convergence begins with studying loss convergence through second-order Taylor expansion of the CE loss:
$\mathcal{L}({\bm{W}}+\bm{\Delta})=\mathcal{L}({\bm{W}})+\langle\nabla\mathcal{L}({\bm{W}}),\bm{\Delta}\rangle$
$\quad+\frac{1}{2n}\sum\nolimits_{i\in[n]}\bm{h}_{i}^{\top}\bm{\Delta}^{\top}{\mathbb{S}}^{\prime}({\bm{W}}\bm{h}_{i})\bm{\Delta}\bm{h}_{i}+o(\|\bm{\Delta}\|_{F}^{3}),$ (6)
where recall that ${\mathbb{S}}^{\prime}({\bm{v}})=\operatorname{diag}({\bm{v}})-{\bm{v}}{\bm{v}}^{\top}$. To bound the loss at ${\bm{W}}_{t+1}={\bm{W}}_{t}-\eta_{t}\bm{\Delta}_{t}$, we must bound both terms in (6). For the NSD updates defined in Eq. (5), the first term evaluates to $-\eta_{t}\|\nabla\mathcal{L}({\bm{W}})\|_{*}$ ($\|\cdot\|_{*}$ is the dual norm). This leads to two key tasks: (1) Lower-bounding the dual gradient norm. (2) Upper-bounding the second-order term.
[…]
Construction of ${\mathcal{G}}({\bm{W}})$
Before showing our construction for the CE loss, it is insightful to discuss how previous works do this in the binary case with labels $y_{b,i}\in{\pm 1}$, classifier vector ${\bm{w}}\in\mathbb{R}^{d}$ and binary margin $\gamma_{b}\coloneqq\max_{\lVert{\bm{w}}\rVert\leq 1}\min_{i\in[n]}y_{b,i}{\bm{w}}^{\top}\bm{h}_{i}$. For exponential loss, Gunasekar et al. (2018) showed that $\lVert\nabla\mathcal{L}({\bm{w}})\rVert\geq\gamma_{b}\mathcal{L}({\bm{w}})$. For logistic loss $\ell(t)=\log(1+\exp(-t))$, Zhang et al. (2024) proved $\lVert\nabla\mathcal{L}({\bm{w}})\rVert_{1}\geq\gamma_{b}{\mathcal{G}}({\bm{w}})$, where ${\mathcal{G}}({\bm{w}})=\frac{1}{n}\sum_{i=1}^{n}|\ell^{\prime}(y_{b,i}{\bm{w}}^{\top}\bm{h}_{i})|$ and $\ell^{\prime}$ is the first-derivative. In both cases, one can take the common form ${\mathcal{G}}_{b}({\bm{w}})=\frac{1}{n}\sum_{i=1}^{n}|\ell^{\prime}(y_{b,i}{\bm{w}}^{\top}\bm{h}_{i})|$. The proof relies on showing $\gamma\leq\min_{\bm{r}\in\triangle^{n-1}}\lVert{\bm{H}}^{T}\bm{r}\rVert$ via Fenchel Duality (Telgarsky, 2013; Gunasekar et al., 2018) and appropriately choosing $\bm{r}$.
In the multiclass setting, where the loss function is vector-valued, it is unclear how to extend the binary proof or definition of ${\mathcal{G}}({\bm{W}})$. To this end, we realize that the key is in the proper manipulation of the gradient inner product $\langle{\bm{A}},-\nabla\mathcal{L}({\bm{W}})\rangle$ (for arbitrary matrix ${\bm{A}}\in\mathbb{R}^{k\times d}$). The CE gradient evaluates to $\nabla\mathcal{L}({\bm{W}})=\frac{1}{n}\sum_{i=1}^{n}(\bm{e}_{y_{i}}-\mathbb{S}({\bm{W}}\bm{h}_{i}))\bm{h}_{i}^{\top}$ and using the fact that $\mathbb{S}({\bm{W}}\bm{h}_{i})\in\triangle^{k-1}$, it turns out that we can express (details in Lemma A.1)
$\langle{\bm{A}},-\nabla\mathcal{L}({\bm{W}})\rangle=\frac{1}{n}\sum_{i\in[n]}\sum_{c\neq y_{i}}\mathbb{S}_{c}({\bm{W}}\bm{h}_{i})(\bm{e}_{y_{i}}-\bm{e}_{c})^{\top}{\bm{A}}\bm{h}_{i}\,.$
This motivates defining ${\mathcal{G}}({\bm{W}})$ as:
${\mathcal{G}}({\bm{W}})\coloneqq\frac{1}{n}\sum_{i\in[n]}(1-\mathbb{S}_{{y_{i}}}({\bm{W}}\bm{h}_{i}))\,.$ (7)
The following key lemma , which directly follows from the inner-product calculation above and our definition of ${\mathcal{G}}({\bm{W}})$, confirms that this is the right choice. For convenience, denote $s_{ic}\coloneqq\mathbb{S}_{c}({\bm{W}}\bm{h}_{i})$, for $i\in[n],c\in[k]$.
Lemma 4.1 (Lower bounding the gradient dual-norm).
For any ${\bm{W}}\in\mathbb{R}^{k\times d}$, it holds that ${\left\|\nabla\mathcal{L}({\bm{W}})\right\|_{\rm{sum}}}\geq\gamma\cdot{\mathcal{G}}({\bm{W}})$.
Proof.
By duality, the above calculated formula for the CE-gradient inner product and the fact that $\sum_{c\in[k]}s_{ic}=1$:
${\left\|\nabla\mathcal{L}({\bm{W}})\right\|_{\rm{sum}}}=\max\nolimits_{{\left\|\bm{A}\right\|_{\max}}\leq 1}\,\langle\bm{A},-\nabla\mathcal{L}({\bm{W}})\rangle$
$=\max\nolimits_{{\left\|\bm{A}\right\|_{\max}}\leq 1}\,\frac{1}{n}\sum\nolimits_{i\in[n]}\sum\nolimits_{c\neq y_{i}}s_{ic}\,(\bm{e}_{{y_{i}}}-\bm{e}_{c})^{\top}\bm{A}\bm{h}_{i}$
$\geq\max_{{\left\|\bm{A}\right\|_{\max}}\leq 1}\frac{1}{n}\sum_{i\in[n]}(1-s_{iy_{i}})\,\min_{c\neq y_{i}}(\bm{e}_{y_{i}}-\bm{e}_{c})^{T}\bm{A}\bm{h}_{i}$
$=\frac{1}{n}\sum_{i\in[n]}(1-s_{iy_{i}})\cdot\max_{{\left\|\bm{A}\right\|_{\max}}\leq 1}\min_{i\in[n],c\neq y_{i}}(\bm{e}_{y_{i}}-\bm{e}_{c})^{T}\bm{A}\bm{h}_{i}\,.$
To finish recall the definitions of ${\mathcal{G}}({\bm{W}})$ and of the max-norm margin $\gamma$. ∎
The lemma completes the first of our two tasks: lower bounding the gradient’s dual norm. Importantly, the factor appearing in front of ${\mathcal{G}}({\bm{W}})$ is the max-norm margin $\gamma$, which will prove crucial in the forthcoming margin analysis.
${\mathcal{G}}({\bm{W}})$ and second-order term
We now show how to bound the second-order term in (6). For this, we establish the following essential lemma.
Lemma 4.2.
For any $\bm{s}\in\Delta^{k-1}$ in the $k$-dimensional simplex, any index $c\in[k]$, and any ${\bm{v}}\in\mathbb{R}^{k}$ it holds:
${\bm{v}}^{\top}\left(\operatorname{diag}(\bm{s})-\bm{s}\bm{s}^{\top}\right){\bm{v}}\leq 4\,(1-s_{c})\,\|{\bm{v}}\|_{\infty}^{2}\,.$
Proof.
By Cauchy-Schwartz,
${\bm{v}}^{\top}$ $\left(\operatorname{diag}(\bm{s})-\bm{s}\bm{s}^{\top}\right){\bm{v}}=\langle\operatorname{diag}(\bm{s})-\bm{s}\bm{s}^{\top},{\bm{v}}{\bm{v}}^{\top}\rangle$
$\leq\,{\left\|\operatorname{diag}(\bm{s})-\bm{s}\bm{s}^{\top}\right\|_{\rm{sum}}}\,\|{\bm{v}}\|_{\infty}^{2}\,.$
Direct calculation yields ${\left\|\operatorname{diag}(\bm{s})-\bm{s}\bm{s}^{\top}\right\|_{\rm{sum}}}=2\sum\nolimits_{c\in[k]}s_{c}(1-s_{c})$. The advertised bound then follows by noting the following $\sum_{c\in[k]}s_{c}(1-s_{c})\leq 2(1-s_{c^{\prime}})$ for any $c^{\prime}\in[k]$ (verified in Lemma A.4). ∎
To bound the second-order term in Eq. (6), we can apply the above lemma with ${\bm{v}}\leftarrow\bm{\Delta}\bm{h}_{i}$ and $c\leftarrow y_{i}$ and further use the inequality $\|{\bm{\Delta}\bm{h}_{i}}\|_{\infty}\leq{\left\|\bm{\Delta}\right\|_{\max}}\|\bm{h}_{i}\|_{1}\leq B{\left\|\bm{\Delta}\right\|_{\max}}$ (the last step invoked Ass. 3.5). This yields an upper bound
$2B^{2}\|{\bm{\Delta}}\|_{\max}^{2}\cdot\frac{1}{n}\sum_{i\in[n]}(1-\mathbb{S}_{y_{i}}({\bm{W}}\bm{h}_{i}))$
for the second-order term in the CE loss expansion in terms of the proxy function ${\mathcal{G}}({\bm{W}})$. Thus, we have fullfilled the first desiderate by finding ${\mathcal{G}}({\bm{W}})$ that simultaneously bounds the first and second-order terms in Eq. (6).
Properties of ${\mathcal{G}}({\bm{W}})$
We now show that the function ${\mathcal{G}}({\bm{W}})$ in Eq. (7) meets the second desiderata: being a good proxy for the loss $\mathcal{L}({\bm{W}})$. This is rooted in elementary relationships between ${\mathcal{G}}({\bm{W}})$ and $\mathcal{L}({\bm{W}})$, which play crucial roles in the various parts of the proof. Below, we summarize this key relationships.
Lemma 4.3 (Properties of ${\mathcal{G}}({\bm{W}})$ and $\mathcal{L}({\bm{W}})$).
Let ${\bm{W}}\in\mathbb{R}^{k\times d}$. The following hold: (i) Under Ass. 3.5, $2B\cdot{\mathcal{G}}({\bm{W}})\geq{\left\|\nabla\mathcal{L}({\bm{W}})\right\|_{\rm{sum}}}$; (ii) $1\geq\frac{{\mathcal{G}}({\bm{W}})}{\mathcal{L}({\bm{W}})}\geq 1-\frac{n\mathcal{L}({\bm{W}})}{2}$; (iii) If ${\bm{W}}$ satisfies $\mathcal{L}({\bm{W}})\leq\frac{\log 2}{n}$ or ${\mathcal{G}}({\bm{W}})\leq\frac{1}{2n}$, then $\mathcal{L}({\bm{W}})\leq 2{\mathcal{G}}({\bm{W}})$.
Lemma 4.3 (i) extends Lemma 4.1 by establishing a sandwich relationship between ${\mathcal{G}}({\bm{W}})$ and the gradient’s dual norm. The lemma’s statements (ii) and (iii) show that ${\mathcal{G}}({\bm{W}})$ can substitute for the loss - it lower bounds $\mathcal{L}({\bm{W}})$ and serves as an upper bound when either $\mathcal{L}({\bm{W}})$ or ${\mathcal{G}}({\bm{W}})$ is sufficiently small. Specifically, the ratio ${{\mathcal{G}}({\bm{W}})}\big{/}{\mathcal{L}({\bm{W}})}$ converges to 1 as the loss decreases, with the convergence rate depending on the rate of decrease. The key property (ii) may seem algebraically complex, but it turns out (Lemma B.3) that both sides of the sandwich relationship follow from the elementary fact that $\forall x>0:1-x\leq e^{-x}\leq 1-x+x^{2}/2$.
5 Implicit Bias of SignGD
We now leverage our construction of ${\mathcal{G}}({\bm{W}})$ to show that the margin of SignGD’s iterates converges to max-norm margin. We only highlight the key steps in the proof and defer details to Appendix C.
SignGD Descent
We start by showing a descent property. By applying Lemmas 4.1 and 4.2 to lower and upper bound the first and second order terms in Eq. (6) yields:
$\mathcal{L}({\bm{W}}_{t+1})$ $\leq\mathcal{L}({\bm{W}}_{t})-\gamma\eta_{t}{\mathcal{G}}({\bm{W}}_{t})+$
$\quad\quad 2\eta_{t}^{2}B^{2}{\mathcal{G}}({\bm{W}}_{t})\sup_{\zeta\in[0,1]}\frac{{\mathcal{G}}({\bm{W}}_{t}+\zeta\bm{\Delta}_{t})}{{\mathcal{G}}({\bm{W}}_{t})}.$
Algebraic manipulations of the definition of ${\mathcal{G}}({\bm{W}})$ allows us to bound the ratio in the right hand side.
Lemma 5.1 (Ratio of ${\mathcal{G}}({\bm{W}})$).
For any $\psi\in[0,1]$, we have the following: $\frac{{\mathcal{G}}({\bm{W}}+\psi\triangle{\bm{W}})}{{\mathcal{G}}({\bm{W}})}\leq e^{2B\psi{\left\|\triangle{\bm{W}}\right\|_{\max}}}$.
From this and the fact ${\left\|\bm{\Delta}_{t}\right\|_{\max}}\leq\eta_{t}$ for SignGD, we obtain
$\mathcal{L}({\bm{W}}_{t+1})$ $\leq\mathcal{L}({\bm{W}}_{t})-\gamma\eta_{t}(1-\alpha_{s_{1}}\eta_{t}){\mathcal{G}}({\bm{W}}_{t}),$ (8)
where $\alpha_{s_{1}}=2B^{2}e^{2B\eta_{0}}/\gamma$. Given a decay learning rate of the form $\eta_{t}=\Theta(\frac{1}{t^{a}})$, we can conclude that the loss starts to monotonically decrease after some time.
SignGD Unnormalized Margin
We now use the descent property in (8) to lower bound the unnormalized margin. An intermediate result towards this is recognizing that sufficiently small loss $\mathcal{L}({\bm{W}})\leq\frac{\log 2}{n}$ guarantees ${\bm{W}}$ separates the data (Lemma B.4). The descent property ensures that SignGD iterates will eventually achieve this loss threshold, thereby guaranteeing separability. The main result of this section, shows that eventually the iterates achieve separability with a substantial margin.
Lemma 5.2 (SignGD Unnormalized Margin).
Assume $\tilde{t}$ such that $\mathcal{L}({\bm{W}}_{t})\leq\frac{\log 2}{n},\forall t>\tilde{t}$. Then, the minimum unnormalized margin $\min_{i\in[n],c\neq y_{i}}(\bm{e}_{y_{i}}-\bm{e}_{c})^{T}{\bm{W}}_{t}\bm{h}_{i}$ of iterates ${\bm{W}}_{t},t\geq\tilde{t}$ is lower bounded by ($\alpha_{s_{2}}=2Be^{2B\eta_{0}}$)
$\gamma\sum_{s=\tilde{t}}^{t-1}\eta_{s}\frac{{\mathcal{G}}({\bm{W}}_{s})}{\mathcal{L}({\bm{W}}_{s})}-\alpha_{s_{2}}\sum_{s=\tilde{t}}^{t-1}\eta_{s}^{2}.$ (9)
Proof.
By exponentiating the unnormalized margin:
$e^{-\min_{i\in[n],c\neq y_{i}}(\bm{e}_{y_{i}}-\bm{e}_{c})^{T}{\bm{W}}_{t}\bm{h}_{i}}=\max_{i\in[n]}e^{-\min_{c\neq y_{i}}(\bm{e}_{y_{i}}-\bm{e}_{c})^{T}{\bm{W}}_{t}\bm{h}_{i}}$
$\stackrel{{\scriptstyle(a)}}{{\leq}}\max_{i\in[n]}\frac{1}{\log 2}\log\bigl{(}1+e^{-\min_{c\neq y_{i}}(\bm{e}_{y_{i}}-\bm{e}_{c})^{T}{\bm{W}}_{t}\bm{h}_{i}}\bigr{)}$
$\leq\max_{i\in[n]}\frac{1}{\log 2}\log(1+\sum_{c\neq y_{i}}e^{-(\bm{e}_{y_{i}}-\bm{e}_{c})^{T}{\bm{W}}_{t}\bm{h}_{i}})\leq\frac{n\mathcal{L}({\bm{W}}_{t})}{\log 2}$
$\stackrel{{\scriptstyle(b)}}{{\leq}}\exp\bigl{(}-\gamma\sum_{s=\tilde{t}}^{t-1}\eta_{s}\frac{{\mathcal{G}}({\bm{W}}_{s})}{\mathcal{L}({\bm{W}}_{s})}+\alpha_{s_{2}}\sum_{s=\tilde{t}}^{t-1}\eta_{s}^{2}\bigr{)}.$
(a) is because $\frac{\log(1+e^{-z})}{e^{-z}}\geq\log 2,z\geq 0$ with $z$ chosen to be $\min_{c\neq y_{i}}(\bm{e}_{y_{i}}-\bm{e}_{c})^{T}{\bm{W}}_{t}\bm{h}_{i}$ for any $i\in[n]$; (b) is by some manipulations of (8) (details in App. C.2). ∎
SignGD Margin Convergence
Proceeding from Eq. (9) requires showing convergence of the ratio $\frac{{\mathcal{G}}({\bm{W}})}{\mathcal{L}({\bm{W}})}$. The two key ingredients are given in Lemma 4.3 (ii) and (iii). Lemma 4.3 (ii) suggests that it is sufficient to study the convergence of $\mathcal{L}({\bm{W}})$, which is captured in (8). However, to obtain an explicit rate via (8), we need to rewrite ${\mathcal{G}}({\bm{W}}_{t})$ in terms of $\mathcal{L}({\bm{W}}_{t})$. This is where Lemma 4.3 (iii) helps. Putting them together, we arrive at the following theorem.
Theorem 5.3.
Suppose that Ass. 3.1, 3.3, and 3.5 hold, then there exists $t_{s_{2}}=t_{s_{2}}(n,\gamma,B,{\bm{W}}_{0})$ such that the margin gap $\gamma-{\min_{i\in[n],c\neq y_{i}}(\bm{e}_{y_{i}}-\bm{e}_{c})^{T}{\bm{W}}_{t}\bm{h}_{i}}/{\left\|{\bm{W}}_{t}\right\|_{\max}}$ of SignGD’s iterates for all $t>t_{s_{2}}$ is upper bounded by
$\mathcal{O}\Big{(}\frac{\sum_{s=t_{s_{2}}}^{t-1}\eta_{s}e^{-\frac{\gamma}{4}\sum_{\tau=t_{s_{2}}}^{s-1}\eta_{\tau}}+\sum_{s=0}^{t_{s_{2}}-1}\eta_{s}+\sum_{s=t_{s_{2}}}^{t-1}\eta_{s}^{2}}{\sum_{s=0}^{t-1}\eta_{s}}\Big{)}.$
Proof.
SignGD’s update rule (4) gives ${\left\|{\bm{W}}_{t}\right\|_{\max}}\leq{\left\|{\bm{W}}_{0}\right\|_{\max}}+\sum_{s=0}^{t-1}\eta_{s}$. We first find iteration $t_{s_{1}}$ via (8) such that $\mathcal{L}({\bm{W}}_{t+1})\leq\mathcal{L}({\bm{W}}_{t})-\frac{\eta_{t}\gamma}{2}{\mathcal{G}}({\bm{W}}_{t})$ for all $t\geq t_{s_{1}}$. Then, we find iteration $t_{s_{2}}>t_{s_{1}}$ such that $\mathcal{L}({\bm{W}}_{t})\leq\frac{\log 2}{n}$ for all $t\geq t_{s_{2}}$. By Lemma 4.3 (iii), this guarantees $\mathcal{L}({\bm{W}}_{t})\leq 2{\mathcal{G}}({\bm{W}}_{t})$. Substituting it to the above recursion on $\mathcal{L}({\bm{W}}_{t})$, we obtain $\mathcal{L}({\bm{W}}_{t})\leq\frac{\log 2}{n}e^{-\frac{\gamma}{4}\sum_{s=t_{s_{2}}}^{t-1}\eta_{s}}$. By Lemma 4.3 (ii), we further obtain $\frac{{\mathcal{G}}({\bm{W}}_{t})}{\mathcal{L}({\bm{W}}_{t})}\geq 1-e^{-\frac{\gamma}{4}\sum_{s=t_{s_{2}}}^{t-1}\eta_{s}}$. Combining this with (9) and the upper bound on ${\left\|{\bm{W}}_{t}\right\|_{\max}}$, the final result is proved (details in App. C.4). ∎
In the proof, $t_{s_{1}}$ can be set to $(4B^{2}e^{2B\eta_{0}}/{\gamma})^{1/a}$, and $t_{s_{2}}$ is chosen such that $t_{s_{2}}\leq\Theta(n^{\frac{1}{1-a}}+n^{\frac{1}{1-a}}\mathcal{L}({\bm{W}}_{0})^{\frac{1}{1-a}})$. Note that both are independent of the problem’s dimensionality, yielding the following convergence rates.
Corollary 5.4.
Set learning rate $\eta_{t}=\Theta(\frac{1}{t^{a}}),a\in(0,1]$. Under the setting of Theorem 5.3, the margin gap of SignGD iterates with $a=1/2$ reduces at rate $\mathcal{O}(\frac{\log t+n}{t^{1/2}})$. [^1] The rates for other values of $a$ can be found in Corollary C.5.
Remark 5.5.
These results generalize to NSD defined in (5). Concretely, Lemma 4.1, 4.2, and 4.3 (i) generalize to any $p$-norm (where margin in (2) is defined w.r.t. the same norm) together with the dual norm on the loss gradient. These results are summarized in Lemmas E.1 and E.3 in App. E. In terms of margin convergence of NSD, (Nacson et al., 2019) showed a rate of $\mathcal{O}(\frac{\log t}{t^{1/2}})$ in the binary setting, limited to the exponential loss. Compared to this, our results hold for the more practical setting of multilcass data and CE loss. The rate in Cor. 5.4 surpasses the max-norm margin convergence rate of Adam in the binary setting (Zhang et al., 2024); thus even in the binary setting it improves the latter and extends (Nacson et al., 2019)’s result to logistic loss. We elaborate on this gap in the next section.
6 Implicit Bias of Adam
Our analysis of Adam for multiclass data, while inspired by the binary setting of (Zhang et al., 2024), requires significant new technical insights to achieve the same convergence rate without class-dependent factors. While the proof follows similar high-level steps as SignGD above and leverages all the key properties of ${\mathcal{G}}({\bm{W}})$ identified in Sec. B, using ${\mathcal{G}}({\bm{W}})$ alone would lead to suboptimal bounds with an extra factor of $k$. Our crucial insight lies in a per-class decomposition of ${\mathcal{G}}({\bm{W}})$ as follows:
${\mathcal{G}}({\bm{W}})$ $=\sum_{c\in[k]}\frac{1}{n}\sum_{i\in[n],y_{i}=c}(1-s_{iyi})$
$=\sum_{c\in[k]}\frac{1}{n}\sum_{i\in[n],y_{i}\neq c}s_{ic},$
where recall $s_{ic}\coloneqq\mathbb{S}_{c}({\bm{W}}\bm{h}_{i})$. This decomposition motivates further defining the “per-class proxies”:
${\mathcal{G}}_{c}({\bm{W}})$ $\coloneqq\frac{1}{n}\sum_{i\in[n],y_{i}=c}(1-s_{iy_{i}})$
$\mathcal{Q}_{c}({\bm{W}})$ $\coloneqq\frac{1}{n}\sum_{i\in[n],y_{i}\neq c}s_{ic}.$
Both terms play crucial roles in our per-class analysis, with this decomposition being key to avoiding the factor $k$ in our bounds. The remainder of this section highlights the technical challenges beyond those encountered in analyzing SignGD (full proof details are provided in Appendix D).
Adam Descent
To show descent of Adam via Taylor expansion (6), the first-order term involves an additional factor when introducing the dual norm of the gradient, i.e.,
$\langle\nabla\mathcal{L}({\bm{W}}_{t}),\bm{\Delta}_{t}\rangle\leq-\eta_{t}{\left\|\nabla\mathcal{L}({\bm{W}}_{t})\right\|_{\rm{sum}}}$
$\qquad+\eta_{t}\underbrace{\bigm{|}\langle\nabla\mathcal{L}({\bm{W}}_{t}),\frac{\bm{M}_{t}}{\sqrt{{\bm{V}}_{t}}}-\frac{\nabla\mathcal{L}({\bm{W}}_{t})}{|\nabla\mathcal{L}({\bm{W}}_{t})|}\rangle\bigm{|}}_{\clubsuit}.$ (10)
Given that we bound the remaining terms in (6) using ${\mathcal{G}}({\bm{W}})$, it seems natural to bound the $\clubsuit$ term similarly. While one could attempt to relate the entries of both matrices $\bm{M}_{t}$ and ${\bm{V}}_{t}$ to ${\mathcal{G}}({\bm{W}}_{t})$, our entry-wise analysis of the $\clubsuit$ term (following (Zhang et al., 2024)’s approach in the binary setting) would inevitably introduce an extra factor of $k$. To avoid this, we instead relate the row vectors $\bm{M}_{t}[c,:]$ and ${\bm{V}}_{t}[c,:]$ to the class-specific functions ${\mathcal{G}}_{c}({\bm{W}})$ and $\mathcal{Q}_{c}({\bm{W}})$ for each $c\in[k]$. This class-wise approach requires careful accounting of interactions between the $k>2$ classes throughout our analysis. We first show how to bound $\bm{M}_{t}$.
Lemma 6.1.
Let $c\in[k]$. Under the setting of Theorem 6.3, there exists time $t_{0}$ such that for all $t\geq t_{0}$:
$|\mathbf{M}_{t}[c,j]-(1-\beta_{1}^{t+1})$ $\nabla\mathcal{L}({\bm{W}}_{t})[c,j]|\leq$
$\alpha_{M}\eta_{t}\bigl{(}{\mathcal{G}}_{c}({\bm{W}}_{t})+\mathcal{Q}_{c}({\bm{W}}_{t})\bigr{)},$
where $j\in[d]$, and $\alpha_{M}:=B(1-\beta_{1})c_{2}$.
Here, we provide a quick proof sketch. By the Adam update rule 3a, we have for all $c\in[k]$ and $j\in[d]$: the following $|\mathbf{M}_{t}[c,j]-(1-\beta_{1}^{t+1})\nabla\mathcal{L}({\bm{W}}_{t})[c,j]|$ can be bounded by
$\sum_{\tau=0}^{t}(1-\beta_{1})\beta_{1}^{\tau}\underbrace{|\nabla\mathcal{L}({\bm{W}}_{t-\tau})[c,j]-\nabla\mathcal{L}({\bm{W}}_{t})[c,j]|}_{\spadesuit}.$
By explicitly writing out the gradient and grouping terms we can show:
$\spadesuit$ $\leq B\frac{1}{n}\sum\nolimits_{i\in[n]}|\mathbb{S}_{c}({\bm{W}}_{t-\tau}\bm{h}_{i})-\mathbb{S}_{c}({\bm{W}}_{t}\bm{h}_{i})|$
$=B\underbrace{\frac{1}{n}\sum\nolimits_{i\in[n]:y_{i}\neq c}|\mathbb{S}_{c}({\bm{W}}_{t-\tau}\bm{h}_{i})-\mathbb{S}_{c}({\bm{W}}_{t}\bm{h}_{i})|}_{\spadesuit_{1}}$
$\quad\quad+B\underbrace{\frac{1}{n}\sum\nolimits_{i\in[n]:y_{i}=c}|\mathbb{S}_{y_{i}}({\bm{W}}_{t-\tau}\bm{h}_{i})-\mathbb{S}_{y_{i}}({\bm{W}}_{t}\bm{h}_{i})|}_{\spadesuit_{2}}.$
In the above equality, we split the sum into two cases: samples where $y_{i}=c$ and the rest. Using the definitions of ${\mathcal{G}}_{c}({\bm{W}})$ and $\mathcal{Q}_{c}({\bm{W}})$, we manipulate and express the first ($\spadesuit_{1}$) and second ($\spadesuit_{2}$) terms via $\mathcal{Q}_{c}({\bm{W}})$ and ${\mathcal{G}}_{c}({\bm{W}})$ respectively. This leaves us to bound expressions of the form $|\mathbb{S}_{c}({\bm{v}}^{\prime})/\mathbb{S}_{c}({\bm{v}})-1|$ and $|(1-\mathbb{S}_{c}({\bm{v}}^{\prime}))/(1-\mathbb{S}_{c}({\bm{v}}))-1|$ respectively for the two terms. We address this technical challenge in Lemma D.10, showing how to bound such terms using $\|{\bm{v}}-{\bm{v}}^{\prime}\|_{\infty}$, which we can control for Adam. We now show how to bound ${\bm{V}}_{t}$.
Lemma 6.2.
Let $c\in[k]$. Under the setting of Theorem 6.3, there exists time $t_{0}$ such that for all $t\geq t_{0}$:
$\bigm{|}\sqrt{{\bm{V}}_{t}[c,j]}-\sqrt{(1-\beta_{2}^{t+1})}|$ $\nabla\mathcal{L}({\bm{W}}_{t})[c,j]|\bigm{|}\leq$
$\alpha_{V}\sqrt{\eta_{t}}\bigl{(}{\mathcal{G}}_{c}({\bm{W}}_{t})+\mathcal{Q}_{c}({\bm{W}}_{t})\bigr{)},$
where $j\in[d]$, and $\alpha_{V}=B\sqrt{(1-\beta_{2})c_{2}}$.
By Adam’s update rule (3c), we have $\forall c\in[k],j\in[d]$ that $|{\bm{V}}_{t}[c,j]-(1-\beta_{2}^{t+1})\nabla\mathcal{L}({\bm{W}}_{t})[c,j]^{2}|$ can be bounded by
$\sum_{\tau=0}^{t}(1-\beta_{2})\beta_{2}^{\tau}\underbrace{|\nabla\mathcal{L}({\bm{W}}_{t-\tau})[c,j]^{2}-\nabla\mathcal{L}({\bm{W}}_{t})[c,j]^{2}|}_{\diamond}.$
For the $\diamond$ term, we compute $\nabla\mathcal{L}({\bm{W}})[c,j]^{2}$ and define the function $f_{c,i,p}({\bm{W}}):=(\delta_{cy_{i}}-\mathbb{S}_{c}({\bm{W}}\bm{h}_{i}))(\delta_{cy_{p}}-\mathbb{S}_{c}({\bm{W}}\bm{h}_{p}))$ to obtain (after some algebra):
$\diamond\leq\frac{B^{2}}{n^{2}}\sum\nolimits_{i\in[n]}\sum\nolimits_{p\in[n]}|f_{c,i,p}({\bm{W}}_{t-\tau})-f_{c,i,p}({\bm{W}}_{t})|.$
Unlike the first moment $\bm{M}_{t}$, this summation involves pairs of sample indices $i$ and $p$ with label $c$ taking values equal to one of four cases: $y_{i},\text{non-}y_{i},y_{p},\text{non-}y_{p}$. Following our first-moment analysis strategy, we decompose the sum into these four components. The resulting bounds involve squared terms ${\mathcal{G}}_{c}({\bm{W}})^{2}$, $\mathcal{Q}_{c}({\bm{W}})^{2}$, and ${\mathcal{G}}_{c}({\bm{W}})\mathcal{Q}_{c}({\bm{W}})$ due to the quadratic softmax expressions in each component (see Lemma D.10 for bounding ratios of softmax quadratics).
With Lemmas 6.1 and 6.2 at hand, we can now bound the Adam-specific term $\clubsuit$ in Eq. (10) in terms of ${\mathcal{G}}({\bm{W}})$. The essence is to do the double sum (over $c\in[k]$ and $j\in[d]$) in $\clubsuit$ in two stages: first sum over $d$ for each $c\in[k]$ with the help of Lemma 6.1 and 6.2; then sum over $c$ with the recognition of ${\mathcal{G}}({\bm{W}})=\sum_{c\in[k]}{\mathcal{G}}_{c}({\bm{W}})=\sum_{c\in[k]}\mathcal{Q}_{c}({\bm{W}})$. The results are summarized in Lemma D.4. From this point on, following similar steps of SignGD, we can show the CE loss (eventually) monotonically decreases (see Lemma D.5).
Adam Implicit Bias
To establish Adam’s implicit bias, we follow the same strategy as SignGD: bound the margin and control weight growth. For the margin bound, we leverage our descent property from above (details in Lemma D.6). Unlike SignGD’s simple updates, Adam’s moment averaging makes bounding ${\left\|{\bm{W}}_{t}\right\|_{\max}}$ more challenging. Our solution again relies on the proxy ${\mathcal{G}}({\bm{W}})$, which we have shown approximates both loss and gradients well. Specifically, using this, we prove that the second moment ${\bm{V}}_{t}$ remains controlled when the loss is small, i.e. $\forall c\in[k],j\in[d]$:
${\bm{V}}_{t}[c,j]$ $\leq\nabla\mathcal{L}({\bm{W}}_{t})[c,j]^{2}+\alpha_{V}\eta_{t}{\mathcal{G}}({\bm{W}}_{t})^{2}$
$\leq 4B^{2}{\mathcal{G}}({\bm{W}}_{t})^{2}+\alpha_{V}\eta_{0}{\mathcal{G}}({\bm{W}}_{t})^{2}$
$\leq(4B^{2}+\alpha_{V}\eta_{0})\mathcal{L}({\bm{W}}_{t})^{2},$
where the penultimate and last inequalities are by Lemma 4.3 (i) and (ii), respectively. This suggests that with sufficiently small loss, all entries of ${\bm{V}}_{t}$ remain bounded by 1. Building on our bound for ${\bm{V}}_{t}$, we apply Zhang et al. (2024, Lemma A.4) (see also Xie & Li (2024, Lemma 4.2)), which connects ${\bm{W}}_{t}$ and $\log({\bm{V}}_{t})$ entry-wise. This allows us to bound ${\left\|{\bm{W}}_{t}\right\|_{\max}}$ using learning rate sums under Ass. 3.2 (details in Lemma D.7). With the convergence $\frac{{\mathcal{G}}({\bm{W}})}{\mathcal{L}({\bm{W}})}\rightarrow 1$ following as in SignGD, we arrive at our main theorem.
Theorem 6.3.
Suppose that Ass. 3.1-3.5 hold, and $\beta_{1}\leq\beta_{2}$. There exists $t_{a_{2}}=t_{a_{2}}(n,d,\gamma,B,{\bm{W}}_{0},\beta_{1},\beta_{2},\omega)$ such that the margin gap $\gamma-{\min_{i\in[n],c\neq y_{i}}(\bm{e}_{y_{i}}-\bm{e}_{c})^{T}{\bm{W}}_{t}\bm{h}_{i}}/{\left\|{\bm{W}}_{t}\right\|_{\max}}$ of Adam’s iterates for for all $t>t_{a_{2}}$ is upper bounded by
$\mathcal{O}(\frac{\sum_{s=t_{a_{2}}}^{t-1}\eta_{s}e^{-\frac{\gamma}{4}\sum_{\tau=t_{a_{2}}}^{s-1}\eta_{\tau}}+\sum_{s=0}^{t_{a_{2}}-1}\eta_{s}+d\sum_{s=t_{a_{2}}}^{t-1}\eta_{s}^{3/2}}{\sum_{s=0}^{t-1}\eta_{s}}).$
The referenced iterations are $t_{a_{1}}=\Theta(d^{2/a})$ and $t_{a_{2}}\leq\Theta(n^{\frac{1}{1-a}}d^{2/a}+n^{\frac{1}{1-a}}\mathcal{L}({\bm{W}}_{0})^{\frac{1}{1-a}}+\log(1/\omega))$. Note that compared to the respective iterations for SignGD, $t_{a_{1}}$ depends on $d$, unlike $t_{s_{1}}$. These values the explicit rates below.
Corollary 6.4.
Set learning rate $\eta_{t}=\Theta(\frac{1}{t^{a}}),a\in(0,1]$. Under the setting of Theorem 6.3, the margin gap of Adam iterates with $a=2/3$ reduces at rate $\mathcal{O}(\frac{d\log(t)+nd+[\log(1/\omega)]^{1/3}}{t^{1/3}})$. [^2] The rates for other values of $a$ can be found in Corollary D.9.
Remark 6.5.
These rates exactly match those in the binary case of (Zhang et al., 2024) with logarithmic dependence on the initialization parameter $\omega$ (Ass. 3.2). This is only made possible through the fine-grained per-class bounding of the first and second moments using both ${\mathcal{G}}_{c}({\bm{W}})$ and $\mathcal{Q}_{c}({\bm{W}})$. Note that Lemma D.4 takes the same form as Zhang et al. (2024, Lemma A.4). However, without the tight per-class bound and the equivalent decomposition of ${\mathcal{G}}({\bm{W}})$ using either $\mathcal{Q}_{c}({\bm{W}})$ or ${\mathcal{G}}_{c}({\bm{W}})$, an extra factor of $k$ would appear. Interestingly, our rates for SignGD in Corollary 5.4 reveal a theoretical gap: Adam’s optimal choice $a=\frac{2}{3}$ yields $\mathcal{O}(\frac{d\log(t)+nd}{t^{1/3}})$ while SignGD achieves $\mathcal{O}(\frac{\log(t)+n}{t^{1/2}})$ with $a=\frac{1}{2}$. Despite achieving tightness w.r.t. class-dimension ($k$), this gap emerges from our entry-wise analysis of the $\clubsuit$ term in (10) across the feature dimension $(d)$ using scalar functions ${\mathcal{G}}_{c}({\bm{W}}),\mathcal{Q}_{c}({\bm{W}})$. Closing this theoretical gap–revealed through our SignGD analysis–that also appears in the binary case (Zhang et al., 2024), forms an important direction for future work.
Experiments
We generate snythetic multiclass separable data as follows: $k=5$ class centers are sampled from a standard normal distribution; within each class, data is sampled from normal distribution $\mathcal{N}(0,\sigma^{2}I),\sigma=0.1$. We set $d=25$, sample $5$ data points for each class, and ensure that margin is positive (thus data is separable). We run different algorithms to minimize CE loss using $\eta_{t}=\frac{\eta_{0}}{t^{a}}$ ($\eta_{0}=0.01$), where (based on our theorems) $a$ is set to $\nicefrac{{1}}{{2}}$, $\nicefrac{{1}}{{2}}$, and $\nicefrac{{2}}{{3}}$ for NGD, SignGD, and Adam, respectively. The stability constant $\epsilon$ for Adam is set from $\{0,10^{-6},10^{-7},10^{-8}\}$. We denote max-margin classifiers defined w.r.t. the 2-norm and the max-norm as ${\bm{V}}_{2}$ and ${\bm{V}}_{\infty}$ respectively. Fig. LABEL:fig:main_fig1 shows the following: (1) SignGD/Adam iterates favor the the max-norm margin over the 2-norm margin. The opposite is true for NGD (Figs. LABEL:fig:l2_margin,LABEL:fig:max_margin); (2) SignGD/Adam iterates correlate well with ${\bm{V}}_{\infty}$ whereas NGD correlates better with ${\bm{V}}_{2}$ (Figs. LABEL:fig:cor_inf,LABEL:fig:cor_v2); (3) Results are consistent for different (small) values of stability constant (curves nearly overlap).
7 Conclusion
We have characterized the implicit bias of SignGD, NSD, and Adam for multiclass separable data with CE loss, providing explicit rates for their margin maximization behavior. While these results establish fundamental theoretical guarantees, they are limited to linear models and separable data. Yet, they open several promising directions for future research: (1) Improving SignGD’s rates through appropriate momentum leveraging ideas from Ji et al. (2021); Wang et al. (2023) for NGD. (2) Extending the results to non-separable data (Ji & Telgarsky, 2019) and soft-label classification (Thrampoulidis, 2024) settings. (3) Extending the analysis to non-linear architectures like self-attention (Tarzanagh et al., 2023b, a; Vasudeva et al., 2024; Julistiono et al., 2024). Our treatment of CE loss through softmax properties in Sec. B is particularly relevant, as the implicit bias in self-attention is fundamentally driven by the softmax map. Moreover, self-attention naturally aligns with our multiclass setting since each token attends to multiple other tokens through softmax. (4) Investigating whether different implicit biases of SignGD/Adam to GD could theoretically explain empirical observations about their superior performance on heavy-tailed, imbalanced multiclass data (Kunstner et al., 2024). (5) From a statistical perspective, identifying specific scenarios where max-norm margin maximization leads to better generalization (building on initial investigations from (Salehi et al., 2019; Varma & Hassibi, 2024; Akhtiamov et al., 2024) on the impact of different regularization forms in linear classification, and recent findings from (Mohamadi et al., 2024) specifically related to infinity-norm regularization in neural networks).
Acknowledgement
This work is funded partially by NSERC Discovery Grants RGPIN-2021-03677 and RGPIN-2022-03669, Alliance GrantALLRP 581098-22, a CIFAR AI Catalyst grant, and the Canada CIFAR AI Chair Program.
Appendix A Facts about CE loss and Softmax
Lemma A.1 is on the gradient of the cross-entropy loss. It will be used for showing the form of ${\mathcal{G}}({\bm{W}})$ in (7) lower bounds ${\left\|\nabla\mathcal{L}({\bm{W}})\right\|_{\max}}$ in Lemma B.1.
Lemma A.1 (Gradient).
Let CE loss
$\mathcal{L}({\bm{W}}):=-\frac{1}{n}\sum_{i\in[n]}\log\big{(}\mathbb{S}_{{y_{i}}}({\bm{W}}\bm{h}_{i})\big{)}.$
For any ${\bm{W}}$, it holds
  • $\nabla\mathcal{L}({\bm{W}})=-\frac{1}{n}\sum_{i\in[n]}\left(\bm{e}_{y_{i}}-\bm{s}_{i}\right)\bm{h}_{i}^{\top}=-\frac{1}{n}(\bm{Y}-\bm{S}){\bm{H}}^{\top}$
  • $\mathds{1}_{k}^{\top}\nabla\mathcal{L}({\bm{W}})=0$
  • For any matrix ${\bm{A}}\in\mathbb{R}^{k\times d}$,
    $\langle\bm{A},-\nabla\mathcal{L}({\bm{W}})\rangle$ $=\frac{1}{n}\sum_{i}\left(1-s_{i{y_{i}}}\right)\left(\bm{e}_{y_{i}}^{\top}\bm{A}\bm{h}_{i}-\frac{\sum_{c\neq{y_{i}}}s_{ic}\,\bm{e}_{c}^{\top}\bm{A}\bm{h}_{i}}{\left(1-s_{i{y_{i}}}\right)}\right)$
    $=\frac{1}{n}\sum_{i\in[n]}{\sum_{c\neq{y_{i}}}s_{ic}\,(\bm{e}_{{y_{i}}}-\bm{e}_{c})^{\top}\bm{A}\bm{h}_{i}}$ (11)
where we simplify $\bm{S}:=\mathbb{S}({\bm{W}}{\bm{H}})=[\bm{s}_{1},\ldots,\bm{s}_{n}]\in\mathbb{R}^{k\times n}$. The last statement yields
$\langle\bm{A},-\nabla\mathcal{L}({\bm{W}})\rangle\geq\frac{1}{n}\sum_{i\in[n]}\left(1-s_{i{y_{i}}}\right)\,\cdot\,\min_{c\neq{y_{i}}}\left(\bm{e}_{y_{i}}-\bm{e}_{c}\right)^{\top}\bm{A}\bm{h}_{i}.$ (12)
Proof.
First bullet is by direct calculation. Second bullet uses the fact that $\mathds{1}^{\top}(\bm{y}_{i}-\bm{s}_{i})=1-1=0$ since $\mathds{1}^{\top}\bm{s}_{i}=1$. The third bullet follows by direct calculation and writing $\bm{s}_{i}^{\top}\bm{A}\bm{h}_{i}=(\sum_{c}s_{ic}\bm{e}_{c})^{\top}\bm{A}\bm{h}_{i}=\sum_{c}s_{ic}\,\bm{e}_{c}^{\top}\bm{A}\bm{h}_{i}$. ∎
Lemma A.2 is on the Taylor expansion of the loss. It will be used in showing the descent properties of SignGD (Lemma C.1) and Adam (Lemma D.5).
Lemma A.2 (Hessian).
Let perturbation $\bm{\Delta}\in\mathbb{R}^{k\times d}$ and denote ${\bm{W}}^{\prime}={\bm{W}}+\bm{\Delta}$. Then,
$\mathcal{L}({\bm{W}}^{\prime})$ $=\mathcal{L}({\bm{W}})-\frac{1}{n}\sum_{i\in[n]}\langle(\bm{e}_{y_{i}}-\mathbb{S}({\bm{W}}\bm{h}_{i}))\bm{h}_{i}^{\top},\bm{\Delta}\rangle$
$\quad+\frac{1}{2n}\sum_{i\in[n]}\bm{h}_{i}^{\top}\bm{\Delta}^{\top}\left(\operatorname{diag}(\mathbb{S}({\bm{W}}\bm{h}_{i}))-\mathbb{S}({\bm{W}}\bm{h}_{i})\mathbb{S}({\bm{W}}\bm{h}_{i})^{\top}\right)\bm{\Delta}\,\bm{h}_{i}+o(\|\bm{\Delta}\|^{3})\,.$ (13)
Proof.
Define function $\ell_{y}:\mathbb{R}^{k}\rightarrow\mathbb{R}$ parameterized by $y\in[k]$ as follows:
$\ell_{y}(\bm{l}):=-\log(\mathbb{S}_{y}(\bm{l}))\,.$
From Lemma A.1,
$\nabla\ell_{y}(\bm{l})=-(\bm{e}_{y}-\mathbb{S}(\bm{l}))\,.$
Thus,
$\nabla^{2}\ell_{y}(\bm{l})=\nabla\mathbb{S}(\bm{l})=\operatorname{diag}(\mathbb{S}(\bm{l}))-\mathbb{S}(\bm{l})\mathbb{S}(\bm{l})^{\top}$
Combining these the second-order taylor expansion of $\ell_{y}$ writes as follows for any $\bm{l},\bm{\delta}\in\mathbb{R}^{k}$:
$\ell_{y}(\bm{l}+\bm{\delta})=\ell_{y}(\bm{l})-(\bm{e}_{y}-\mathbb{S}(\bm{l}))^{\top}\bm{\delta}+\frac{1}{2}\bm{\delta}^{\top}\left(\operatorname{diag}(\mathbb{S}(\bm{l}))-\mathbb{S}(\bm{l})\mathbb{S}(\bm{l})^{\top}\right)\bm{\delta}+o(\|\bm{\delta}\|^{3})\,.$
To evaluate this with respect to a change on the classifier parameters, set $\bm{l}={\bm{W}}\bm{h}$ and $\bm{\delta}=\bm{\Delta}\bm{h}$ for $\bm{\Delta}\in\mathbb{R}^{k\times d}$. Denoting ${\bm{W}}^{\prime}={\bm{W}}+\bm{\Delta}$, we then have
$\ell_{y}({\bm{W}}^{\prime})=\ell_{y}({\bm{W}})-\langle(\bm{e}_{y}-\mathbb{S}(\bm{l}))\bm{h}^{\top},\bm{\Delta}\rangle+\frac{1}{2}\bm{h}^{\top}\bm{\Delta}^{\top}\left(\operatorname{diag}(\mathbb{S}(\bm{l}))-\mathbb{S}(\bm{l})\mathbb{S}(\bm{l})^{\top}\right)\bm{\Delta}\bm{h}+o(\|\bm{\Delta}\|^{3})\,.$
This shows the desired since $n\mathcal{L}({\bm{W}}):=\sum_{i\in[n]}\ell_{y_{i}}({\bm{W}}\bm{h}_{i})$  and we can further obtain
$\ell_{y}({\bm{W}}^{\prime})=\ell_{y}({\bm{W}})-\langle(\bm{e}_{y}-\mathbb{S}(\bm{l}))\bm{h}^{\top},\bm{\Delta}\rangle+\frac{1}{2}\bm{h}^{\top}\bm{\Delta}^{\top}\left(\operatorname{diag}(\mathbb{S}(\bm{l}^{\prime}))-\mathbb{S}(\bm{l}^{\prime})\mathbb{S}(\bm{l}^{\prime})^{\top}\right)\bm{\Delta}\bm{h},$ (14)
where $\bm{l}^{\prime}=\bm{l}+\zeta\bm{\delta}$ for some $\zeta\in[0,1]$. ∎
Lemma A.3 is used in bounding the second order term in the Taylor expansion of $\mathcal{L}({\bm{W}})$.
Lemma A.3.
For any $\bm{s}\in\Delta^{k-1}$ in the $k$-dimensional simplex, any index $c\in[k]$, and any ${\bm{v}}\in\mathbb{R}^{k}$ it holds:
${\bm{v}}^{\top}\left(\operatorname{diag}(\bm{s})-\bm{s}\bm{s}^{\top}\right){\bm{v}}\leq 4\,\|{\bm{v}}\|_{\infty}^{2}\,(1-s_{c})$
Proof.
By Cauchy-Schwartz,
${\bm{v}}^{\top}\left(\operatorname{diag}(\bm{s})-\bm{s}\bm{s}^{\top}\right){\bm{v}}$ $=\operatorname{vec}\big{(}\operatorname{diag}(\bm{s})-\bm{s}\bm{s}^{\top}\big{)}^{\top}\operatorname{vec}\big{(}{\bm{v}}{\bm{v}}^{\top}\big{)}$
$\leq\|\operatorname{vec}\big{(}\operatorname{diag}(\bm{s})-\bm{s}\bm{s}^{\top}\big{)}\|_{1}\|\operatorname{vec}\big{(}{\bm{v}}{\bm{v}}^{\top}\big{)}\|_{\infty}$
$\leq\|{\bm{v}}\|_{\infty}^{2}\,\|\operatorname{vec}\big{(}\operatorname{diag}(\bm{s})-\bm{s}\bm{s}^{\top}\big{)}\|_{1}\,.$
But,
$\|\operatorname{vec}\big{(}\operatorname{diag}(\bm{s})-\bm{s}\bm{s}^{\top}\big{)}\|_{1}$ $=\sum_{c\in[k]}s_{c}(1-s_{c})+\sum_{c^{\prime}\neq c\in[k]}s_{c}s_{c^{\prime}}$
$=2\sum_{c\in[k]}s_{c}(1-s_{c})$
where the last line uses $\sum_{c\neq c}s_{c^{\prime}}=1-s_{c}$.
The proof completes by applying Lemma A.4 to the above.
Lemma A.4 is used in the proof of Lemma A.3.
Lemma A.4.
For any $\bm{s}\in\Delta^{k-1}$ in the $k$-dimensional simplex and any index $c\in[k]$ it holds that
$\sum_{c^{\prime}}s_{c^{\prime}}(1-s_{c^{\prime}})\leq 2(1-s_{c})\,.$
Proof.
With a bit of algebra and using $\sum_{c^{\prime}\neq c}s_{c^{\prime}}=1-s_{c}$ the claim becomes equivalent to
$\sum_{c^{\prime}\neq c}s_{c^{\prime}}^{2}+s_{c}^{2}-2s_{c}+1\geq 0.$
Since this holds true, the lemma holds. ∎
Appendix B Lemmas on $\mathcal{L}({\bm{W}})$ and ${\mathcal{G}}({\bm{W}})$
Lemma B.1 shows that ${\mathcal{G}}({\bm{W}})$ lower bounds the loss gradient and that the bound becomes tight as the loss approaches zero.
Lemma B.1 (${\mathcal{G}}({\bm{W}})$ as proxy to the loss-gradient norm).
Recall the margin $\gamma$ and the assumption $\|\bm{h}_{i}\|_{1}\leq B$. Then, for any ${\bm{W}}$ it holds that
$2B\cdot{\mathcal{G}}({\bm{W}})\geq{\left\|\nabla\mathcal{L}({\bm{W}})\right\|_{\rm{sum}}}\geq\gamma\cdot{\mathcal{G}}({\bm{W}})\,.$
Proof.
First, we prove the lower bound. By duality and direct application of (12)
${\left\|\nabla\mathcal{L}({\bm{W}})\right\|_{\rm{sum}}}$ $=\max_{{\left\|\bm{A}\right\|_{\max}}\leq 1}\langle\bm{A},-\nabla\mathcal{L}({\bm{W}})\rangle$
$\geq\max_{{\left\|\bm{A}\right\|_{\max}}\leq 1}\frac{1}{n}\sum_{i\in[n]}(1-s_{iy_{i}})\min_{c\neq y_{i}}(\bm{e}_{y_{i}}-\bm{e}_{c})^{T}\bm{A}\bm{h}_{i}$
$\geq\frac{1}{n}\sum_{i\in[n]}(1-s_{iy_{i}})\cdot\max_{{\left\|\bm{A}\right\|_{\max}}\leq 1}\min_{i\in[n],c\neq y_{i}}(\bm{e}_{y_{i}}-\bm{e}_{c})^{T}\bm{A}\bm{h}_{i}.$
Second, for the upper bound, start with noting by triangle inequality that
${\left\|\nabla\mathcal{L}({\bm{W}})\right\|_{\rm{sum}}}\leq\frac{1}{n}\sum_{i\in[n]}{\left\|\nabla\ell_{i}({\bm{W}})\right\|_{\rm{sum}}}\,,$
where $\ell_{i}({\bm{W}})=-\log(\mathbb{S}_{y_{i}}({\bm{W}}\bm{h}_{i}))$. Recall that
$\nabla\ell_{i}({\bm{W}})=-(\bm{e}_{y}-\mathbb{S}_{y_{i}}({\bm{W}}\bm{h}_{i}))\bm{h}_{i}^{\top},$
and, for two vectors ${\bm{v}},\bm{u}$: ${\left\|\bm{u}{\bm{v}}^{\top}\right\|_{\rm{sum}}}=\|\bm{u}\|_{1}\|{\bm{v}}\|_{1}$. Combining these and noting that
$\|\bm{e}_{y_{i}}-\mathbb{S}_{y_{i}}({\bm{W}}\bm{h}_{i})\|_{1}=2(1-s_{y_{i}})$
together with using the assumption $\|\bm{h}_{i}\|\leq B$ yields the advertised upper bound. ∎
Built upon B.1, we obtain a simple bound on the difference between losses at two different points using the max-norm on the iterates.
Lemma B.2.
For any ${\bm{W}},{\bm{W}}_{0}\in\mathbb{R}^{k\times d}$, suppose that $\mathcal{L}({\bm{W}})$ is convex, we have
$|\mathcal{L}({\bm{W}})-\mathcal{L}({\bm{W}}_{0})|\leq 2B{\left\|{\bm{W}}-{\bm{W}}_{0}\right\|_{\max}}.$
Proof.
By convexity of $\mathcal{L}$, we have
$\mathcal{L}({\bm{W}}_{0})-\mathcal{L}({\bm{W}})\leq\langle\nabla\mathcal{L}({\bm{W}}_{0}),{\bm{W}}_{0}-{\bm{W}}\rangle\leq{\left\|\nabla\mathcal{L}({\bm{W}}_{0})\right\|_{\rm{sum}}}{\left\|{\bm{W}}_{0}-{\bm{W}}\right\|_{\max}}\leq 2B{\left\|{\bm{W}}_{0}-{\bm{W}}\right\|_{\max}}\,,$
where the last inequality is by Lemma B.1. Similarly, we can also show that $\mathcal{L}({\bm{W}})-\mathcal{L}({\bm{W}}_{0})\leq 2B{\left\|{\bm{W}}_{0}-{\bm{W}}\right\|_{\max}}$. ∎
Lemma B.3 shows the close relationship between ${\mathcal{G}}({\bm{W}})$ and $\mathcal{L}({\bm{W}})$. ${\mathcal{G}}({\bm{W}})$ not only lower bounds $\mathcal{L}({\bm{W}})$, but also upper bounds $\mathcal{L}({\bm{W}})$ up to a constant provided that the loss is sufficiently small. Moreover, the rate of convergence $\frac{{\mathcal{G}}({\bm{W}})}{\mathcal{L}({\bm{W}})}$ depends on the rate of decrease in the loss.
Lemma B.3 (${\mathcal{G}}({\bm{W}})$ as proxy to the loss).
Let ${\bm{W}}\in\mathbb{R}^{k\times d}$, we have
  1. (i)
    $1\geq\frac{{\mathcal{G}}({\bm{W}})}{\mathcal{L}({\bm{W}})}\geq 1-\frac{n\mathcal{L}({\bm{W}})}{2}$
  2. (ii)
    Suppose that ${\bm{W}}$ satisfies $\mathcal{L}({\bm{W}})\leq\frac{\log 2}{n}$ or ${\mathcal{G}}({\bm{W}})\leq\frac{1}{2n}$, then $\mathcal{L}({\bm{W}})\leq 2{\mathcal{G}}({\bm{W}}).$
Proof.
(i) Denote for simplicity $s_{i}:=s_{i{y_{i}}}=\mathbb{S}_{{y_{i}}}({\bm{W}}\bm{h}_{i})$, thus $\mathcal{L}({\bm{W}})=\frac{1}{n}\sum_{i\in[n]}\log(1/s_{i})$ and ${\mathcal{G}}(W)=\frac{1}{n}\sum_{i\in[n]}(1-s_{i})$. For the upper bound, simply use the fact that $e^{x-1}\geq x,$ forall $x\in[0,1],$ thus $\log(1/s_{i})\geq 1-s_{i}$ for all $i\in[n]$.
The lower bound can be proved using the exact same arguments in the proof of Zhang et al. (2024, Lemma C.7) for the binary case. For completeness, we provide an alternative elementary proof. It suffices to prove for $n=1$ that for $s\in(0,1)$:
$1-s\geq\log(1/s)-\frac{1}{2}\log^{2}(1/s).$ (15)
The general case follows by summing over $s=s_{i}$ and using $\sum_{i\in[n]}\log^{2}(1/s_{i})\leq\left(\sum_{i\in[n]}\log(1/s_{i})\right)^{2}$ since $\log(1/s_{i})>0$. For (15), let $x=\log(1/s)>0$. The inequality becomes $e^{-x}\leq 1-x+x^{2}/2$, which holds for $x>0$ by the second-order Taylor expansion of $e^{-x}$ around 00.
(ii) The sufficiency of $\mathcal{L}({\bm{W}})\leq\frac{\log 2}{n}$ (to guarantee that $\mathcal{L}({\bm{W}})\leq 2{\mathcal{G}}({\bm{W}})$) follows from (i) and $\mathcal{L}({\bm{W}})\leq\frac{\log 2}{n}\leq\frac{1}{n}$. The inequality $\log(\frac{1}{x})\leq 2(1-x)$ holds when $x\in[0.2032,1]$. This translates to the following sufficient condition on $s_{iy_{i}}$
$s_{i}=\frac{e^{\bm{\ell}_{i}[y_{i}]}}{\sum_{c\in[k]}e^{\bm{\ell}_{i}[c]}}=\frac{1}{1+\sum_{c\in[k],c\neq y_{i}}e^{\bm{\ell}_{i}[c]-\bm{\ell}_{i}[y_{i}]}}\geq 0.2032.$
Under the assumption ${\mathcal{G}}({\bm{W}})\leq\frac{1}{2n}$, we have $1-s_{i}\leq\sum_{i\in[n]}(1-s_{i})=n{\mathcal{G}}({\bm{W}})\leq\frac{1}{2}$, from which we obtain $s_{i}\geq\frac{1}{2}\geq 0.2032$ for all $i\in[n]$. ∎
Lemma B.4 shows that weights ${\bm{W}}$ of low loss separate the data. It is used in deriving the lower bound on the unnormalized margin.
Lemma B.4 (Low $\mathcal{L}({\bm{W}})$ implies separability).
Suppose that there exists ${\bm{W}}\in\mathbb{R}^{k\times d}$ such that $\mathcal{L}({\bm{W}})\leq\frac{\log 2}{n}$, then we have
$(\bm{e}_{y_{i}}-\bm{e}_{c})^{T}{\bm{W}}\bm{h}_{i}\geq 0,\quad\text{for all $i\in[n]$ and for all $c\in[k]$ such that $c\neq y_{i}$}.$ (16)
Proof.
We rewrite the loss into the form:
$\mathcal{L}({\bm{W}})=-\frac{1}{n}\sum_{i\in[n]}\log(\frac{e^{\bm{\ell}_{i}[y_{i}]}}{\sum_{c\in[k]}e^{\bm{\ell}_{i}[c]}})=\frac{1}{n}\sum_{i\in[n]}\log(1+\sum_{c\neq y_{i}}e^{-(\bm{\ell}_{i}[y_{i}]-\bm{\ell}_{i}[c])}).$
Fix any $i\in[n]$, by the assumption that $\mathcal{L}({\bm{W}})\leq\frac{\log 2}{n}$, we have the following:
$\log(1+\sum_{c\neq y_{i}}e^{-(\bm{\ell}_{i}[y_{i}]-\bm{\ell}_{i}[c])})\leq n\mathcal{L}({\bm{W}})\leq\log(2).$
This implies:
$e^{-\min_{c\neq y_{i}}(\bm{\ell}_{i}[y_{i}]-\bm{\ell}_{i}[c])}=\max_{c\neq y_{i}}e^{-(\bm{\ell}_{i}[y_{i}]-\bm{\ell}_{i}[c])\leq}\leq\sum_{c\neq y_{i}}e^{-(\bm{\ell}_{i}[y_{i}]-\bm{\ell}_{i}[c])}\leq 1.$
After taking $\log$ on both sides, we obtain the following: $\bm{\ell}_{i}[y_{i}]-\bm{\ell}_{i}[c]=(\bm{e}_{y_{i}}-\bm{e}_{c})^{T}{\bm{W}}\bm{h}_{i}\geq 0$ for any $c\in[k]$ such that $c\neq y_{i}$. ∎
Lemma B.5 shows that the ratio of ${\mathcal{G}}({\bm{W}})$ at two points can be bounded by the exponential of the max-norm of their differences. It is used in handling the second order term in the Taylor expansion of the loss.
Lemma B.5 (Ratio of ${\mathcal{G}}({\bm{W}})$).
For any $\psi\in[0,1]$, we have the following:
$\frac{{\mathcal{G}}({\bm{W}}+\psi\triangle{\bm{W}})}{{\mathcal{G}}({\bm{W}})}\leq e^{2B\psi{\left\|\triangle{\bm{W}}\right\|_{\max}}}$
Proof.
By the definition of ${\mathcal{G}}({\bm{W}})$, we have:
$\frac{{\mathcal{G}}({\bm{W}}+\psi\triangle{\bm{W}})}{{\mathcal{G}}({\bm{W}})}=\frac{\sum_{i\in[n]}\bigl{(}1-\mathbb{S}_{y_{i}}(({\bm{W}}+\psi\triangle{\bm{W}})\bm{h}_{i})\bigr{)}}{\sum_{i\in[n]}\bigl{(}1-\mathbb{S}_{y_{i}}({\bm{W}}\bm{h}_{i})\bigr{)}}.$
For any $c\in[k]$ and ${\bm{v}},{\bm{v}}^{\prime}\in\mathbb{R}^{k}$, we have:
$\frac{1-\mathbb{S}_{c}({\bm{v}}^{\prime})}{1-\mathbb{S}_{c}({\bm{v}})}$ $=\frac{1-\frac{e^{v^{\prime}_{c}}}{\sum_{i\in[k]}e^{v^{\prime}_{i}}}}{1-\frac{e^{v_{c}}}{\sum_{i\in[k]}e^{v_{i}}}}$
$=\frac{\frac{\sum_{j\in[k],j\neq c}e^{v^{\prime}_{j}}}{\sum_{i\in[k]}e^{v^{\prime}_{i}}}}{\frac{\sum_{j\in[k],j\neq c}e^{v_{j}}}{\sum_{i\in[k]}e^{v_{i}}}}$
$=\frac{\sum_{j\in[k],j\neq c}\sum_{i\in[k]}e^{v^{\prime}_{j}+v_{i}}}{\sum_{j\in[k],j\neq c}\sum_{i\in[k]}e^{v_{j}+v^{\prime}_{i}}}$
$\leq e^{2\lVert{\bm{v}}-{\bm{v}}^{\prime}\rVert_{\infty}}.$
The last inequality is because $\frac{e^{v^{\prime}_{j}+v_{i}}}{e^{v_{j}+v^{\prime}_{i}}}\leq e^{|v^{\prime}_{j}-v_{j}|+|v_{i}-v^{\prime}_{i}|}\leq e^{2\lVert{\bm{v}}-{\bm{v}}^{\prime}\rVert_{\infty}}$, which implies that $\sum_{j\in[k],j\neq c}\sum_{i\in[k]}e^{v^{\prime}_{j}+v_{i}}\leq e^{2\lVert{\bm{v}}-{\bm{v}}^{\prime}\rVert_{\infty}}\sum_{j\in[k],j\neq c}\sum_{i\in[k]}e^{v_{j}+v^{\prime}_{i}}$. Next, we specialize this result to ${\bm{v}}^{\prime}=({\bm{W}}+\psi\triangle{\bm{W}})\bm{h}_{i}$, ${\bm{v}}={\bm{W}}\bm{h}_{i}$, and $c=y_{i}$ for any $i\in[n]$ to obtain:
$\frac{1-\mathbb{S}_{y_{i}}(({\bm{W}}+\psi\triangle{\bm{W}})\bm{h}_{i})\bigr{)}}{1-\mathbb{S}_{y_{i}}({\bm{W}}\bm{h}_{i})}\leq e^{2\psi\lVert\triangle{\bm{W}}\bm{h}_{i}\rVert_{\infty}}\leq e^{2B\psi{\left\|\triangle{\bm{W}}\right\|_{\max}}}.$
Then, we rearrange and sum over $i\in[n]$ to obtain: $\sum_{i\in[n]}\bigl{(}1-\mathbb{S}_{y_{i}}(({\bm{W}}+\psi\triangle{\bm{W}})\bm{h}_{i})\bigr{)}\leq e^{2B\psi{\left\|\triangle{\bm{W}}\right\|_{\max}}}\sum_{i\in[n]}\bigl{(}1-\mathbb{S}_{y_{i}}({\bm{W}}\bm{h}_{i})\bigr{)}$, from which the desired inequality follows. ∎
Proof Overview
We consider a decay learning rate schedule of the form $\eta_{t}=\Theta(\frac{1}{t^{a}})$ where $a\in(0,1]$. The first step is to show that the loss monotonically decreases after certain time and the rate depends on ${\mathcal{G}}({\bm{W}})$. To obtain this, we apply Lemma B.1 and Lemma A.3 to upper bound the first-order and second-order terms in the Taylor expansion of the loss 17, respectively. The difference between SignGD and Adam is that Adam involves an additional term $\bigm{|}\langle\nabla\mathcal{L}({\bm{W}}_{t}),\frac{\bm{M}_{t}}{\sqrt{{\bm{V}}_{t}}}-\frac{\nabla\mathcal{L}({\bm{W}}_{t})}{|\nabla\mathcal{L}({\bm{W}}_{t})|}\rangle\bigm{|}$. To handle it, we apply Lemma D.2 and Lemma D.3 to bound the first ($\bm{M}_{t}$) and second moment (${\bm{V}}_{t}$) buffer of Adam using ${\mathcal{G}}({\bm{W}})$. Next, we use loss monotonically to derive a lower bound on the unnormalized margin which involves the ratio $\frac{{\mathcal{G}}({\bm{W}})}{\mathcal{L}({\bm{W}})}$. A crucial step involved is to find a time $\bar{t}_{2}$ such that separability (36) holds for all $t\geq\bar{t}_{2}$, and the existence of $\bar{t}_{2}$ is guaranteed by loss monotonicity given $\mathcal{L}({\bm{W}}_{t})\leq\frac{\log 2}{n}$ implying separability (36) proved in Lemma B.4.
Then, we argue that the ratio $\frac{{\mathcal{G}}({\bm{W}}_{t})}{\mathcal{L}({\bm{W}}_{t})}$ converges to $1$ exponentially fast (recalling that $1\geq\frac{{\mathcal{G}}({\bm{W}}_{t})}{\mathcal{L}({\bm{W}}_{t})}\geq 1-\frac{n\mathcal{L}({\bm{W}}_{t})}{2}$) by proving the same rate for the decrease of the loss $\mathcal{L}({\bm{W}}_{t})$. We first choose a time $t_{1}$ after $t_{0}$ (recall that $t_{0}$ is the time that satisfies Assumption 3.4) such that $\mathcal{L}({\bm{W}}_{t+1})\leq\mathcal{L}({\bm{W}}_{t})-\frac{\eta_{t}\gamma}{2}{\mathcal{G}}({\bm{W}}_{T})$ for all $t\geq t_{1}$. Next, we lower bound $G({\bm{W}}_{t})$ using $\mathcal{L}({\bm{W}}_{t})$. By Lemma B.3, there are two sufficient conditions (namely, $\mathcal{L}({\bm{W}}_{t})\leq\frac{\log 2}{n}\eqqcolon\tilde{\mathcal{L}}$ or $\bm{G}({\bm{W}}_{t})\leq\frac{1}{2n}$) that guarantee $\mathcal{L}({\bm{W}}_{t})\leq 2{\mathcal{G}}({\bm{W}}_{t})$. We choose a time $t_{2}$ (after $t_{1}$) that is sufficiently large such that there exists $t^{*}\in[t_{1},t_{2}]$ for which we have ${\mathcal{G}}({\bm{W}}_{t^{*}})\leq\frac{\tilde{\mathcal{L}}}{2}\leq\frac{1}{2n}$. This not only guarantees that $\mathcal{L}({\bm{W}}_{t^{*}})\leq 2{\mathcal{G}}({\bm{W}}_{t^{*}})$ at time $t^{*}$, but also (crucially due to monotonicity) implies that $\mathcal{L}({\bm{W}}_{t})\leq\mathcal{L}({\bm{W}}_{t^{*}})\leq 2{\mathcal{G}}({\bm{W}}_{t^{*}})\leq\frac{\log 2}{n}$ for all $t\geq t_{2}$. Thus, we observe that the other sufficient condition $\mathcal{L}({\bm{W}}_{t})\leq\frac{\log 2}{n}$ is satisfied, from which we conclude that $\mathcal{L}({\bm{W}}_{t})\leq 2{\mathcal{G}}({\bm{W}}_{t})$ for all $t\geq t_{2}$. We remark that the choice of $t_{2}$ depends on $\mathcal{L}({\bm{W}}_{t_{1}})$ (whose magnitude is bounded using Lemma B.2), and $t_{2}$ can be used as $\bar{t}_{2}$ above. To recap, $t_{1}$ is the time (after $t_{0}$) after which the successive loss decrease is lower bounded by the product $\eta_{t}\gamma{\mathcal{G}}({\bm{W}}_{t})$; $t_{2}$ (after $t_{1}$) is the time after which $\mathcal{L}({\bm{W}}_{t})\leq\frac{\log 2}{n}$ (thus, both $\mathcal{L}({\bm{W}}_{t})\leq 2{\mathcal{G}}({\bm{W}}_{t})$ and separability condition (36) hold for all $t\geq t_{2}$).
The next step is to bound the max-norm of the iterates in terms of learning rates. The SignGD case is straightforward given ${\left\|\texttt{sign}(\nabla\mathcal{L}({\bm{W}}))\right\|_{\max}}=1$. In the case of Adam, the important step is to show that ${\bm{V}}_{t}[c,j]\leq\mathcal{O}(\mathcal{L}({\bm{W}}_{t})^{2})$ for all $c\in[k]$ and $j\in[d]$. Thus, ${\bm{V}}_{t}[c,j]\leq 1$ is feasible again due to loss monotonicity. To prove this, we use Lemma B.1 that lower bounds ${\left\|\nabla\mathcal{L}({\bm{W}})\right\|_{\max}}$ using ${\mathcal{G}}({\bm{W}})$ up to some constant. Here, we provide a summary of the lemmas (in sections A and B) on the properties of $\mathcal{L}({\bm{W}})$ and ${\mathcal{G}}({\bm{W}})$ that are used in proving the various key steps in the Proof Overview:
  • Loss Monotonicity: Lemma A.2 is on the Taylor expansion of the loss; Lemma A.3 and B.5 are for handling the second-order term in the expansion; Lemma B.1 proves that ${\left\|\nabla\mathcal{L}({\bm{W}})\right\|_{\rm{sum}}}\geq\gamma{\mathcal{G}}({\bm{W}})$.
  • Unnormalized Margin: Lemma B.4 is on the separability condition (36) implied by a low loss ($\mathcal{L}({\bm{W}})\leq\frac{\log 2}{n}$); Lemma B.3 proves that $\frac{{\mathcal{G}}({\bm{W}})}{\mathcal{L}({\bm{W}})}\leq 1$.
  • Convergence of $\frac{{\mathcal{G}}({\bm{W}}_{t})}{\mathcal{L}({\bm{W}}_{t})}$: Lemma B.3 proves that $\frac{{\mathcal{G}}({\bm{W}}_{t})}{\mathcal{L}({\bm{W}}_{t})}\geq 1-\frac{n\mathcal{L}({\bm{W}}_{t})}{2}$, and it provides the sufficient conditions that gaurantee $\mathcal{L}({\bm{W}})\leq 2{\mathcal{G}}({\bm{W}})$; Lemma B.2 provides a simple bound on the loss using the max-norm of the iterates.
  • Iterate Bound: Lemma B.1 proves that $2B\cdot{\mathcal{G}}({\bm{W}})\leq{\left\|\nabla\mathcal{L}({\bm{W}})\right\|_{\max}}$.
Appendix C Proof of SignGD
In this section, we break the proof of implicit bias of SignGD into several parts following the arguments in the Proof Overview. Lemma C.1 shows the descent properties of SignGD. It is used in Lemma C.2 to lower bound the unnormalized margin, and in the proof of Theorem C.4 to show the convergence of $\frac{{\mathcal{G}}({\bm{W}}_{t})}{\mathcal{L}({\bm{W}}_{t})}$.
Lemma C.1 (SignGD Descent).
Under the same setting as Theorem C.4, it holds for all $t\geq 0$,
$\mathcal{L}({\bm{W}}_{t+1})$ $\leq\mathcal{L}({\bm{W}}_{t})-\gamma\eta_{t}(1-\alpha_{s_{1}}\eta_{t}){\mathcal{G}}({\bm{W}}_{t}),$
where $\alpha_{s_{1}}$ is some constant that depends on $B$ and $\gamma$.
Proof.
By Lemma A.2, we let ${\bm{W}}^{\prime}={\bm{W}}_{t+1}$, ${\bm{W}}={\bm{W}}_{t}$, $\bm{\Delta}_{t}={\bm{W}}_{t+1}-{\bm{W}}_{t}$, and define ${\bm{W}}_{t,t+1,\zeta}:={\bm{W}}_{t}+\zeta({\bm{W}}_{t+1}-{\bm{W}}_{t})$. We choose $\zeta^{*}$ such that ${\bm{W}}_{t,t+1,\zeta^{*}}$ satisfies (14), we have:
$\mathcal{L}({\bm{W}}_{t+1})$ $=\mathcal{L}({\bm{W}}_{t})+\underbrace{\langle\nabla\mathcal{L}({\bm{W}}_{t}),\bm{\Delta}_{t}\rangle}_{\spadesuit_{t}}$
$\quad+\frac{1}{2n}\sum_{i\in[n]}\underbrace{\bm{h}_{i}^{\top}\bm{\Delta}_{t}^{\top}\left(\operatorname{diag}(\mathbb{S}({\bm{W}}_{t,t+1,\gamma}\bm{h}_{i}))-\mathbb{S}({\bm{W}}_{t,t+1,\zeta^{*}}\bm{h}_{i})\mathbb{S}({\bm{W}}_{t,t+1,\zeta^{*}}\bm{h}_{i})^{\top}\right)\bm{\Delta}_{t}\,\bm{h}_{i}}_{\clubsuit_{t}}\,.$ (17)
For the $\spadesuit_{t}$ term, we have by Lemma B.1:
$\spadesuit_{t}=\langle\nabla\mathcal{L}({\bm{W}}_{t}),-\eta_{t}\frac{\nabla\mathcal{L}({\bm{W}}_{t})}{\sqrt{\nabla\mathcal{L}^{2}({\bm{W}}_{t})}}\rangle=-\eta_{t}{\left\|\nabla\mathcal{L}({\bm{W}}_{t})\right\|_{\rm{sum}}}\leq-\eta_{t}\gamma G({\bm{W}}_{t})$
For the $\clubsuit_{t}$ term, we let ${\bm{v}}=\bm{\Delta}_{t}\bm{h}_{i}$ and $\bm{s}=\mathbb{S}({\bm{W}}_{t,t+1,\zeta^{*}}\bm{h}_{i})$, and apply Lemma A.3 to obtain
$\clubsuit_{t}\leq 4\|\bm{\Delta}_{t}\bm{h}_{i}\|_{\infty}^{2}(1-\mathbb{S}_{y_{i}}({\bm{W}}_{t,t+1,\zeta^{*}}\bm{h}_{i}))\leq 4\eta_{t}^{2}B^{2}(1-\mathbb{S}_{y_{i}}({\bm{W}}_{t,t+1,\zeta^{*}}\bm{h}_{i})),$
where in the second inequality we have used $\|\bm{\Delta}_{t}\bm{h}_{i}\|_{\infty}\leq{\left\|\bm{\Delta}_{t}\right\|_{\max}}\|\bm{h}_{i}\|_{1}$, $\|\bm{h}_{i}\|_{1}\leq B$, and ${\left\|\bm{\Delta}_{t}\right\|_{\max}}=\eta_{t}{\left\|\texttt{sign}(\nabla\mathcal{L}({\bm{W}}_{t}))\right\|_{\max}}\leq\eta_{t}$. Putting these two pieces together, we obtain
$\mathcal{L}({\bm{W}}_{t+1})$ $\leq\mathcal{L}({\bm{W}}_{t})-\gamma\eta_{t}{\mathcal{G}}({\bm{W}}_{t})+2\eta_{t}^{2}B^{2}\frac{1}{n}\sum_{i\in[n]}(1-\mathbb{S}_{y_{i}}({\bm{W}}_{t,t+1,\zeta^{*}}\bm{h}_{i}))$
$=\mathcal{L}({\bm{W}}_{t})-\gamma\eta_{t}{\mathcal{G}}({\bm{W}}_{t})+2\eta_{t}^{2}B^{2}{\mathcal{G}}({\bm{W}}_{t,t+1,\zeta^{*}})$
$\leq\mathcal{L}({\bm{W}}_{t})-\gamma\eta_{t}{\mathcal{G}}({\bm{W}}_{t})+2\eta_{t}^{2}B^{2}\sup_{\zeta\in[0,1]}{\mathcal{G}}({\bm{W}}_{t,t+1,\zeta})$
$=\mathcal{L}({\bm{W}}_{t})-\gamma\eta_{t}{\mathcal{G}}({\bm{W}}_{t})+2\eta_{t}^{2}B^{2}{\mathcal{G}}({\bm{W}}_{t})\sup_{\zeta\in[0,1]}\frac{{\mathcal{G}}({\bm{W}}_{t}+\zeta\bm{\Delta}_{t})}{{\mathcal{G}}({\bm{W}}_{t})}$
$\stackrel{{\scriptstyle(a)}}{{\leq}}\mathcal{L}({\bm{W}}_{t})-\gamma\eta_{t}{\mathcal{G}}({\bm{W}}_{t})+2\eta_{t}^{2}B^{2}{\mathcal{G}}({\bm{W}}_{t})\sup_{\zeta\in[0,1]}e^{2B\zeta{\left\|\bm{\Delta}_{t}\right\|_{\max}}}$
$\stackrel{{\scriptstyle(b)}}{{\leq}}\mathcal{L}({\bm{W}}_{t})-\gamma\eta_{t}{\mathcal{G}}({\bm{W}}_{t})+2\eta_{t}^{2}B^{2}e^{2B\eta_{0}}{\mathcal{G}}({\bm{W}}_{t}),$ (18)
where (a) is by Lemma B.5 and (b) is by ${\left\|\bm{\Delta}\right\|_{\max}}\leq\eta_{t}$. Letting $\alpha_{s_{1}}=\frac{2B^{2}e^{2B\eta_{0}}}{\gamma}$, Eq. (18) simplifies to:
$\mathcal{L}({\bm{W}}_{t+1})$ $\leq\mathcal{L}({\bm{W}}_{t})-\gamma\eta_{t}(1-\alpha_{s_{1}}\eta_{t}){\mathcal{G}}({\bm{W}}_{t}),$
from which we observe that the loss starts to monotonically decrease after $\eta_{t}$ satisfies $\eta_{t}\leq\frac{1}{\alpha_{s_{1}}}$ for a decreasing learning rate schedule. ∎
For a decaying learning rate schedule, Lemma C.1 implies that the loss monotonically decreases after a certain time. Thus, we know that the assumption of Lemma C.2 can be satisfied. In the proof of Theorem C.4, we will specify a concrete form of $\tilde{t}$ in Lemma C.2.
Lemma C.2 (SignGD Unnormalized Margin).
Suppose that there exist $\tilde{t}$ such that $\mathcal{L}({\bm{W}}_{t})\leq\frac{\log 2}{n}$ for all $t>\tilde{t}$, then we have
$\min_{i\in[n],c\neq y_{i}}(\bm{e}_{y_{i}}-\bm{e}_{c})^{T}{\bm{W}}_{t}\bm{h}_{i}\geq\gamma\sum_{s=\tilde{t}}^{t-1}\eta_{s}\frac{{\mathcal{G}}({\bm{W}}_{s})}{\mathcal{L}({\bm{W}}_{s})}-\alpha_{s_{2}}\sum_{s=\tilde{t}}^{t-1}\eta_{s}^{2},$
where $\alpha_{s_{2}}$ is some constant that depends on $B$.
Proof.
We let $\alpha_{s_{2}}=2Be^{2B\eta_{0}}$, then from (18), we have for $t>\tilde{t}$:
$\mathcal{L}({\bm{W}}_{t+1})$ $\leq\mathcal{L}({\bm{W}}_{t})-\gamma\eta_{t}{\mathcal{G}}({\bm{W}}_{t})+\alpha_{s_{2}}\eta_{t}^{2}{\mathcal{G}}({\bm{W}}_{t})$
$=\mathcal{L}({\bm{W}}_{t})\bigl{(}1-\gamma\eta_{t}\frac{{\mathcal{G}}({\bm{W}}_{t})}{\mathcal{L}({\bm{W}}_{t})}+\alpha_{s_{2}}\eta_{t}^{2}\frac{{\mathcal{G}}({\bm{W}}_{t})}{\mathcal{L}({\bm{W}}_{t})}\bigr{)}$
$\leq\mathcal{L}({\bm{W}}_{t})\exp\bigl{(}-\gamma\eta_{t}\frac{{\mathcal{G}}({\bm{W}}_{t})}{\mathcal{L}({\bm{W}}_{t})}+\alpha_{s_{2}}\eta_{t}^{2}\frac{{\mathcal{G}}({\bm{W}}_{t})}{\mathcal{L}({\bm{W}}_{t})}\bigr{)}$
$\leq\mathcal{L}({\bm{W}}_{\tilde{t}})\exp\bigl{(}-\gamma\sum_{s=\tilde{t}}^{t}\eta_{s}\frac{{\mathcal{G}}({\bm{W}}_{s})}{\mathcal{L}({\bm{W}}_{s})}+\alpha_{s_{2}}\sum_{s=\tilde{t}}^{t}\eta_{s}^{2}\bigr{)}.$
$\leq\frac{\log 2}{n}\exp\bigl{(}-\gamma\sum_{s=\tilde{t}}^{t}\eta_{s}\frac{{\mathcal{G}}({\bm{W}}_{s})}{\mathcal{L}({\bm{W}}_{s})}+\alpha_{s_{2}}\sum_{s=\tilde{t}}^{t}\eta_{s}^{2}\bigr{)},$ (19)
where the penultimate inequality uses Lemma B.3, and the last inequality uses the assumption that $\mathcal{L}({\bm{W}}_{t})\leq\frac{\log 2}{n}$ for all $t\geq\tilde{t}$. Then, we have for all $t>\tilde{t}$:
$e^{-\min_{i\in[n],c\neq y_{i}}(\bm{e}_{y_{i}}-\bm{e}_{c})^{T}{\bm{W}}_{t}\bm{h}_{i}}$ $=\max_{i\in[n]}e^{-\min_{c\neq y_{i}}(\bm{e}_{y_{i}}-\bm{e}_{c})^{T}{\bm{W}}_{t}\bm{h}_{i}}$
$\stackrel{{\scriptstyle(a)}}{{\leq}}\max_{i\in[n]}\frac{1}{\log 2}\log\bigl{(}1+e^{-\min_{c\neq y_{i}}(\bm{e}_{y_{i}}-\bm{e}_{c})^{T}{\bm{W}}_{t}\bm{h}_{i}}\bigr{)}$
$\leq\max_{i\in[n]}\frac{1}{\log 2}\log(1+\sum_{c\neq y_{i}}e^{-(\bm{e}_{y_{i}}-\bm{e}_{c})^{T}{\bm{W}}_{t}\bm{h}_{i}})\leq\frac{n\mathcal{L}({\bm{W}}_{t})}{\log 2}$
$\stackrel{{\scriptstyle(b)}}{{\leq}}\exp\bigl{(}-\gamma\sum_{s=\tilde{t}}^{t-1}\eta_{s}\frac{{\mathcal{G}}({\bm{W}}_{s})}{\mathcal{L}({\bm{W}}_{s})}+\alpha_{s_{2}}\sum_{s=\tilde{t}}^{t-1}\eta_{s}^{2}\bigr{)}.$
(a) is by the following: the assumption $\mathcal{L}({\bm{W}}_{t})\leq\frac{\log 2}{n}$ implies that $\min_{c\neq y_{i}}(\bm{e}_{y_{i}}-\bm{e}_{c})^{T}{\bm{W}}_{t}\bm{h}_{i}\geq 0$ for all $i\in[n]$ by Lemma B.4. We also know the inequality $\frac{\log(1+e^{-z})}{e^{-z}}\geq\log 2$ holds for any $z\geq 0$. Then, for any $i\in[n]$, we can set $z=\min_{c\neq y_{i}}(\bm{e}_{y_{i}}-\bm{e}_{c})^{T}{\bm{W}}_{t}\bm{h}_{i}$ to obtain the desired inequality; and (b) is by (19). Finally, taking $\log$ on both sides leads to the result. ∎
Next Lemma upper bounds the max-norm of SignGD iterates using learning rates. It is used in the proof of Theorem C.4.
Lemma C.3 (SignGD ${\left\|{\bm{W}}_{t}\right\|_{\max}}$).
For SignGD, we have for any $t>0$ that
${\left\|{\bm{W}}_{t}\right\|_{\max}}\leq{\left\|{\bm{W}}_{0}\right\|_{\max}}+\sum_{s=0}^{t-1}\eta_{s}.$
Proof.
By the SignGD update rule (4), we have
${\bm{W}}_{t+1}={\bm{W}}_{0}-\sum_{s=0}^{t}\eta_{s}\texttt{sign}(\nabla\mathcal{L}({\bm{W}}_{t})).$
This leads to ${\left\|{\bm{W}}_{t}\right\|_{\max}}\leq{\left\|{\bm{W}}_{0}\right\|_{\max}}+\sum_{s=0}^{t-1}\eta_{s}$ ∎
The main step in the proof of Theorem C.4 is to determine the time that satisfies the assumption in Lemma C.2 and show the convergence of $\frac{{\mathcal{G}}({\bm{W}}_{t})}{\mathcal{L}({\bm{W}}_{t})}$. After this, Lemma C.2 and Lemma C.3 will be combined to obtain the final result.
Theorem C.4.
Suppose that Assumption 3.1, 3.3, and 3.5 hold, then there exists $t_{s_{2}}=t_{s_{2}}(n,\gamma,B,{\bm{W}}_{0})$ such that SignGD achieves the following for all $t>t_{s_{2}}$
$\left|\frac{\min_{i\in[n],c\neq y_{i}}(\bm{e}_{y_{i}}-\bm{e}_{c})^{T}{\bm{W}}_{t}\bm{h}_{i}}{{\left\|{\bm{W}}_{t}\right\|_{\max}}}-\gamma\right|$ $\leq\mathcal{O}\Bigg{(}\frac{\sum_{s=t_{s_{2}}}^{t-1}\eta_{s}e^{-\frac{\gamma}{4}\sum_{\tau=t_{s_{2}}}^{s-1}\eta_{\tau}}+\sum_{s=0}^{t_{s_{2}}-1}\eta_{s}+\sum_{s=t_{s_{2}}}^{t-1}\eta_{s}^{2}}{\sum_{s=0}^{t-1}\eta_{s}}\Bigg{)}.$
Proof.
Determination of $t_{s_{1}}$. In Lemma C.1 we choose $t_{s_{1}}$ such that $\eta_{t}\leq\frac{1}{2\alpha_{s_{1}}}$ for all $t\geq t_{s_{1}}$. Considering $\eta_{t}=\Theta(\frac{1}{t^{a}})$ (where $a\in(0,1]$), we set $t_{s_{1}}=(2\alpha_{s_{1}})^{\frac{1}{a}}=(\frac{4B^{2}e^{2B\eta_{0}}}{\gamma})^{\frac{1}{a}}$. Then, we have for all $t\geq t_{s_{1}}$
$\mathcal{L}({\bm{W}}_{t+1})\leq\mathcal{L}({\bm{W}}_{t})-\frac{\eta_{t}\gamma}{2}{\mathcal{G}}({\bm{W}}_{t}).$ (20)
Rearranging this equation and using non-negativity of the loss we obtain $\gamma\sum_{s=t_{s_{1}}}^{t}\eta_{s}{\mathcal{G}}({\bm{W}}_{s})\leq 2\mathcal{L}({\bm{W}}_{t_{s_{1}}})$.
Determination of $t_{s_{2}}$. By Lemma B.2, we can bound $\mathcal{L}({\bm{W}}_{t_{s_{1}}})$ as follows
$|\mathcal{L}({\bm{W}}_{t_{s_{1}}})-\mathcal{L}({\bm{W}}_{0})|\leq 2B{\left\|{\bm{W}}_{t_{s_{1}}}-{\bm{W}}_{0}\right\|_{\max}}\leq 2B\sum_{s=0}^{t_{s_{1}}-1}\eta_{s}{\left\|\texttt{sign}(\nabla\mathcal{L}({\bm{W}}_{s}))\right\|_{\max}}\leq 2B\sum_{s=0}^{t_{s_{1}}-1}\eta_{s},$
where the last inequality is by Lemma D.1. Combining this with the result above and letting ~$\tilde{\mathcal{L}}\coloneqq\frac{\log 2}{n}$, we obtain
${\mathcal{G}}({\bm{W}}_{t^{*}})=\min_{s\in[t_{s_{1}},t_{s_{2}}]}{\mathcal{G}}({\bm{W}}_{s})\leq\frac{2\mathcal{L}({\bm{W}}_{0})+4B\sum_{s=0}^{t_{s_{1}}-1}\eta_{s}}{\gamma\sum_{s=t_{s_{1}}}^{t_{s_{2}}}\eta_{s}}\leq\frac{\tilde{\mathcal{L}}}{2}\leq\frac{1}{2n},$
from which we derive the sufficient condition on $t_{s_{2}}$ to be $\sum_{s=t_{s_{1}}}^{t_{s_{2}}}\eta_{s}\geq\frac{4\mathcal{L}({\bm{W}}_{0})+8B\sum_{s=0}^{t_{s_{1}}-1}\eta_{s}}{\gamma\tilde{\mathcal{L}}}$.
Convergence of $\frac{{\mathcal{G}}({\bm{W}}_{t})}{\mathcal{L}({\bm{W}}_{t})}$ Given ${\mathcal{G}}({\bm{W}}_{t^{*}})\leq\frac{\tilde{\mathcal{L}}}{2}\leq\frac{1}{2n}$, we obtain that $\mathcal{L}({\bm{W}}_{t})\leq\mathcal{L}({\bm{W}}_{t^{*}})\leq 2{\mathcal{G}}({\bm{W}}_{t^{*}})\leq\tilde{L}$ for all $t\geq t_{s_{2}}$, where the first and second inequalities are due to monotonicity in the risk and Lemma B.3, respectively. Thus, the other sufficient condition $\mathcal{L}({\bm{W}}_{t})\leq\frac{\log 2}{n}$ in Lemma B.3 is satisfied, from which we conclude that $\mathcal{L}({\bm{W}}_{t})\leq 2{\mathcal{G}}({\bm{W}}_{t})$ for all $t\geq t_{s_{2}}$. Substituting this into (20), we obtain for all $t>t_{s_{2}}$
$\mathcal{L}({\bm{W}}_{t})\leq(1-\frac{\gamma\eta_{t-1}}{4})\mathcal{L}({\bm{W}}_{t-1})\leq\mathcal{L}({\bm{W}}_{t_{s_{2}}})e^{-\frac{\gamma}{4}\sum_{s=t_{s_{2}}}^{t-1}\eta_{s}}\leq\tilde{\mathcal{L}}e^{-\frac{\gamma}{4}\sum_{s=t_{s_{2}}}^{t-1}\eta_{s}}$
Then, by Lemma B.3, we obtain
$\frac{{\mathcal{G}}({\bm{W}}_{t})}{\mathcal{L}({\bm{W}}_{t})}\geq 1-\frac{n\mathcal{L}({\bm{W}}_{t})}{2}\geq 1-\frac{n\tilde{\mathcal{L}}e^{-\frac{\gamma}{4}\sum_{s=t_{s_{2}}}^{t-1}\eta_{s}}}{2}\geq 1-e^{-\frac{\gamma}{4}\sum_{s=t_{s_{2}}}^{t-1}\eta_{s}}.$ (21)
Margin Convergence Finally, we combine Lemma C.2, Lemma C.3, and (21) to obtain
$|\frac{\min_{i\in[n],c\neq y_{i}}(\bm{e}_{y_{i}}-\bm{e}_{c})^{T}{\bm{W}}_{t}\bm{h}_{i}}{{\left\|{\bm{W}}_{t}\right\|_{\max}}}-\gamma|$ $\leq\frac{\gamma\bigl{(}{\left\|{\bm{W}}_{0}\right\|_{\max}}+\sum_{s=t_{s_{2}}}^{t-1}\eta_{s}e^{-\frac{\gamma}{4}\sum_{\tau=t_{s_{2}}}^{s-1}\eta_{\tau}}+\sum_{s=0}^{t_{s_{2}}-1}\eta_{s}\bigr{)}+\alpha_{s_{2}}\sum_{s=t_{s_{2}}}^{t-1}\eta_{s}^{2}}{{\left\|{\bm{W}}_{0}\right\|_{\max}}+\sum_{s=0}^{t-1}\eta_{s}}$
$\leq\mathcal{O}(\frac{\sum_{s=t_{s_{2}}}^{t-1}\eta_{s}e^{-\frac{\gamma}{4}\sum_{\tau=t_{s_{2}}}^{s-1}\eta_{\tau}}+\sum_{s=0}^{t_{s_{2}}-1}\eta_{s}+\sum_{s=t_{s_{2}}}^{t-1}\eta_{s}^{2}}{\sum_{s=0}^{t-1}\eta_{s}})$
Next, we explicitly upper bound $t_{s_{2}}$ in Theorem C.4 and derive the margin convergence rates of SignGD.
Corollary C.5.
Consider learning rate schedule of the form $\eta_{t}=\Theta(\frac{1}{t^{a}})$ where $a\in(0,1]$, under the same setting as Theorem C.4, then we have for SignGD
$|\frac{\min_{i\in[n],c\neq y_{i}}(\bm{e}_{y_{i}}-\bm{e}_{c})^{T}{\bm{W}}_{t}\bm{h}_{i}}{{\left\|{\bm{W}}_{t}\right\|_{\max}}}-\gamma|=\left\{\begin{array}[]{ll}\mathcal{O}(\frac{t^{1-2a}+n}{t^{1-a}})&\text{if}\quad a<\frac{1}{2}\\ \mathcal{O}(\frac{\log t+n}{t^{1/2}})&\text{if}\quad a=\frac{1}{2}\\ \mathcal{O}(\frac{n}{t^{1-a}})&\text{if}\quad\frac{1}{2}<a<1\\ \mathcal{O}(\frac{n}{\log t})&\text{if}\quad a=1\end{array}\right.$
Proof.
Recall that $t_{s_{1}}=(\frac{4B^{2}e^{2B\eta_{0}}}{\gamma})^{\frac{1}{a}}=:C_{s_{1}}$, and the condition on $t_{s_{2}}$ is $\sum_{s=t_{s_{1}}}^{t_{s_{2}}}\eta_{s}\geq\frac{4\mathcal{L}({\bm{W}}_{0})+8B\sum_{s=0}^{t_{s_{1}}-1}\eta_{s}}{\gamma\tilde{\mathcal{L}}}$, where $\tilde{L}=\frac{\log 2}{n}$. We can apply integral approximations to the terms that involve sums of learning rates to obtain
$t_{s_{2}}\leq C_{s_{2}}n^{\frac{1}{1-a}}t_{s_{1}}+C_{s_{3}}n^{\frac{1}{1-a}}\mathcal{L}({\bm{W}}_{0})^{\frac{1}{1-a}}.$
Given $t_{s_{1}}$ is some constant, this further implies that
$\sum_{s=0}^{t_{s_{2}}-1}\eta_{s}=\mathcal{O}(t_{s_{2}}^{1-a})=\mathcal{O}(n+n\mathcal{L}({\bm{W}}_{0})).$
Next, we focus on the term $\sum_{s=t_{s_{2}}}^{t-1}\eta_{s}^{2}$. For $a>\frac{1}{2}$, this term can be bounded by some constant. For $a<\frac{1}{2}$, we have $\sum_{s=t_{s_{2}}}^{t-1}\eta_{s}^{2}=\mathcal{O}(t^{1-2a})$, and it evaluates to $\mathcal{O}(\log t)$ for $a=\frac{1}{2}$. Finally, we have that $\sum_{s=0}^{t-1}\eta_{s}=\mathcal{O}(t^{1-a})$ for $a<1$ and $\sum_{s=0}^{t-1}\eta_{s}=\mathcal{O}(\log t)$ for $a=1$. The rest arguments can be found in Zhang et al. (2024, Corollary 4.7 and Lemma C.1), including showing the learning rate schedule $\eta_{t}=\frac{1}{(t+2)^{a}}$ satisfying Assumption 3.4, and the term $\sum_{s=t_{s_{2}}}^{t-1}\eta_{s}e^{-\frac{\gamma}{4}\sum_{\tau=t_{s_{2}}}^{s-1}\eta_{\tau}}$ is bounded by some constant for all $a\in(0,1]$. ∎
Appendix D Proof of Adam
The proof of Adam follows the similar approach as SignGD. The key challenge is to connect $\bm{M}_{t}$ and ${\bm{V}}_{t}$ to a per-class decomposition of ${\mathcal{G}}({\bm{W}}_{t})$. The following Lemma in ([)Lemma 6.5]zhang2024implicit is useful. It provides an entry-wise bound on the ratio between the first moment and square root of the second moment.
Lemma D.1.
Considering the Adam updates given in (3a), (3b), and (3c), suppose that $\beta_{1}\leq\beta_{2}$ and set $\alpha=\sqrt{\frac{\beta_{2}(1-\beta_{1})^{2}}{(1-\beta_{2})(\beta_{2}-\beta_{1}^{2})^{2}}}$, then we obtain $\bm{M}_{t}[c,j]\leq\alpha\cdot\sqrt{{\bm{V}}_{t}[c,j]}$ for all $c\in[k]$ and $j\in[d]$.
The following Lemma bounds the first moment buffer ($\bm{M}_{t}$) of Adam in terms of the product of $\eta_{t}$ with ${\mathcal{G}}_{c}({\bm{W}}_{t})$ and $\mathcal{Q}_{c}({\bm{W}}_{t})$. It is used in the proof of Lemma D.4.
Lemma D.2.
Let $c\in[k]$. Under the same setting as Theorem D.8, there exists a time $t_{0}$ such that the following holds for all $t\geq t_{0}$
$|\mathbf{M}_{t}[c,j]-(1-\beta_{1}^{t+1})\nabla\mathcal{L}({\bm{W}}_{t})[c,j]|$ $\leq\alpha_{M}\eta_{t}({\mathcal{G}}_{c}({\bm{W}}_{t})+\mathcal{Q}_{c}({\bm{W}}_{t})),$
where $j\in[d]$ and $\alpha_{M}$ is some constant that depends on $B$ and $\beta_{1}$.
Proof.
For any fixed $c\in[k]$ and $j\in[d]$,
$|\mathbf{M}_{t}[c,j]-(1-\beta_{1}^{t+1})\nabla\mathcal{L}({\bm{W}}_{t})[c,j]|$ $=|\sum_{\tau=0}^{t}(1-\beta_{1})\beta_{1}^{\tau}\bigl{(}\nabla\mathcal{L}({\bm{W}}_{t-\tau})[c,j]-\nabla\mathcal{L}({\bm{W}}_{t})[c,j]\bigr{)}|$
$\leq\sum_{\tau=0}^{t}(1-\beta_{1})\beta_{1}^{\tau}\underbrace{|\nabla\mathcal{L}({\bm{W}}_{t-\tau})[c,j]-\nabla\mathcal{L}({\bm{W}}_{t})[c,j]|}_{\clubsuit}.$ (22)
We first notice that for any ${\bm{W}}\in\mathbb{R}^{k\times d}$, we have $\nabla\mathcal{L}({\bm{W}})[c,j]=\bm{e}_{c}^{T}\nabla\mathcal{L}({\bm{W}})\bm{e}_{j}=-\frac{1}{n}\sum_{i\in[n]}\bm{e}_{c}^{T}\bigl{(}\bm{e}_{y_{i}}-\mathbb{S}({\bm{W}}\bm{h}_{i})\bigr{)}\bm{h}_{i}^{T}\bm{e}_{j}=-\frac{1}{n}\sum_{i\in[n]}\bm{e}_{c}^{T}\bigl{(}\bm{e}_{y_{i}}-\mathbb{S}({\bm{W}}\bm{h}_{i})\bigr{)}h_{ij}$. Then, the gradient difference term becomes
$\clubsuit$ $=|-\frac{1}{n}\sum_{i\in[n]}\bm{e}_{c}^{T}\bigl{(}\bm{e}_{y_{i}}-\mathbb{S}({\bm{W}}_{t-\tau}\bm{h}_{i})\bigr{)}h_{ij}+\frac{1}{n}\sum_{i\in[n]}\bm{e}_{c}^{T}\bigl{(}\bm{e}_{y_{i}}-\mathbb{S}({\bm{W}}_{t}\bm{h}_{i})\bigr{)}h_{ij}|$
$=|\frac{1}{n}\sum_{i\in[n]}\bm{e}_{c}^{T}\bigl{(}\mathbb{S}({\bm{W}}_{t-\tau}\bm{h}_{i})-\mathbb{S}({\bm{W}}_{t}\bm{h}_{i})\bigr{)}h_{ij}|$
$=|\frac{1}{n}\sum_{i\in[n]}\bigl{(}\mathbb{S}_{c}({\bm{W}}_{t-\tau}\bm{h}_{i})-\mathbb{S}_{c}({\bm{W}}_{t}\bm{h}_{i})\bigr{)}h_{ij}|$
$\leq B\frac{1}{n}\sum_{i\in[n]}|\mathbb{S}_{c}({\bm{W}}_{t-\tau}\bm{h}_{i})-\mathbb{S}_{c}({\bm{W}}_{t}\bm{h}_{i})|$
$=B\underbrace{\frac{1}{n}\sum_{i\in[n],y_{i}\neq c}|\mathbb{S}_{c}({\bm{W}}_{t-\tau}\bm{h}_{i})-\mathbb{S}_{c}({\bm{W}}_{t}\bm{h}_{i})|}_{\clubsuit_{1}}+B\underbrace{\frac{1}{n}\sum_{i\in[n],y_{i}=c}|\mathbb{S}_{c}({\bm{W}}_{t-\tau}\bm{h}_{i})-\mathbb{S}_{c}({\bm{W}}_{t}\bm{h}_{i})|}_{\clubsuit_{2}}$
Next, we link the $\clubsuit_{1}$ and $\clubsuit_{2}$ terms with ${\mathcal{G}}({\bm{W}})$. Starting with the first term, we obtain:
$\clubsuit_{1}$ $=\frac{1}{n}\sum_{i\in[n],y_{i}\neq c}\mathbb{S}_{c}({\bm{W}}_{t}\bm{h}_{i})|\frac{\mathbb{S}_{c}({\bm{W}}_{t-\tau}\bm{h}_{i})}{\mathbb{S}_{c}({\bm{W}}_{t}\bm{h}_{i})}-1|$
$\stackrel{{\scriptstyle(a)}}{{\leq}}\frac{1}{n}\sum_{i\in[n],y_{i}\neq c}\mathbb{S}_{c}({\bm{W}}_{t}\bm{h}_{i})(e^{2\lVert({\bm{W}}_{t-\tau}-{\bm{W}}_{t})\bm{h}_{i}\rVert_{\infty}}-1)$
$\stackrel{{\scriptstyle(b)}}{{\leq}}\frac{1}{n}\sum_{i\in[n],y_{i}\neq c}\mathbb{S}_{c}({\bm{W}}_{t}\bm{h}_{i})(e^{2B{\left\|{\bm{W}}_{t-\tau}-{\bm{W}}_{t}\right\|_{\max}}}-1)$
$\stackrel{{\scriptstyle(c)}}{{\leq}}\bigl{(}e^{2B\sum_{s=1}^{\tau}\eta_{t-s}{\left\|\frac{\bm{M}_{t-s}}{\sqrt{{\bm{V}}_{t-s}}}\right\|_{\max}}}-1\bigr{)}\bigl{(}\frac{1}{n}\sum_{i\in[n],y_{i}\neq c}\mathbb{S}_{c}({\bm{W}}_{t}\bm{h}_{i})\bigr{)}$
$\stackrel{{\scriptstyle(d)}}{{\leq}}\bigl{(}e^{2\alpha B\sum_{s=1}^{\tau}\eta_{t-s}}-1\bigr{)}\mathcal{Q}_{c}({\bm{W}}_{t}),$
where (a) is by Lemma D.10, (b) is by $\lVert\bm{h}_{i}\rVert_{1}\leq B$ for all $i\in[n]$, (c) is by (3c) and triangle inequality, and $(d)$ is by Lemma D.1 and the definition of ${\mathcal{G}}({\bm{W}}_{t})$. For the second term, we obtain:
$\clubsuit_{2}$ $=\frac{1}{n}\sum_{i\in[n],y_{i}=c}|\mathbb{S}_{c}({\bm{W}}_{t-\tau}\bm{h}_{i})-\mathbb{S}_{c}({\bm{W}}_{t}\bm{h}_{i})|$
$=\frac{1}{n}\sum_{i\in[n],y_{i}=c}|\mathbb{S}_{y_{i}}({\bm{W}}_{t-\tau}\bm{h}_{i})-1+1-\mathbb{S}_{y_{i}}({\bm{W}}_{t}\bm{h}_{i})|$
$=\frac{1}{n}\sum_{i\in[n],y_{i}=c}\bigl{(}1-\mathbb{S}_{y_{i}}({\bm{W}}_{t}\bm{h}_{i})\bigr{)}|\frac{\mathbb{S}_{y_{i}}({\bm{W}}_{t-\tau}\bm{h}_{i})-1}{1-\mathbb{S}_{y_{i}}({\bm{W}}_{t}\bm{h}_{i})}+1|$
$=\frac{1}{n}\sum_{i\in[n],y_{i}=c}\bigl{(}1-\mathbb{S}_{y_{i}}({\bm{W}}_{t}\bm{h}_{i})\bigr{)}|\frac{1-\mathbb{S}_{y_{i}}({\bm{W}}_{t-\tau}\bm{h}_{i})}{1-\mathbb{S}_{y_{i}}({\bm{W}}_{t}\bm{h}_{i})}-1|$
$\stackrel{{\scriptstyle(e)}}{{\leq}}\frac{1}{n}\sum_{i\in[n],y_{i}=c}\bigl{(}1-\mathbb{S}_{y_{i}}({\bm{W}}_{t}\bm{h}_{i})\bigr{)}(e^{2\lVert({\bm{W}}_{t-\tau}-{\bm{W}}_{t})\bm{h}_{i}\rVert_{\infty}}-1)$
$\stackrel{{\scriptstyle(f)}}{{\leq}}\bigl{(}e^{2\alpha B\sum_{s=1}^{\tau}\eta_{t-s}}-1\bigr{)}{\mathcal{G}}_{c}({\bm{W}}_{t}),$
where (e) is by Lemma D.10, and (f) is by the same approach taken for $\clubsuit_{1}$. Based on the upper bounds for $\clubsuit_{1}$ and $\clubsuit_{2}$, we obtain the following: $\clubsuit\leq 2B\bigl{(}e^{2\alpha B\sum_{s=1}^{\tau}\eta_{t-s}}-1\bigr{)}({\mathcal{G}}_{c}({\bm{W}}_{t})+\mathcal{Q}_{c}({\bm{W}}_{t}))$. Then, we substitute this into (22) to obtain:
$|\mathbf{M}_{t}[c,j]-(1-\beta_{1}^{t+1})\nabla\mathcal{L}({\bm{W}}_{t})[c,j]|$ $\leq B(1-\beta_{1})({\mathcal{G}}_{c}({\bm{W}}_{t})+\mathcal{Q}_{c}({\bm{W}}_{t}))\sum_{\tau=0}^{t}\beta_{1}^{\tau}\bigl{(}e^{2\alpha B\sum_{s=1}^{\tau}\eta_{t-s}}-1\bigr{)}$
$\stackrel{{\scriptstyle(g)}}{{\leq}}B(1-\beta_{1})c_{2}\eta_{t}({\mathcal{G}}_{c}({\bm{W}}_{t})+\mathcal{Q}_{c}({\bm{W}}_{t})),$
where (g) is by the Assumption 3.4. ∎
The following Lemma bounds the first moment buffer (${\bm{V}}_{t}$) of Adam in terms of the product of $\eta_{t}$ and with ${\mathcal{G}}_{c}({\bm{W}}_{t})$ and $\mathcal{Q}_{c}({\bm{W}}_{t})$. It is used in the proof of Lemma D.4.
Lemma D.3.
Let $c\in[k]$. Under the same setting as Theorem D.8, there exists a time $t_{0}$ such that the following holds for all $t\geq t_{0}$
$\bigm{|}\sqrt{{\bm{V}}_{t}[c,j]}-\sqrt{(1-\beta_{2}^{t+1})}|\nabla\mathcal{L}({\bm{W}}_{t})[c,j]|\bigm{|}$ $\leq\alpha_{V}\sqrt{\eta_{t}}(\mathcal{Q}_{c}({\bm{W}}_{t})+{\mathcal{G}}_{c}({\bm{W}}_{t})),$
where $j\in[d]$, and $\alpha_{V}$ is some constant that depends on $B$ and $\beta_{2}$.
Proof.
Consider any fixed $c\in[k]$ and $j\in[d]$,
$|{\bm{V}}_{t}[c,j]-(1-\beta_{2}^{t+1})\nabla\mathcal{L}({\bm{W}}_{t})[c,j]^{2}|$ $=|\sum_{\tau=0}^{t}(1-\beta_{2})\beta_{2}^{\tau}\bigl{(}\nabla\mathcal{L}({\bm{W}}_{t-\tau})[c,j]^{2}-\nabla\mathcal{L}({\bm{W}}_{t})[c,j]^{2}\bigr{)}|$
$\leq\sum_{\tau=0}^{t}(1-\beta_{2})\beta_{2}^{\tau}\underbrace{|\nabla\mathcal{L}({\bm{W}}_{t-\tau})[c,j]^{2}-\nabla\mathcal{L}({\bm{W}}_{t})[c,j]^{2}|}_{\spadesuit}.$ (23)
For any ${\bm{W}}\in\mathbb{R}^{k\times d}$, recall that $-\frac{1}{n}\sum_{i\in[n]}\bm{e}_{c}^{T}\bigl{(}\bm{e}_{y_{i}}-\mathbb{S}({\bm{W}}\bm{h}_{i})\bigr{)}h_{ij}$. Then, we can obtain $\nabla\mathcal{L}({\bm{W}})[c,j]^{2}=\frac{1}{n^{2}}\sum_{i\in[n]}\sum_{p\in[n]}h_{ij}h_{pj}(\delta_{cy_{i}}-\mathbb{S}_{c}({\bm{W}}\bm{h}_{i}))(\delta_{cy_{p}}-\mathbb{S}_{c}({\bm{W}}\bm{h}_{p}))$ where $\delta_{cy}=1$ if and only if $c=y$. Next, we define the function $f_{c,i,p}$ to be $f_{c,i,p}({\bm{W}}):=(\delta_{cy_{i}}-\mathbb{S}_{c}({\bm{W}}\bm{h}_{i}))(\delta_{cy_{p}}-\mathbb{S}_{c}({\bm{W}}\bm{h}_{p}))$. Then, we have
$|f_{c,i,p}({\bm{W}}_{t-\tau})-f_{c,i,p}({\bm{W}}_{t})|$ $=\delta_{cy_{i}}\bigl{(}\mathbb{S}_{c}({\bm{W}}_{t}\bm{h}_{p})-\mathbb{S}_{c}({\bm{W}}_{t-\tau}\bm{h}_{p})\bigr{)}+\delta_{cy_{p}}\bigl{(}\mathbb{S}_{c}({\bm{W}}_{t}\bm{h}_{i})-\mathbb{S}_{c}({\bm{W}}_{t-\tau}\bm{h}_{i})\bigr{)}$
$\quad\quad+\bigl{(}\mathbb{S}_{c}({\bm{W}}_{t-\tau}\bm{h}_{i})\mathbb{S}_{c}({\bm{W}}_{t-\tau}\bm{h}_{p})-\mathbb{S}_{c}({\bm{W}}_{t}\bm{h}_{i})\mathbb{S}_{c}({\bm{W}}_{t}\bm{h}_{p})\bigr{)}$
We can substitute this result into $\spadesuit$ to obtain
$\spadesuit$ $=|\frac{1}{n^{2}}\sum_{i\in[n]}\sum_{p\in[n]}h_{ij}h_{pj}(f_{c,i,p}({\bm{W}}_{t-\tau})-f_{c,i,p}({\bm{W}}_{t}))|$
$\leq\frac{B^{2}}{n^{2}}\sum_{i\in[n]}\sum_{p\in[n]}|f_{c,i,p}({\bm{W}}_{t-\tau})-f_{c,i,p}({\bm{W}}_{t})|$
$=B^{2}\underbrace{\frac{1}{n^{2}}\sum_{i\in[n],y_{i}\neq c}\sum_{p\in[n],y_{p}\neq c}|f_{c,i,p}({\bm{W}}_{t-\tau})-f_{c,i,p}({\bm{W}}_{t})|}_{\spadesuit_{1}}$
$\quad\quad+B^{2}\underbrace{\frac{1}{n^{2}}\sum_{i\in[n],y_{i}\neq c}\sum_{p\in[n],y_{p}=c}|f_{c,i,p}({\bm{W}}_{t-\tau})-f_{c,i,p}({\bm{W}}_{t})|}_{\spadesuit_{2}}$
$\quad\quad+B^{2}\underbrace{\frac{1}{n^{2}}\sum_{i\in[n],y_{i}=c}\sum_{p\in[n],y_{p}\neq c}|f_{c,i,p}({\bm{W}}_{t-\tau})-f_{c,i,p}({\bm{W}}_{t})|}_{\spadesuit_{3}}$
$\quad\quad+B^{2}\underbrace{\frac{1}{n^{2}}\sum_{i\in[n],y_{i}=c}\sum_{p\in[n],y_{p}=c}|f_{c,i,p}({\bm{W}}_{t-\tau})-f_{c,i,p}({\bm{W}}_{t})|}_{\spadesuit_{4}}$
We deal with the $4$ terms $\spadesuit_{1},\spadesuit_{2},\spadesuit_{3}$, and $\spadesuit_{4}$ separately. Starting with the first term, we have
$\spadesuit_{1}$ $=\frac{1}{n^{2}}\sum_{i\in[n],y_{i}\neq c}\sum_{p\in[n],y_{p}\neq c}|\mathbb{S}_{c}({\bm{W}}_{t-\tau}\bm{h}_{i})\mathbb{S}_{c}({\bm{W}}_{t-\tau}\bm{h}_{p})-\mathbb{S}_{c}({\bm{W}}_{t}\bm{h}_{i})\mathbb{S}_{c}({\bm{W}}_{t}\bm{h}_{p})|$
$=\frac{1}{n^{2}}\sum_{i\in[n],y_{i}\neq c}\sum_{p\in[n],y_{p}\neq c}\mathbb{S}_{c}({\bm{W}}_{t}\bm{h}_{i})\mathbb{S}_{c}({\bm{W}}_{t}\bm{h}_{p})|\frac{\mathbb{S}_{c}({\bm{W}}_{t-\tau}\bm{h}_{i})\mathbb{S}_{c}({\bm{W}}_{t-\tau}\bm{h}_{p})}{\mathbb{S}_{c}({\bm{W}}_{t}\bm{h}_{i})\mathbb{S}_{c}({\bm{W}}_{t}\bm{h}_{p})}-1|$
$\stackrel{{\scriptstyle(a)}}{{\leq}}\frac{1}{n^{2}}\sum_{i\in[n],y_{i}\neq c}\sum_{p\in[n],y_{p}\neq c}\mathbb{S}_{c}({\bm{W}}_{t}\bm{h}_{i})\mathbb{S}_{c}({\bm{W}}_{t}\bm{h}_{p})\bigl{(}e^{2\bigl{(}\lVert({\bm{W}}_{t-\tau}-{\bm{W}}_{t})\bm{h}_{i}\rVert_{\infty}+\lVert({\bm{W}}_{t-\tau}-{\bm{W}}_{t})\bm{h}_{p}\rVert_{\infty}\bigr{)}}-1\bigr{)}$
$\stackrel{{\scriptstyle(b)}}{{\leq}}\frac{1}{n^{2}}\sum_{i\in[n],y_{i}\neq c}\sum_{p\in[n],y_{p}\neq c}\mathbb{S}_{c}({\bm{W}}_{t}\bm{h}_{i})\mathbb{S}_{c}({\bm{W}}_{t}\bm{h}_{p})\bigl{(}e^{4B{\left\|{\bm{W}}_{t-\tau}-{\bm{W}}_{t}\right\|_{\max}}}-1\bigr{)}$
$\stackrel{{\scriptstyle(c)}}{{\leq}}\bigl{(}e^{4B\sum_{s=1}^{\tau}\eta_{t-s}{\left\|\frac{\bm{M}_{t-s}}{\sqrt{{\bm{V}}_{t-s}}}\right\|_{\max}}}-1\bigr{)}\frac{1}{n^{2}}\sum_{i\in[n],y_{i}\neq c}\sum_{p\in[n],y_{p}\neq c}\mathbb{S}_{c}({\bm{W}}_{t}\bm{h}_{i})\mathbb{S}_{c}({\bm{W}}_{t}\bm{h}_{p})$
$\stackrel{{\scriptstyle(d)}}{{\leq}}\bigl{(}e^{4B\alpha\sum_{s=1}^{\tau}\eta_{t-s}}-1\bigr{)}\mathcal{Q}_{c}({\bm{W}}_{t})^{2},$
where (a) is by Lemma D.10, (b) is by $\lVert\bm{h}_{i}\rVert_{1}\leq B$ for all $i\in[n]$, (c) is by (3c) and the triangle inequality, and (d) is by Lemma D.1 and the definition of ${\mathcal{G}}({\bm{W}}_{t})$. For the second term, we have
$\spadesuit_{2}$ $=\frac{1}{n^{2}}\sum_{i\in[n],y_{i}\neq c}\sum_{p\in[n],y_{p}=c}|\bigl{(}\mathbb{S}_{c}({\bm{W}}_{t}\bm{h}_{i})-\mathbb{S}_{c}({\bm{W}}_{t-\tau}\bm{h}_{i})\bigr{)}$
$\quad\quad\quad\quad+\bigl{(}\mathbb{S}_{c}({\bm{W}}_{t-\tau}\bm{h}_{i})\mathbb{S}_{c}({\bm{W}}_{t-\tau}\bm{h}_{p})-\mathbb{S}_{c}({\bm{W}}_{t}\bm{h}_{i})\mathbb{S}_{c}({\bm{W}}_{t}\bm{h}_{p})\bigr{)}|$
$=\frac{1}{n^{2}}\sum_{i\in[n],y_{i}\neq c}\sum_{p\in[n],y_{p}=c}|\mathbb{S}_{c}({\bm{W}}_{t}\bm{h}_{i})\bigl{(}1-\mathbb{S}_{c}({\bm{W}}_{t}\bm{h}_{p})\bigr{)}-\bigl{(}1-\mathbb{S}_{c}({\bm{W}}_{t-\tau}\bm{h}_{p})\bigr{)}\mathbb{S}_{c}({\bm{W}}_{t-\tau}\bm{h}_{i})|$
$=\frac{1}{n^{2}}\sum_{i\in[n],y_{i}\neq c}\sum_{p\in[n],y_{p}=c}\mathbb{S}_{c}({\bm{W}}_{t}\bm{h}_{i})\bigl{(}1-\mathbb{S}_{c}({\bm{W}}_{t}\bm{h}_{p})\bigr{)}|1-\frac{\bigl{(}1-\mathbb{S}_{c}({\bm{W}}_{t-\tau}\bm{h}_{p})\bigr{)}\mathbb{S}_{c}({\bm{W}}_{t-\tau}\bm{h}_{i})}{\bigl{(}1-\mathbb{S}_{c}({\bm{W}}_{t}\bm{h}_{p})\bigr{)}\mathbb{S}_{c}({\bm{W}}_{t}\bm{h}_{i})}|$
$\leq\bigl{(}e^{4B\alpha\sum_{s=1}^{\tau}\eta_{t-s}}-1\bigr{)}\mathcal{Q}_{c}({\bm{W}}_{t}){\mathcal{G}}_{c}({\bm{W}}_{t}),$
where the last inequality is by Lemma D.10 and the same steps taken for $\spadesuit_{1}$. The third term can be derived similarly as the second term and we can obtain the same bound as follows:
$\spadesuit_{3}$ $=\frac{1}{n^{2}}\sum_{i\in[n],y_{i}=c}\sum_{p\in[n],y_{p}\neq c}\mathbb{S}_{c}({\bm{W}}_{t}\bm{h}_{p})\bigl{(}1-\mathbb{S}_{c}({\bm{W}}_{t}\bm{h}_{i})\bigr{)}|1-\frac{\bigl{(}1-\mathbb{S}_{c}({\bm{W}}_{t-\tau}\bm{h}_{p})\bigr{)}\mathbb{S}_{c}({\bm{W}}_{t-\tau}\bm{h}_{i})}{\bigl{(}1-\mathbb{S}_{c}({\bm{W}}_{t}\bm{h}_{i})\bigr{)}\mathbb{S}_{c}({\bm{W}}_{t}\bm{h}_{p})}|$
$\leq\bigl{(}e^{4B\alpha\sum_{s=1}^{\tau}\eta_{t-s}}-1\bigr{)}\mathcal{Q}_{c}({\bm{W}}_{t}){\mathcal{G}}_{c}({\bm{W}}_{t}).$
For the fourth term, we obtain:
$\spadesuit_{4}$ $=\frac{1}{n^{2}}\sum_{i\in[n],y_{i}=c}\sum_{p\in[n],y_{p}=c}|\bigl{(}1-\mathbb{S}_{c}({\bm{W}}_{t-\tau}\bm{h}_{i})\bigr{)}\bigl{(}1-\mathbb{S}_{c}({\bm{W}}_{t-\tau}\bm{h}_{p})\bigr{)}-\bigl{(}1-\mathbb{S}_{c}({\bm{W}}_{t}\bm{h}_{i})\bigr{)}\bigl{(}1-\mathbb{S}_{c}({\bm{W}}_{t}\bm{h}_{p})\bigr{)}|$
$=\frac{1}{n^{2}}\sum_{i\in[n],y_{i}=c}\sum_{p\in[n],y_{p}=c}\bigl{(}1-\mathbb{S}_{c}({\bm{W}}_{t}\bm{h}_{i})\bigr{)}\bigl{(}1-\mathbb{S}_{c}({\bm{W}}_{t}\bm{h}_{p})\bigr{)}|\frac{\bigl{(}1-\mathbb{S}_{c}({\bm{W}}_{t-\tau}\bm{h}_{i})\bigr{)}\bigl{(}1-\mathbb{S}_{c}({\bm{W}}_{t-\tau}\bm{h}_{p})\bigr{)}}{\bigl{(}1-\mathbb{S}_{c}({\bm{W}}_{t}\bm{h}_{i})\bigr{)}\bigl{(}1-\mathbb{S}_{c}({\bm{W}}_{t}\bm{h}_{p})\bigr{)}}-1|$
$\leq\bigl{(}e^{4B\alpha\sum_{s=1}^{\tau}\eta_{t-s}}-1\bigr{)}{\mathcal{G}}_{c}({\bm{W}}_{t})^{2},$
where the last inequality is by Lemma D.10 and the same steps taken for $\spadesuit_{1}$. We combine the bounds for $\spadesuit_{1}$, $\spadesuit_{2}$, $\spadesuit_{3}$, and $\spadesuit_{4}$ to obtain: $\spadesuit\leq 4B^{2}\bigl{(}e^{4B\alpha\sum_{s=1}^{\tau}\eta_{t-s}}-1\bigr{)}({\mathcal{G}}_{c}({\bm{W}}_{t})+\mathcal{Q}_{c}({\bm{W}}_{t}))^{2}$. Then, we substitute this into (23) to obtain:
$|{\bm{V}}_{t}[c,j]-(1-\beta_{2}^{t+1})\nabla\mathcal{L}({\bm{W}}_{t})[c,j]^{2}|$ $\leq B^{2}(1-\beta_{2})(\mathcal{Q}_{c}({\bm{W}}_{t})+{\mathcal{G}}_{c}({\bm{W}}_{t}))^{2}\sum_{\tau=0}^{t}\beta_{2}^{\tau}\bigl{(}e^{4\alpha B\sum_{s=1}^{\tau}\eta_{t-s}}-1\bigr{)}$
$\leq B^{2}(1-\beta_{2})c_{2}\eta_{t}(\mathcal{Q}_{c}({\bm{W}}_{t})+{\mathcal{G}}_{c}({\bm{W}}_{t}))^{2},$
where the last inequality is by the Assumption 3.4. The final result follows from the fact that $|p-q|^{2}\leq|p^{2}-q^{2}|$ when both $p$ and $q$ are positive. ∎
The following Lemma bounds the term $\bigm{|}\langle\nabla\mathcal{L}({\bm{W}}_{t}),\frac{\bm{M}_{t}}{\sqrt{{\bm{V}}_{t}}}-\frac{\nabla\mathcal{L}({\bm{W}}_{t})}{|\nabla\mathcal{L}({\bm{W}}_{t})|}\rangle\bigm{|}$ using ${\mathcal{G}}({\bm{W}}_{t})$. It is used in Lemma D.5 to show the decrease in the risk. The proof is similar to that of Zhang et al. (2024, Lemma A.3), but here we need to carefully track the index $c\in[k]$ using both ${\mathcal{G}}_{c}({\bm{W}})$ and $\mathcal{Q}_{c}({\bm{W}})$ to avoid $k$ dependence. The final result crucially relies on the decomposition ${\mathcal{G}}({\bm{W}}_{t})=\sum_{c\in[k]}\mathcal{T}_{c}({\bm{W}}_{t})=\sum_{c\in[k]}\mathcal{Q}_{c}({\bm{W}}_{t})$.
Lemma D.4.
Under the same setting as Theorem C.4, we have
$\bigm{|}\langle\nabla\mathcal{L}({\bm{W}}_{t}),\frac{\bm{M}_{t}}{\sqrt{{\bm{V}}_{t}}}-\frac{\nabla\mathcal{L}({\bm{W}}_{t})}{|\nabla\mathcal{L}({\bm{W}}_{t})|}\rangle\bigm{|}$ $\leq 4\sqrt{\frac{\beta_{1}^{t+1}}{1-\beta_{2}^{t+1}}}{\left\|\nabla\mathcal{L}({\bm{W}}_{t})\right\|_{\rm{sum}}}+$
$\quad\quad\frac{2d}{\sqrt{1-\beta_{2}}}\bigl{(}\frac{6\alpha_{V}}{\sqrt{1-\beta_{2}^{t+1}}}\sqrt{\eta_{t}}+3\alpha_{M}\eta_{t}\bigr{)}{\mathcal{G}}({\bm{W}}_{t}).$
Proof.
For simplicity, we drop the subscripts $t$. Denote $\mathcal{T}_{c}({\bm{W}})\coloneqq{\mathcal{G}}_{c}({\bm{W}})+\mathcal{Q}_{c}({\bm{W}})$. Then, by Lemmas D.2 and D.3, we have for any $c\in[k]$ and $j\in[d]$:
$\bm{M}[c,j]$ $=(1-\beta_{1}^{t+1})\nabla\mathcal{L}({\bm{W}})[c,j]+\alpha_{M}\eta_{t}\mathcal{T}_{c}({\bm{W}})\epsilon_{m,c,j},$ (24)
$\sqrt{{\bm{V}}[c,j]}$ $=\sqrt{1-\beta_{2}^{t+1}}|\nabla\mathcal{L}({\bm{W}})[c,j]|+\alpha_{V}\sqrt{\eta_{t}}\mathcal{T}_{c}({\bm{W}})\epsilon_{v,c,j},$ (25)
where $|\epsilon_{m,c,j}|\leq 1$ and $|\epsilon_{v,c,j}|\leq 1$ are some residual terms. We denote $\psi_{c,j}\coloneqq\nabla\mathcal{L}({\bm{W}})[c,j](\frac{\bm{M}[c,j]}{\sqrt{{\bm{V}}[c,j]}}-\frac{\nabla\mathcal{L}({\bm{W}})[c,j]}{|\nabla\mathcal{L}({\bm{W}})[c,j]|})$, the set of index $E_{c,j}\coloneqq\{j\in[d]\Bigm{|}\sqrt{1-\beta_{2}^{t+1}}|\nabla\mathcal{L}({\bm{W}})[c,j]|\geq 2\alpha_{V}\sqrt{\eta}_{t}\mathcal{T}_{c}({\bm{W}})|\epsilon_{v,c,j}|\}$, and its complement $E_{c,j}^{c}=[d]\backslash E_{c,j}$. The goal is to bound $|\psi_{c,j}|$ when $j\in E_{c,j}^{c}$ or $j\in E_{c,j}$ using $\mathcal{T}_{c}({\bm{W}})$. We start with the indices in $E_{c,j}^{c}$:
$\sum_{j\in E_{c,j}^{c}}|\psi_{c,j}|$ $\leq\sum_{j\in E_{c,j}^{c}}|\nabla\mathcal{L}({\bm{W}})[c,j]|\bigl{(}\frac{|\bm{M}[c,j]|}{\sqrt{{\bm{V}}[c,j]}}+1\bigr{)}$
$\stackrel{{\scriptstyle(a)}}{{\leq}}\sum_{j\in E_{c,j}^{c}}|\nabla\mathcal{L}({\bm{W}})[c,j]|\bigl{(}\frac{(1-\beta_{1}^{t+1})|\nabla\mathcal{L}({\bm{W}})[c,j]|+\alpha_{M}\eta_{t}\mathcal{T}_{c}({\bm{W}})}{\sqrt{1-\beta_{2}}|\nabla\mathcal{L}({\bm{W}})[c,j]|}+1\bigr{)}$
$\stackrel{{\scriptstyle(b)}}{{\leq}}\sum_{j\in E_{c,j}^{c}}(\frac{1-\beta_{1}^{t+1}}{\sqrt{1-\beta_{2}}}+1)\frac{2\alpha_{V}\sqrt{\eta_{t}}\mathcal{T}_{c}({\bm{W}})}{\sqrt{1-\beta_{2}^{t+1}}}+\frac{\alpha_{M}\eta_{t}\mathcal{T}_{c}({\bm{W}})}{\sqrt{1-\beta_{2}}}$
$\leq\frac{d}{\sqrt{1-\beta_{2}}}\bigl{(}\frac{4\alpha_{V}}{\sqrt{1-\beta_{2}^{t+1}}}\sqrt{\eta_{t}}+\alpha_{M}\eta_{t}\bigr{)}\mathcal{T}_{c}({\bm{W}}),$
where (a) is by (24), $|\epsilon_{m,c,j}|\leq 1$, and ${\bm{V}}[c,j]\geq(1-\beta_{2})\nabla\mathcal{L}({\bm{W}})[c,j]^{2}$; and (b) is by $j\in E^{c}_{c,j}$ s.t. $|\nabla\mathcal{L}({\bm{W}})[c,j]|\leq\frac{2\alpha_{V}\sqrt{\eta_{t}}\mathcal{T}_{c}({\bm{W}})}{\sqrt{1-\beta_{2}^{t+1}}}$. Next, we focus on the indices $j\in E_{c,j}$. In this case, we have
$\psi_{c,j}$ $=\nabla\mathcal{L}({\bm{W}})[c,j]\bigl{(}\frac{\bm{M}[c,j]}{\sqrt{1-\beta_{2}^{t+1}}|\nabla\mathcal{L}({\bm{W}})[c,j]|+\alpha_{V}\sqrt{\eta_{t}}\mathcal{T}_{c}({\bm{W}})\epsilon_{v,c,j}}-\frac{\mathcal{L}({\bm{W}})[c,j]}{|\nabla\mathcal{L}({\bm{W}})[c,j]|}\bigr{)}$
$=\nabla\mathcal{L}({\bm{W}})[c,j]\underbrace{\frac{\bm{M}[c,j]|\nabla\mathcal{L}({\bm{W}})[c,j]|-\bigl{(}\sqrt{1-\beta_{2}^{t+1}}|\nabla\mathcal{L}({\bm{W}})[c,j]|+\alpha_{V}\sqrt{\eta_{t}}\mathcal{T}_{c}({\bm{W}})\epsilon_{v,c,j}\bigr{)}\nabla\mathcal{L}({\bm{W}})[c,j]}{\bigl{(}\sqrt{1-\beta_{2}^{t+1}}|\nabla\mathcal{L}({\bm{W}})[c,j]|+\alpha_{V}\sqrt{\eta_{t}}\mathcal{T}_{c}({\bm{W}})\epsilon_{v,c,j}\bigr{)}|\nabla\mathcal{L}({\bm{W}})[c,j]|}}_{\frac{\spadesuit_{1}}{\spadesuit_{2}}},$
where
$\Bigm{|}\spadesuit_{1}\Bigm{|}$ $=\Bigm{|}\bigl{(}1-\beta_{1}^{t+1}-\sqrt{1-\beta_{2}^{t+1}}\bigr{)}\nabla\mathcal{L}({\bm{W}})[c,j]|\nabla\mathcal{L}({\bm{W}})[c,j]|+$
$\quad\quad\alpha_{M}\eta_{t}\mathcal{T}_{c}({\bm{W}})\epsilon_{m,c,j}|\nabla\mathcal{L}({\bm{W}})[c,j]|-\alpha_{V}\sqrt{\eta_{t}}\mathcal{T}_{c}({\bm{W}})\epsilon_{v,c,j}\nabla\mathcal{L}({\bm{W}})[c,j]\Bigm{|}$
$\stackrel{{\scriptstyle(c)}}{{\leq}}\bigm{|}1-\beta_{1}^{t+1}-\sqrt{1-\beta_{2}^{t+1}}||\nabla\mathcal{L}({\bm{W}})[c,j]|^{3}+(\alpha_{M}\eta_{t}+\alpha_{V}\sqrt{\eta_{t}})\mathcal{T}_{c}({\bm{W}})|\nabla\mathcal{L}({\bm{W}})[c,j]|^{2},$
and
$\Bigm{|}\spadesuit_{2}\Bigm{|}=\spadesuit_{2}\stackrel{{\scriptstyle(d)}}{{\geq}}\frac{1}{2}\sqrt{1-\beta_{2}^{t+1}}|\nabla\mathcal{L}({\bm{W}})[c,j]|^{2}.$
Inequality (c) is by $|\epsilon_{m,c,j}|\leq 1$ and $|\epsilon_{v,c,j}|\leq 1$, and (d) is by $\alpha_{V}\sqrt{\eta_{t}}\mathcal{T}_{c}({\bm{W}})\epsilon_{v,c,j}\geq-\frac{1}{2}\sqrt{1-\beta_{2}^{t+1}}|\nabla\mathcal{L}({\bm{W}})[c,j]|$ for any $j\in E_{c,j}$. Putting these two pieces together, we obtain
$\sum_{j\in E_{c,j}}|\psi_{c,j}|$ $\leq\sum_{j\in E_{c,j}}\frac{|1-\beta_{1}^{t+1}-\sqrt{1-\beta_{2}^{t+1}}||\nabla\mathcal{L}({\bm{W}})[c,j]|^{3}+(\alpha_{M}\eta_{t}+\alpha_{V}\sqrt{\eta_{t}})\mathcal{T}_{c}({\bm{W}})|\nabla\mathcal{L}({\bm{W}})[c,j]|^{2}}{\frac{1}{2}\sqrt{1-\beta_{2}^{t+1}}|\nabla\mathcal{L}({\bm{W}})[c,j]|^{2}}$
$\stackrel{{\scriptstyle(e)}}{{\leq}}\Bigl{(}\sum_{j\in E_{c,j}}4\sqrt{\frac{\beta_{1}^{t+1}}{1-\beta_{2}^{t+1}}}|\nabla\mathcal{L}({\bm{W}})[c,j]|\Bigr{)}+d\bigl{(}\frac{2\alpha_{V}}{\sqrt{1-\beta_{2}^{t+1}}}\sqrt{\eta_{t}}+\frac{2\alpha_{M}}{\sqrt{1-\beta_{2}^{t+1}}}\eta_{t}\bigr{)}\mathcal{T}_{c}({\bm{W}})$
$\leq 4\sqrt{\frac{\beta_{1}^{t+1}}{1-\beta_{2}^{t+1}}}{\left\|\nabla\mathcal{L}({\bm{W}})[c,:]\right\|_{\rm{sum}}}+\frac{d}{\sqrt{1-\beta_{2}}}\bigl{(}\frac{2\alpha_{V}}{\sqrt{1-\beta_{2}^{t+1}}}\sqrt{\eta_{t}}+2\alpha_{M}\eta_{t}\bigr{)}\mathcal{T}_{c}({\bm{W}}),$
where (e) is by $\sqrt{a}\leq\sqrt{a-b}+\sqrt{b}$ implying $1-\sqrt{1-\beta_{2}^{t+1}}\leq\beta_{1}^{\frac{t+1}{2}}$, and $\nabla\mathcal{L}({\bm{W}})[c,:]$ denotes the $c$th row of $\nabla\mathcal{L}({\bm{W}})$. Finally, we note that $|\langle\nabla\mathcal{L}({\bm{W}}),\frac{\bm{M}}{\sqrt{{\bm{V}}}}-\frac{\nabla\mathcal{L}({\bm{W}})}{|\nabla\mathcal{L}({\bm{W}})|}\rangle|=|\sum_{(c,j)}\psi_{c,j}|\leq\sum_{(c,j)}|\psi_{c,j}|$. Then, we obtain
$\sum_{c,j}|\psi_{c,j}|$ $=\sum_{c\in[k]}\bigl{(}\sum_{j\in E^{c}_{c,j}}|\psi_{c,j}|+\sum_{j\in E_{c,j}}|\psi_{c,j}|\bigr{)}$
$=\sum_{c\in[k]}4\sqrt{\frac{\beta_{1}^{t+1}}{1-\beta_{2}^{t+1}}}{\left\|\nabla\mathcal{L}({\bm{W}})[c,:]\right\|_{\rm{sum}}}+\sum_{c\in[k]}\frac{d}{\sqrt{1-\beta_{2}}}\bigl{(}\frac{2\alpha_{V}}{\sqrt{1-\beta_{2}^{t+1}}}\sqrt{\eta_{t}}+2\alpha_{M}\eta_{t}\bigr{)}\mathcal{T}_{c}({\bm{W}})$
=$\stackrel{{\scriptstyle(f)}}{{=}}4\sqrt{\frac{\beta_{1}^{t+1}}{1-\beta_{2}^{t+1}}}{\left\|\nabla\mathcal{L}({\bm{W}})\right\|_{\rm{sum}}}+\frac{2d}{\sqrt{1-\beta_{2}}}\bigl{(}\frac{2\alpha_{V}}{\sqrt{1-\beta_{2}^{t+1}}}\sqrt{\eta_{t}}+2\alpha_{M}\eta_{t}\bigr{)}{\mathcal{G}}({\bm{W}}),$
where (f) is by $\sum_{c\in[k]}\mathcal{T}_{c}({\bm{W}})=\sum_{c\in[k]}\mathcal{Q}_{c}({\bm{W}})+{\mathcal{G}}_{c}({\bm{W}})=2{\mathcal{G}}({\bm{W}})$. ∎
Lemma D.5 (Adam Descent).
Under the same setting as Theorem D.8, set $t_{A}\coloneqq\frac{2\log(\frac{\sqrt{1-\beta_{2}}}{4})}{\log(\beta_{1})}$, then we have for all $t\geq t_{A}$
$\mathcal{L}({\bm{W}}_{t+1})$ $\leq\mathcal{L}({\bm{W}}_{t})-\eta_{t}\gamma\bigl{(}1-\alpha_{a_{1}}\beta_{1}^{t/2}-\alpha_{a_{2}}d\eta_{t}^{\frac{1}{2}}-\alpha_{a_{3}}d\eta_{t}\bigr{)}{\mathcal{G}}({\bm{W}}_{t}),$
where $\alpha_{a_{1}}$, $\alpha_{a_{2}}$, and $\alpha_{a_{3}}$ are some constants that depend on $B$, $\gamma$,$\beta_{1}$, and $\beta_{2}$.
Proof.
We follow the same notations and strategy of Lemma C.1, and recall the definitions $\spadesuit_{t}=\langle\nabla\mathcal{L}({\bm{W}}_{t}),\bm{\Delta}_{t}\rangle$ and $\clubsuit_{t}=\bm{h}_{i}^{\top}\bm{\Delta}_{t}^{\top}\left(\operatorname{diag}(\mathbb{S}({\bm{W}}_{t,t+1,\gamma}\bm{h}_{i}))-\mathbb{S}({\bm{W}}_{t,t+1,\zeta^{*}}\bm{h}_{i})\mathbb{S}({\bm{W}}_{t,t+1,\zeta^{*}}\bm{h}_{i})^{\top}\right)\bm{\Delta}_{t}\,\bm{h}_{i}$. In the case of Adam, we have $\bm{\Delta}_{t}=\frac{\bm{M}_{t}}{\sqrt{{\bm{V}}_{t}}}$. We bound $\spadesuit_{t}$ and $\clubsuit_{t}$ separately. Starting with $\spadesuit_{t}$, we have for all $t\geq t_{A}$
$\spadesuit_{t}$ $=-\eta_{t}\langle\nabla\mathcal{L}({\bm{W}}_{t}),\frac{\bm{M}_{t}}{\sqrt{{\bm{V}}_{t}}}\rangle$
$=-\eta_{t}\bigl{(}\langle\nabla\mathcal{L}({\bm{W}}_{t}),\frac{\bm{M}_{t}}{\sqrt{{\bm{V}}_{t}}}-\frac{\nabla\mathcal{L}({\bm{W}}_{t})}{|\nabla\mathcal{L}({\bm{W}}_{t})|}\rangle+\langle\nabla\mathcal{L}({\bm{W}}_{t}),\frac{\nabla\mathcal{L}({\bm{W}}_{t})}{|\nabla\mathcal{L}({\bm{W}}_{t})|}\rangle\bigr{)}$
$\leq-\eta_{t}{\left\|\nabla\mathcal{L}({\bm{W}}_{t})\right\|_{\rm{sum}}}+\eta_{t}\bigm{|}\langle\nabla\mathcal{L}({\bm{W}}_{t}),\frac{\bm{M}_{t}}{\sqrt{{\bm{V}}_{t}}}-\frac{\nabla\mathcal{L}({\bm{W}}_{t})}{|\nabla\mathcal{L}({\bm{W}}_{t})|}\rangle\bigm{|}$
$\stackrel{{\scriptstyle(a)}}{{\leq}}-\eta_{t}\bigl{(}1-4\sqrt{\frac{\beta_{1}^{t+1}}{1-\beta_{2}^{t+1}}}\bigr{)}{\left\|\nabla\mathcal{L}({\bm{W}}_{t})\right\|_{\rm{sum}}}+\frac{2d}{\sqrt{1-\beta_{2}}}\bigl{(}\frac{6\alpha_{V}}{\sqrt{1-\beta_{2}^{t+1}}}\eta_{t}^{\frac{3}{2}}+3\alpha_{M}\eta_{t}^{2}\bigr{)}{\mathcal{G}}({\bm{W}}_{t})$
$\leq-\eta_{t}\bigl{(}1-4\frac{\beta_{1}^{\frac{t}{2}}}{\sqrt{1-\beta_{2}}}\bigr{)}{\left\|\nabla\mathcal{L}({\bm{W}}_{t})\right\|_{\rm{sum}}}+\frac{12\alpha_{V}}{1-\beta_{2}}d\eta_{t}^{3/2}{\mathcal{G}}({\bm{W}}_{t})+\frac{6\alpha_{M}}{\sqrt{1-\beta_{2}}}d\eta_{t}^{2}{\mathcal{G}}({\bm{W}}_{t})$
$\stackrel{{\scriptstyle(b)}}{{\leq}}-\eta_{t}\gamma\bigl{(}1-4\frac{\beta_{1}^{\frac{t}{2}}}{\sqrt{1-\beta_{2}}}\bigr{)}{\mathcal{G}}({\bm{W}}_{t})+\frac{12\alpha_{V}}{1-\beta_{2}}d\eta_{t}^{3/2}{\mathcal{G}}({\bm{W}}_{t})+\frac{6\alpha_{M}}{\sqrt{1-\beta_{2}}}d\eta_{t}^{2}{\mathcal{G}}({\bm{W}}_{t}),$
where (a) is by Lemma D.4, and (b) is by Lemma B.1. For $\clubsuit_{t}$, we apply Lemma A.3 to obtain
$\clubsuit_{t}\leq 4\|\bm{\Delta}_{t}\bm{h}_{i}\|_{\infty}^{2}(1-\mathbb{S}_{y_{i}}({\bm{W}}_{t,t+1,\zeta^{*}}\bm{h}_{i}))\leq 4\eta_{t}^{2}\alpha^{2}B^{2}(1-\mathbb{S}_{y_{i}}({\bm{W}}_{t,t+1,\zeta^{*}}\bm{h}_{i})),$
where in the second inequality we have used $\|\bm{\Delta}_{t}\bm{h}_{i}\|_{\infty}\leq{\left\|\bm{\Delta}_{t}\right\|_{\max}}\|\bm{h}_{i}\|_{1}$, $\|\bm{h}_{i}\|_{1}\leq B$, and ${\left\|\bm{\Delta}_{t}\right\|_{\max}}=\eta_{t}{\left\|\frac{\bm{M}_{t}}{\sqrt{{\bm{V}}_{t}}}\right\|_{\max}}\leq\eta_{t}\alpha$ by Lemma D.1 given $t\geq t_{A}$ implying that $1\geq 4\frac{\beta_{1}^{\frac{t}{2}}}{\sqrt{1-\beta_{2}}}$. Combing this with Lemma A.3, we obtain
$\frac{1}{2n}\sum_{i\in[n]}\bm{h}_{i}^{\top}\bm{\Delta}_{t}^{\top}$ $\left(\operatorname{diag}(\mathbb{S}({\bm{W}}_{t,t+1,\gamma}\bm{h}_{i}))-\mathbb{S}({\bm{W}}_{t,t+1,\zeta^{*}}\bm{h}_{i})\mathbb{S}({\bm{W}}_{t,t+1,\zeta^{*}}\bm{h}_{i})^{\top}\right)\bm{\Delta}_{t}\,\bm{h}_{i}$
$\leq\frac{1}{2n}\sum_{i\in[n]}4\eta_{t}^{2}\alpha^{2}B^{2}(1-\mathbb{S}_{y_{i}}({\bm{W}}_{t,t+1,\zeta^{*}}\bm{h}_{i}))\leq 2\alpha^{2}\eta_{t}^{2}B^{2}e^{2B\eta_{0}}{\mathcal{G}}({\bm{W}}_{t}),$
where the derivation of the second inequality can be found in the derivation of 18. Putting everything together, we obtain
$\mathcal{L}({\bm{W}}_{t+1})$ $\leq\mathcal{L}({\bm{W}}_{t})-\eta_{t}\gamma{\mathcal{G}}({\bm{W}}_{t})+4\frac{\beta_{1}^{\frac{t}{2}}}{\sqrt{1-\beta_{2}}}\gamma\eta_{t}{\mathcal{G}}({\bm{W}}_{t})+\frac{12\alpha_{V}}{1-\beta_{2}}d\eta_{t}^{3/2}{\mathcal{G}}({\bm{W}}_{t})+$
$\quad\quad\quad\bigl{(}\frac{6\alpha_{M}}{\sqrt{1-\beta_{2}}}+2\alpha^{2}B^{2}e^{2B\eta_{0}}\bigr{)}d\eta_{t}^{2}{\mathcal{G}}({\bm{W}}_{t})$
$=\mathcal{L}({\bm{W}}_{t})-\eta_{t}\gamma\bigl{(}1-\alpha_{a_{1}}\beta_{1}^{t/2}-\alpha_{a_{2}}d\eta_{t}^{\frac{1}{2}}-\alpha_{a_{3}}d\eta_{t}\bigr{)}{\mathcal{G}}({\bm{W}}_{t}),$
where we have defined $\alpha_{a_{1}}\coloneqq\frac{4}{\sqrt{1-\beta_{2}}}$, $\alpha_{a_{2}}\coloneqq\frac{12\alpha_{V}}{\gamma(1-\beta_{2})}$, and $\alpha_{a_{3}}\coloneqq\frac{6\alpha_{M}}{\gamma\sqrt{1-\beta_{2}}}+\frac{2\alpha^{2}B^{2}e^{2B\eta_{0}}}{\gamma}$. ∎
Built upon Lemma D.5, we can further lower bound the unnormalized margin of Adam iterates for a sufficiently large $t$. The proof is similar to that of SignGD (i.e., Lemma C.2), which crucially depends on the separability condition obtained after achieving a low loss (Lemma B.4). The time $\tilde{t}_{A}$ will be specified in the proof of Theorem D.8.
Lemma D.6 (Adam Unnormalized Margin).
Under the same setting as Theorem D.8, suppose that there exist $\tilde{t}$ such that $\mathcal{L}({\bm{W}}_{t})\leq\frac{\log 2}{n}$ for all $t>\tilde{t}$, then we have for all $t\geq\tilde{t}_{A}\coloneqq\max\{t_{A},\tilde{t}\}$
$\min_{i\in[n],c\neq y_{i}}(\bm{e}_{y_{i}}-\bm{e}_{c})^{T}{\bm{W}}_{t}\bm{h}_{i}\geq\gamma\sum_{s=\tilde{t}_{A}}^{t-1}\eta_{s}\frac{{\mathcal{G}}({\bm{W}}_{s})}{\mathcal{L}({\bm{W}}_{s})}-\alpha_{a_{5}}d\sum_{s=\tilde{t}_{A}}^{t-1}\eta_{s}^{\frac{3}{2}}-\alpha_{a_{6}}d\sum_{s=\tilde{t}_{A}}^{t-1}\eta_{s}^{2}-\alpha_{a_{7}},$
where $t_{A}=\frac{2\log(\frac{\sqrt{1-\beta_{2}}}{4})}{\log(\beta_{1})}$, and $\alpha_{a_{5}}$, $\alpha_{a_{6}}$, and $\alpha_{a_{7}}$ are some constants that depend on $B$, $\beta_{1}$, and $\beta_{2}$.
Proof.
We denote $\alpha_{a_{4}}\coloneqq\frac{4}{\sqrt{1-\beta_{2}}}$, $\alpha_{a_{5}}\coloneqq\frac{12\alpha_{V}}{1-\beta_{2}}$, and $\alpha_{a_{6}}\coloneqq\frac{6\alpha_{M}}{\sqrt{1-\beta_{2}}}+2\alpha^{2}B^{2}e^{2B\eta_{0}}$. Under the assumption that $\mathcal{L}({\bm{W}}_{t})\leq\frac{\log 2}{n}$ for all $t\geq\tilde{t}$, we have for all $t\geq\tilde{t}_{A}\coloneqq\max\{t_{A},\tilde{t}\}$ (recall that $t_{A}=\frac{2\log(\frac{\sqrt{1-\beta_{2}}}{4})}{\log(\beta_{1})}$)
$\mathcal{L}({\bm{W}}_{t+1})$ $\stackrel{{\scriptstyle(a)}}{{\leq}}\mathcal{L}({\bm{W}}_{t})-\eta_{t}\gamma{\mathcal{G}}({\bm{W}}_{t})+\alpha_{a_{4}}\beta_{1}^{\frac{t}{2}}\gamma\eta_{t}{\mathcal{G}}({\bm{W}}_{t})+\alpha_{a_{5}}kd\eta_{t}^{\frac{3}{2}}{\mathcal{G}}({\bm{W}}_{t})+\alpha_{a_{6}}kd\eta_{t}^{2}{\mathcal{G}}({\bm{W}}_{t})$
$\stackrel{{\scriptstyle(b)}}{{\leq}}\mathcal{L}({\bm{W}}_{t})\bigl{(}1-\eta_{t}\gamma\frac{{\mathcal{G}}({\bm{W}}_{t})}{\mathcal{L}({\bm{W}}_{t})}+\alpha_{a_{4}}\beta_{1}^{\frac{t}{2}}\gamma\eta_{t}+\alpha_{a_{5}}d\eta_{t}^{\frac{3}{2}}+\alpha_{a_{6}}d\eta_{t}^{2}\bigr{)}$
$\leq\mathcal{L}({\bm{W}}_{\tilde{t}_{A}})\exp\bigl{(}-\gamma\sum_{s=\tilde{t}_{A}}^{t}\eta_{s}\frac{{\mathcal{G}}({\bm{W}}_{s})}{\mathcal{L}({\bm{W}}_{s})}+\alpha_{a_{4}}\gamma\sum_{s=\tilde{t}_{A}}^{t}\beta_{1}^{\frac{s}{2}}\eta_{s}+\alpha_{a_{5}}d\sum_{s=\tilde{t}_{A}}^{t}\eta_{s}^{\frac{3}{2}}+\alpha_{a_{6}}d\sum_{s=\tilde{t}_{A}}^{t}\eta_{s}^{2}\bigr{)}$
$\stackrel{{\scriptstyle(c)}}{{\leq}}\frac{\log 2}{n}\exp\bigl{(}-\gamma\sum_{s=\tilde{t}_{A}}^{t}\eta_{s}\frac{{\mathcal{G}}({\bm{W}}_{s})}{\mathcal{L}({\bm{W}}_{s})}+\alpha_{a_{5}}d\sum_{s=\tilde{t}_{A}}^{t}\eta_{s}^{\frac{3}{2}}+\alpha_{a_{6}}d\sum_{s=\tilde{t}_{A}}^{t}\eta_{s}^{2}+\alpha_{a_{7}}\bigr{)},$
where (a) is by Lemma D.5, (b) is by $\frac{{\mathcal{G}}({\bm{W}}_{t})}{\mathcal{L}({\bm{W}}_{t})}\leq 1$ (shown in Lemma B.3), and (c) is by $\mathcal{L}({\bm{W}}_{\tilde{t}_{A}})\leq\frac{\log 2}{n}$ and $\alpha_{a_{4}}\gamma\sum_{s=\tilde{t}_{A}}^{t}\beta_{1}^{\frac{s}{2}}\eta_{s}\leq\frac{\alpha_{a_{4}}\gamma\eta_{0}}{1-\beta_{1}^{\frac{1}{2}}}\eqqcolon\alpha_{a_{7}}$ The rest of the proof follows the same arguments in Lemma C.2. Namely, the assumption $\mathcal{L}({\bm{W}}_{t})\leq\frac{\log 2}{n}$ implies that $\min_{c\neq y_{i}}(\bm{e}_{y_{i}}-\bm{e}_{c})^{T}{\bm{W}}_{t}\bm{h}_{i}\geq 0$ for all $i\in[n]$. This separability condition can be used further to show that for all $t\geq\tilde{t}_{A}$
$e^{-\min_{i\in[n],c\neq y_{i}}(\bm{e}_{y_{i}}-\bm{e}_{c})^{T}{\bm{W}}_{t}\bm{h}_{i}}\leq\exp\bigl{(}-\gamma\sum_{s=\tilde{t}_{A}}^{t-1}\eta_{s}\frac{{\mathcal{G}}({\bm{W}}_{s})}{\mathcal{L}({\bm{W}}_{s})}+\alpha_{a_{5}}d\sum_{s=\tilde{t}_{A}}^{t-1}\eta_{s}^{\frac{3}{2}}+\alpha_{a_{6}}d\sum_{s=\tilde{t}_{A}}^{t-1}\eta_{s}^{2}+\alpha_{a_{7}}\bigr{)}.$
Taking the $\log$ on both sides leads to the final result. ∎
Next lemma upper bounds the max-norm of Adam iterates. It involves showing that the risk upper bounds entry-wise second moment, which will become small after the risk starts to monotonically decrease. Its proof can be found in Zhang et al. (2024, Lemma 6.4). Here, we only show the steps that are specific in our settings.
Lemma D.7 (Adam ${\left\|{\bm{W}}_{t}\right\|_{\max}}$).
Under the same setting as Theorem D.8, suppose that there exists $\tilde{t}_{B}>\log(\frac{1}{\omega})$ such that $\mathcal{L}({\bm{W}}_{t})\leq\frac{1}{\sqrt{4B^{2}+\alpha_{V}\eta_{0}}}$ for all $t\geq\tilde{t}_{B}$, then we have
${\left\|{\bm{W}}_{t}\right\|_{\max}}\leq\alpha_{a_{8}}\sum_{s=0}^{\tilde{t}_{B}-1}\eta_{s}+\sum_{s=\tilde{t}_{B}}^{t-1}\eta_{s}+{\left\|{\bm{W}}_{0}\right\|_{\max}},$
where $\alpha_{a_{8}}$ is some constant that depends on $B$, $\beta_{1}$, and $\beta_{2}$.
Proof.
For any $c\in[k]$ and $j\in[d]$, we have for all $t\geq\tilde{t}_{B}$
${\bm{V}}_{t}[c,j]$ $\stackrel{{\scriptstyle(a)}}{{\leq}}(1-\beta_{2}^{t+1})\nabla\mathcal{L}({\bm{W}}_{t})[c,j]^{2}+\alpha_{V}\eta_{t}{\mathcal{G}}({\bm{W}}_{t})^{2}$
$\leq\nabla\mathcal{L}({\bm{W}}_{t})[c,j]^{2}+\alpha_{V}\eta_{t}{\mathcal{G}}({\bm{W}}_{t})^{2}$
$\stackrel{{\scriptstyle(b)}}{{\leq}}4B^{2}{\mathcal{G}}({\bm{W}}_{t})^{2}+\alpha_{V}\eta_{0}{\mathcal{G}}({\bm{W}}_{t})^{2}$
$\stackrel{{\scriptstyle(c)}}{{\leq}}(4B^{2}+\alpha_{V}\eta_{0})\mathcal{L}({\bm{W}}_{t})^{2}\stackrel{{\scriptstyle(d)}}{{\leq}}1,$
where (a) is by Lemma D.2, (b) is by Lemma B.1, (c) is by Lemma B.3, and (d) is by the assumption. This implies that for all $t\geq\tilde{t}_{B}$
$0\geq\log({\bm{V}}_{t}[c,j])\geq\log(\beta_{2}^{t}(1-\beta_{2})\mathcal{L}({\bm{W}}_{t})[c,j]^{2})\stackrel{{\scriptstyle(e)}}{{\geq}}t\log(\beta_{2})+\log(1-\beta_{2})+\log(\omega),$
where (e) is by the Assumption 3.2. The rest proof follows the same arguments in Zhang et al. (2024, Lemma 6.4). ∎
Theorem D.8.
Suppose that Assumption 3.1, 3.2, 3.3, 3.4, and 3.5 hold, and $\beta_{1}\leq\beta_{2}$, then there exists $t_{a_{2}}=t_{a_{2}}(n,d,\gamma,B,{\bm{W}}_{0},\beta_{1},\beta_{2},\omega)$ such that Adam achieves the following for all $t>t_{a_{2}}$
$\left|\frac{\min_{i\in[n],c\neq y_{i}}(\bm{e}_{y_{i}}-\bm{e}_{c})^{T}{\bm{W}}_{t}\bm{h}_{i}}{{\left\|{\bm{W}}_{t}\right\|_{\max}}}-\gamma\right|$ $\leq\mathcal{O}(\frac{\sum_{s=t_{a_{2}}}^{t-1}\eta_{s}e^{-\frac{\gamma}{4}\sum_{\tau=t_{a_{2}}}^{s-1}\eta_{\tau}}+\sum_{s=0}^{t_{a_{2}}-1}\eta_{s}+d\sum_{s=t_{a_{2}}}^{t-1}\eta_{s}^{3/2}}{\sum_{s=0}^{t-1}\eta_{s}}).$
Proof.
Determination of $t_{a_{1}}$. Here, we consider learning rate schedule of the form $\eta_{t}=\Theta(\frac{1}{t^{a}})$ where $a\in(0,1]$. We choose $t_{a_{1}}$ after ($\max\{t_{0},t_{A},\log(\frac{1}{\omega})\}$ where $t_{0}$ satisfies Assumption 3.4 and $t_{A}=\frac{2\log(\frac{\sqrt{1-\beta_{2}}}{4})}{\log\beta_{1}}$) such that the following conditions are met: $\alpha_{a_{1}}\beta_{1}^{t/2}\leq\frac{1}{6}$, $\alpha_{a_{2}}d\eta_{t}^{1/2}\leq\frac{1}{6}$, and $\alpha_{a_{3}}d\eta_{t}\leq\frac{1}{6}$. Concretely, we can set $t_{a_{1}}=\max\{\frac{-2\log(6\alpha_{a_{1}})}{\log\beta_{1}},(36\alpha_{a_{2}}^{2}d^{2})^{1/a},(6\alpha_{a_{3}}d)^{1/a}\}=\Theta(d^{2/a})$. Then, we have for all $t\geq t_{a_{1}}$
$\mathcal{L}({\bm{W}}_{t+1})\leq\mathcal{L}({\bm{W}}_{t})-\frac{\eta_{t}\gamma}{2}{\mathcal{G}}({\bm{W}}_{t}).$ (26)
Rearranging this equation and using non-negativity of the loss we obtain $\gamma\sum_{s=t_{a_{1}}}^{t}\eta_{s}{\mathcal{G}}({\bm{W}}_{s})\leq 2\mathcal{L}({\bm{W}}_{t_{a_{1}}})$.
Determination of $t_{a_{2}}$. By Lemma B.2, we can bound $\mathcal{L}({\bm{W}}_{t_{s_{1}}})$ as follows
$|\mathcal{L}({\bm{W}}_{t_{a_{1}}})-\mathcal{L}({\bm{W}}_{0})|\leq 2B{\left\|{\bm{W}}_{t_{a_{1}}}-{\bm{W}}_{0}\right\|_{\max}}\leq 2B\sum_{s=0}^{t_{a_{1}}-1}\eta_{s}{\left\|\frac{\bm{M}_{s}}{\sqrt{{\bm{V}}_{s}}}\right\|_{\max}}\leq 2B\alpha\sum_{s=0}^{t_{a_{1}}-1}\eta_{s},$
where the last inequality is by Lemma D.1. Combining this with the result above and letting ~$\tilde{\mathcal{L}}=\min\{\frac{\log 2}{n},\frac{1}{\sqrt{4B^{2}+\alpha_{V}\eta_{0}}}\}$), we obtain
${\mathcal{G}}({\bm{W}}_{t^{*}})=\min_{s\in[t_{a_{1}},t_{a_{2}}]}{\mathcal{G}}({\bm{W}}_{s})\leq\frac{2\mathcal{L}({\bm{W}}_{0})+4B\alpha\sum_{s=1}^{t_{a_{1}}-1}\eta_{s}}{\gamma\sum_{s=t_{a_{1}}}^{t_{a_{2}}}\eta_{s}}\leq\frac{\tilde{\mathcal{L}}}{2}\leq\frac{1}{2n},$
from which we derive the sufficient condition on $t_{a_{2}}$ to be $\sum_{s=t_{a_{1}}}^{t_{a_{2}}}\eta_{s}\geq\frac{4\mathcal{L}({\bm{W}}_{0})+8B\alpha\sum_{s=1}^{t_{a_{1}}-1}\eta_{s}}{\gamma\tilde{\mathcal{L}}}$.
Convergence of $\frac{{\mathcal{G}}({\bm{W}}_{t})}{\mathcal{L}({\bm{W}}_{t})}$ We follow the same arguments in the proof of SignGD (Theorem C.4) to conclude that
$\frac{{\mathcal{G}}({\bm{W}}_{t})}{\mathcal{L}({\bm{W}}_{t})}\geq 1-e^{-\frac{\gamma}{4}\sum_{s=t_{a_{2}}}^{t-1}\eta_{s}}.$ (27)
We note that $t_{a_{2}}$ satisfies the assumptions in Lemma D.6 and Lemma D.7.
Margin Convergence Finally, we combine Lemma D.6, Lemma D.7, and (27) to obtain
$|\frac{\min_{i\in[n],c\neq y_{i}}(\bm{e}_{y_{i}}-\bm{e}_{c})^{T}{\bm{W}}_{t}\bm{h}_{i}}{{\left\|{\bm{W}}_{t}\right\|_{\max}}}-\gamma|$ $\leq\mathcal{O}(\frac{\sum_{s=t_{a_{2}}}^{t-1}\eta_{s}e^{-\frac{\gamma}{4}\sum_{\tau=t_{a_{2}}}^{s-1}\eta_{\tau}}+\sum_{s=0}^{t_{a_{2}}-1}\eta_{s}+d\sum_{s=t_{a_{2}}}^{t-1}\eta_{s}^{3/2}+d\sum_{s=t_{a_{2}}}^{t-1}\eta_{s}^{2}}{\sum_{s=0}^{t-1}\eta_{s}})$
$\leq\mathcal{O}(\frac{\sum_{s=t_{a_{2}}}^{t-1}\eta_{s}e^{-\frac{\gamma}{4}\sum_{\tau=t_{a_{2}}}^{s-1}\eta_{\tau}}+\sum_{s=0}^{t_{a_{2}}-1}\eta_{s}+d\sum_{s=t_{a_{2}}}^{t-1}\eta_{s}^{3/2}}{\sum_{s=0}^{t-1}\eta_{s}})$
Similar to the case of SignGD, we can derive the margin convergence rates for Adam.
Corollary D.9.
Consider learning rate schedule of the form $\eta_{t}=\Theta(\frac{1}{t^{a}})$ where $a\in(0,1]$, under the same setting as Theorem D.8, then we have for Adam
$|\frac{\min_{i\in[n],c\neq y_{i}}(\bm{e}_{y_{i}}-\bm{e}_{c})^{T}{\bm{W}}_{t}\bm{h}_{i}}{{\left\|{\bm{W}}_{t}\right\|_{\max}}}-\gamma|=\left\{\begin{array}[]{ll}\mathcal{O}(\frac{dt^{1-\frac{3a}{2}}+nd^{\frac{2(1-a)}{a}}+n\mathcal{L}({\bm{W}}_{0})+[\log(1/\omega)]^{1-a}}{t^{1-a}})&\text{if}\quad a<\frac{2}{3}\\ \mathcal{O}(\frac{d\log(t)+nd+n\mathcal{L}({\bm{W}}_{0})+[\log(1/\omega)]^{1/3}}{t^{1/3}})&\text{if}\quad a=\frac{2}{3}\\ \mathcal{O}(\frac{d+nd^{\frac{2(1-a)}{a}}+n\mathcal{L}({\bm{W}}_{0})+[\log(1/\omega)]^{1-a}}{t^{1-a}})&\text{if}\quad\frac{2}{3}<a<1\\ \mathcal{O}(\frac{d+n\log(d)+n\mathcal{L}({\bm{W}}_{0})+\log\log(1/\omega)}{\log t})&\text{if}\quad a=1\end{array}\right.$
Proof.
Recall that $t_{a_{1}}=\Theta(d^{2/a})=C_{a_{1}}d^{2/a}$, and the condition on $t_{a_{2}}$ is $\frac{2\mathcal{L}({\bm{W}}_{0})+4B\alpha\sum_{s=1}^{t_{a_{1}}-1}\eta_{s}}{\gamma\sum_{s=t_{a_{1}}}^{t_{a_{2}}}\eta_{s}}\leq\frac{\tilde{\mathcal{L}}}{2}$, where ~$\tilde{\mathcal{L}}=\min\{\frac{\log 2}{n},\frac{1}{\sqrt{4B^{2}+\alpha_{V}\eta_{0}}}\}$. Then, we apply integral approximations and the rest of the proof can be found in ([)Corollary 4.7 and Lemma C.1]zhang2024implicit. ∎
Lemma D.10.
For any ${\bm{v}},{\bm{v}}^{\prime},{\bm{q}},{\bm{q}}^{\prime}\in\mathbb{R}^{k}$ and $c\in[k]$, the following inequalities hold:
  1. (i)
    $|\frac{\mathbb{S}_{c}({\bm{v}}^{\prime})}{\mathbb{S}_{c}({\bm{v}})}-1|\leq e^{2\lVert{\bm{v}}-{\bm{v}}^{\prime}\rVert_{\infty}}-1$
  2. (ii)
    $|\frac{1-\mathbb{S}_{c}({\bm{v}}^{\prime})}{1-\mathbb{S}_{c}({\bm{v}})}-1|\leq e^{2\lVert{\bm{v}}-{\bm{v}}^{\prime}\rVert_{\infty}}-1$
  3. (iii)
    $|\frac{\mathbb{S}_{c}({\bm{v}}^{\prime})\mathbb{S}_{c}({\bm{q}}^{\prime})}{\mathbb{S}_{c}({\bm{v}})\mathbb{S}_{c}({\bm{q}})}-1|\leq e^{2(\lVert{\bm{v}}^{\prime}-{\bm{v}}\rVert_{\infty}+\lVert{\bm{q}}^{\prime}-{\bm{q}}\rVert_{\infty})}-1$
  4. (iv)
    $|\frac{\mathbb{S}_{c}({\bm{v}}^{\prime})(1-\mathbb{S}_{c}({\bm{q}}^{\prime}))}{\mathbb{S}_{c}({\bm{v}})(1-\mathbb{S}_{c}({\bm{q}}))}-1|\leq e^{2(\lVert{\bm{v}}^{\prime}-{\bm{v}}\rVert_{\infty}+\lVert{\bm{q}}^{\prime}-{\bm{q}}\rVert_{\infty})}-1$
  5. (v)
    $|\frac{(1-\mathbb{S}_{c}({\bm{v}}^{\prime}))(1-\mathbb{S}_{c}({\bm{q}}^{\prime}))}{(1-\mathbb{S}_{c}({\bm{v}}))(1-\mathbb{S}_{c}({\bm{q}}))}-1|\leq e^{2(\lVert{\bm{v}}^{\prime}-{\bm{v}}\rVert_{\infty}+\lVert{\bm{q}}^{\prime}-{\bm{q}}\rVert_{\infty})}-1$
Proof.
We prove each inequality:
(i) First, observe that
$|\frac{\mathbb{S}_{c}({\bm{v}}^{\prime})}{\mathbb{S}_{c}({\bm{v}})}-1|$ $=|\frac{e^{v^{\prime}_{c}}}{e^{v_{c}}}\frac{\sum_{i\in[k]}e^{v_{i}}}{\sum_{i\in[k]}e^{v^{\prime}_{i}}}-1|$
$=|\frac{\sum_{i\in[k]}e^{v^{\prime}_{c}+v_{i}}-\sum_{i\in[k]}e^{v_{c}+v^{\prime}_{i}}}{\sum_{i\in[k]}e^{v_{c}+v^{\prime}_{i}}}|$
$\leq\frac{\sum_{i\in[k]}|e^{v^{\prime}_{c}+v_{i}}-e^{v_{c}+v^{\prime}_{i}}|}{\sum_{i\in[k]}e^{v_{c}+v^{\prime}_{i}}}$
For any $i\in[k]$, we have $\frac{|e^{v^{\prime}_{c}+v_{i}}-e^{v_{c}+v^{\prime}_{i}}|}{e^{v_{c}+v^{\prime}_{i}}}=|e^{v^{\prime}_{c}-v_{c}+v_{i}-v^{\prime}_{i}}-1|\leq e^{|v^{\prime}_{c}-v_{c}+v_{i}-v^{\prime}_{i}|}-1\leq e^{2\lVert{\bm{v}}-{\bm{v}}^{\prime}\rVert_{\infty}}-1$. This implies $\sum_{i\in[k]}|e^{v^{\prime}_{c}+v_{i}}-e^{v_{c}+v^{\prime}_{i}}|\leq\bigl{(}e^{2\lVert{\bm{v}}-{\bm{v}}^{\prime}\rVert_{\infty}}-1\bigr{)}\sum_{i\in[k]}e^{v_{c}+v^{\prime}_{i}}$, from which we obtain the desired inequality.
(ii) For the second inequality:
$|\frac{1-\mathbb{S}_{c}({\bm{v}}^{\prime})}{1-\mathbb{S}_{c}({\bm{v}})}-1|$ $=|\frac{1-\frac{e^{v^{\prime}_{c}}}{\sum_{i\in[k]}e^{v^{\prime}_{i}}}}{1-\frac{e^{v_{c}}}{\sum_{i\in[k]}e^{v_{i}}}}-1|$
$=|\frac{(\sum_{j\in[k],j\neq c}e^{v^{\prime}_{j}})(\sum_{i\in[k]}e^{v_{i}})}{(\sum_{j\in[k],j\neq c}e^{v_{j}})(\sum_{i\in[k]}e^{v^{\prime}_{i}})}-1|$
$=|\frac{\sum_{j\in[k],j\neq c}\sum_{i\in[k]}\bigl{[}e^{v^{\prime}_{j}+v_{i}}-e^{v_{j}+v^{\prime}_{i}}\bigl{]}}{\sum_{j\in[k],j\neq c}\sum_{i\in[k]}e^{v_{j}+v^{\prime}_{i}}}|$
$\leq\frac{\sum_{j\in[k],j\neq c}\sum_{i\in[k]}|e^{v^{\prime}_{j}+v_{i}}-e^{v_{j}+v^{\prime}_{i}}|}{\sum_{j\in[k],j\neq c}\sum_{i\in[k]}e^{v_{j}+v^{\prime}_{i}}}$
For any $j\in[k]$, $j\neq c$, and $i\in[k]$, we have $\frac{|e^{v_{j}^{\prime}+v_{i}}-e^{v_{j}+v_{i}^{\prime}}|}{e^{v_{j}+v_{i}^{\prime}}}\leq e^{2\lVert{\bm{v}}-{\bm{v}}^{\prime}\rVert_{\infty}}-1$. This implies that $\sum_{j\in[k],j\neq c}\sum_{i\in[k]}|e^{v_{j}^{\prime}+v_{i}}-e^{v_{j}+v_{i}^{\prime}}|\leq(e^{2\lVert{\bm{v}}-{\bm{v}}^{\prime}\rVert_{\infty}}-1)\sum_{j\in[k],j\neq c}\sum_{i\in[k]}e^{v_{j}+v_{i}^{\prime}}$, from which the result follows.
(iii) For the third inequality:
$|\frac{\mathbb{S}_{c}({\bm{v}}^{\prime})\mathbb{S}_{c}({\bm{q}}^{\prime})}{\mathbb{S}_{c}({\bm{v}})\mathbb{S}_{c}({\bm{q}})}-1|$ $=|\frac{\frac{e^{v^{\prime}_{c}}}{\sum_{i\in[k]}e^{v^{\prime}_{i}}}\frac{e^{q^{\prime}_{c}}}{\sum_{i\in[k]}e^{q^{\prime}_{i}}}}{\frac{e^{v_{c}}}{\sum_{i\in[k]}e^{v_{i}}}\frac{e^{q_{c}}}{\sum_{i\in[k]}e^{q_{i}}}}-1|$
$=|\frac{\frac{e^{v^{\prime}_{c}}}{\sum_{i\in[k]}e^{v^{\prime}_{i}}}\frac{e^{q^{\prime}_{c}}}{\sum_{i\in[k]}e^{q^{\prime}_{i}}}}{\frac{e^{v_{c}}}{\sum_{i\in[k]}e^{v_{i}}}\frac{e^{q_{c}}}{\sum_{i\in[k]}e^{q^{\prime}_{i}}}}-\frac{\frac{e^{v_{c}}}{\sum_{i\in[k]}e^{v^{\prime}_{i}}}\frac{e^{q_{c}}}{\sum_{i\in[k]}e^{q^{\prime}_{i}}}}{\frac{e^{v_{c}}}{\sum_{i\in[k]}e^{v^{\prime}_{i}}}\frac{e^{q_{c}}}{\sum_{i\in[k]}e^{q^{\prime}_{i}}}}|$
$=|\frac{e^{v^{\prime}_{c}}e^{q^{\prime}_{c}}\sum_{i\in[k]}e^{v_{i}}\sum_{j\in[k]}e^{q_{j}}}{e^{v_{c}}e^{q_{c}}\sum_{i\in[k]}e^{v^{\prime}_{i}}\sum_{j\in[k]}e^{q^{\prime}_{j}}}-\frac{e^{v_{c}}e^{q_{c}}\sum_{i\in[k]}e^{v^{\prime}_{i}}\sum_{j\in[k]}e^{q^{\prime}_{j}}}{e^{v_{c}}e^{q_{c}}\sum_{i\in[k]}e^{v^{\prime}_{i}}\sum_{j\in[k]}e^{q^{\prime}_{j}}}|$
$=|\frac{\sum_{i\in[k]}\sum_{j\in[k]}\bigl{[}e^{v^{\prime}_{c}+v_{i}+q^{\prime}_{c}+q_{j}}-e^{v_{c}+v^{\prime}_{i}+q_{c}+q^{\prime}_{j}}\bigr{]}}{\sum_{i\in[k]}\sum_{j\in[k]}e^{v_{c}+v^{\prime}_{i}+q_{c}+q^{\prime}_{j}}}|$
$\leq\frac{\sum_{i\in[k]}\sum_{j\in[k]}|e^{v^{\prime}_{c}+v_{i}+q^{\prime}_{c}+q_{j}}-e^{v_{c}+v^{\prime}_{i}+q_{c}+q^{\prime}_{j}}|}{\sum_{i\in[k]}\sum_{j\in[k]}e^{v_{c}+v^{\prime}_{i}+q_{c}+q^{\prime}_{j}}}$
For any $i\in[k]$ and $j\in[k]$, $\frac{|e^{v^{\prime}_{c}+v_{i}+q^{\prime}_{c}+q_{j}}-e^{v_{c}+v^{\prime}_{i}+q_{c}+q^{\prime}_{j}}|}{e^{v_{c}+v^{\prime}_{i}+q_{c}+q^{\prime}_{j}}}=|e^{v^{\prime}_{c}-v_{c}+v_{i}-v^{\prime}_{i}+q^{\prime}_{c}-q_{c}+q_{j}-q^{\prime}_{j}}-1|\leq e^{|v^{\prime}_{c}-v_{c}|+|v_{i}-v^{\prime}_{i}|+|q^{\prime}_{c}-q_{c}|+|q_{j}-q^{\prime}_{j}|}-1\leq e^{2(\lVert{\bm{v}}^{\prime}-{\bm{v}}\rVert_{\infty}+\lVert{\bm{q}}^{\prime}-{\bm{q}}\rVert_{\infty})}-1$. Then, rearranging and summing over $i$ and $j$ leads to the result.
(iv) For the fourth inequality:
$|\frac{\mathbb{S}_{c}({\bm{v}}^{\prime})(1-\mathbb{S}_{c}({\bm{q}}^{\prime}))}{\mathbb{S}_{c}({\bm{v}})(1-\mathbb{S}_{c}({\bm{q}}))}-1|$ $=|\frac{\frac{e^{v^{\prime}_{c}}}{\sum_{s\in[k]}e^{v^{\prime}_{s}}}(1-\frac{e^{q^{\prime}_{c}}}{\sum_{t\in[k]}e^{q^{\prime}_{t}}})}{\frac{e^{v_{c}}}{\sum_{s\in[k]}e^{v_{s}}}(1-\frac{e^{q_{c}}}{\sum_{t\in[k]}e^{q_{t}}})}-1|$
$=|\frac{\frac{e^{v^{\prime}_{c}}}{\sum_{s\in[k]}e^{v^{\prime}_{s}}}\frac{\sum_{i\in[k],i\neq c}e^{q^{\prime}_{i}}}{\sum_{t\in[k]}e^{q^{\prime}_{t}}}}{\frac{e^{v_{c}}}{\sum_{s\in[k]}e^{v_{s}}}\frac{\sum_{i\in[k],i\neq c}e^{q_{t}}}{\sum_{t\in[k]}e^{q_{t}}}}-1|$
$=|\frac{\sum_{i\in[k],i\neq c}\sum_{t\in[k]}\sum_{s\in[k]}e^{v^{\prime}_{c}+q^{\prime}_{i}+v_{s}+q_{t}}}{\sum_{i\in[k],i\neq c}\sum_{t\in[k]}\sum_{s\in[k]}e^{v_{c}+q_{i}+v^{\prime}_{s}+q^{\prime}_{t}}}-1|$
$\leq\frac{\sum_{i\in[k],i\neq c}\sum_{t\in[k]}\sum_{s\in[k]}|e^{v^{\prime}_{c}+q^{\prime}_{i}+v_{s}+q_{t}}-e^{v_{c}+q_{i}+v^{\prime}_{s}+q^{\prime}_{t}}|}{\sum_{i\in[k],i\neq c}\sum_{t\in[k]}\sum_{s\in[k]}e^{v_{c}+q_{i}+v^{\prime}_{s}+q^{\prime}_{t}}}$
For each $i\in[k],i\neq c$, $s\in[k]$, and $t\in[k]$, we obtain $\frac{|e^{v^{\prime}_{c}+q^{\prime}_{i}+v_{s}+q_{t}}-e^{v_{c}+q_{i}+v^{\prime}_{s}+q^{\prime}_{t}}|}{e^{v_{c}+q_{i}+v^{\prime}_{s}+q^{\prime}_{t}}}\leq e^{2(\lVert{\bm{v}}^{\prime}-{\bm{v}}\rVert_{\infty}+\lVert{\bm{q}}^{\prime}-{\bm{q}}\rVert_{\infty})}-1$. Then, rearranging and summing over $i$, $s$, and $t$ leads to the result.
(v) Finally, for the fifth inequality:
$|\frac{(1-\mathbb{S}_{c}({\bm{v}}^{\prime}))(1-\mathbb{S}_{c}({\bm{q}}^{\prime}))}{(1-\mathbb{S}_{c}({\bm{v}}))(1-\mathbb{S}_{c}({\bm{q}}))}-1|$ $=|\frac{(1-\frac{e^{v^{\prime}_{c}}}{\sum_{s\in[k]}e^{v^{\prime}_{s}}})(1-\frac{e^{q^{\prime}_{c}}}{\sum_{t\in[k]}e^{q^{\prime}_{t}}})}{(1-\frac{e^{v_{c}}}{\sum_{s\in[k]}e^{v_{s}}})(1-\frac{e^{q_{c}}}{\sum_{t\in[k]}e^{q_{t}}})}-1|$
$=|\frac{\frac{\sum_{j\in[k],j\neq c}e^{v^{\prime}_{j}}}{\sum_{s\in[k]}e^{v^{\prime}_{s}}}\frac{\sum_{i\in[k],i\neq c}e^{q^{\prime}_{i}}}{\sum_{t\in[k]}e^{q^{\prime}_{t}}}}{\frac{\sum_{j\in[k],j\neq c}e^{v_{j}}}{\sum_{s\in[k]}e^{v_{s}}}\frac{\sum_{i\in[k],i\neq c}e^{q_{i}}}{\sum_{t\in[k]}e^{q_{t}}}}-1|$
$=|\frac{\sum_{j\in[k],j\neq c}\sum_{i\in[k],i\neq c}\sum_{t\in[k]}\sum_{s\in[k]}e^{v^{\prime}_{j}+q^{\prime}_{i}+v_{s}+q_{t}}}{\sum_{j\in[k],j\neq c}\sum_{i\in[k],i\neq c}\sum_{t\in[k]}\sum_{s\in[k]}e^{v_{j}+q_{i}+v^{\prime}_{s}+q^{\prime}_{t}}}-1|$
$\leq\frac{\sum_{j\in[k],j\neq c}\sum_{i\in[k],i\neq c}\sum_{t\in[k]}\sum_{s\in[k]}|e^{v^{\prime}_{j}+q^{\prime}_{i}+v_{s}+q_{t}}-e^{v_{j}+q_{i}+v^{\prime}_{s}+q^{\prime}_{t}}|}{\sum_{j\in[k],j\neq c}\sum_{i\in[k],i\neq c}\sum_{t\in[k]}\sum_{s\in[k]}e^{v_{j}+q_{i}+v^{\prime}_{s}+q^{\prime}_{t}}}.$
For each $j\in[k]$ ($j\neq c$), $i\in[k]$ ($i\neq c$), $s\in[k]$, and $t\in[k]$, we have
$\frac{|e^{v^{\prime}_{j}+q^{\prime}_{i}+v_{s}+q_{t}}-e^{v_{j}+q_{i}+v^{\prime}_{s}+q^{\prime}_{t}}|}{e^{v_{j}+q_{i}+v^{\prime}_{s}+q^{\prime}_{t}}}$ $=|e^{v^{\prime}_{j}-v_{j}+q^{\prime}_{i}-q_{i}+v_{s}-v^{\prime}_{s}+q_{t}-q^{\prime}_{t}}-1|$
$\leq e^{|v^{\prime}_{j}-v_{j}|+|q^{\prime}_{i}-q_{i}|+|v_{s}-v^{\prime}_{s}|+|q_{t}-q^{\prime}_{t}|}-1$
$\leq e^{2(\lVert{\bm{v}}^{\prime}-{\bm{v}}\rVert_{\infty}+\lVert{\bm{q}}^{\prime}-{\bm{q}}\rVert_{\infty})}-1$
Then, rearranging and summing over $j$, $i$, $s$, and $t$ leads to the result. ∎
Appendix E Normalized $p$-Nrom Steepest Descent
[…]
Consider normalized steepest-descent with respect to the $p$-norm:
${\bm{W}}_{t+1}={\bm{W}}_{t}-\eta_{t}\bm{\Delta}_{t},\qquad\text{where}~{}~{}\bm{\Delta}_{t}:=\arg\max_{\|\bm{\Delta}\|\leq 1}\langle\nabla\mathcal{L}({\bm{W}}_{t}),\bm{\Delta}\rangle\,.$ (28)
Lemma E.1 generalizes Lemma B.1.
Lemma E.1 (${\mathcal{G}}({\bm{W}})$ as proxy to the loss-gradient norm).
Define the margin $\gamma$ with respect to $p$-norm, for $p\geq 1$ and denote $\|\cdot\|_{\star}$ its dual norm. Then, for any ${\bm{W}}$ it holds that
$2B\cdot{\mathcal{G}}({\bm{W}})\geq\|{\nabla\mathcal{L}({\bm{W}})}\|_{\star}\geq\gamma\cdot{\mathcal{G}}({\bm{W}})\,.$
Proof.
First, we prove the lower bound. By duality and direct application of (12)
$\|{\nabla\mathcal{L}({\bm{W}})}\|_{\star}$ $=\max_{\|{\bm{A}}\|\leq 1}\langle\bm{A},-\nabla\mathcal{L}({\bm{W}})\rangle$
$\geq\max_{\|\bm{A}\|\leq 1}\frac{1}{n}\sum_{i\in[n]}(1-s_{iy_{i}})\min_{c\neq y_{i}}(\bm{e}_{y_{i}}-\bm{e}_{c})^{T}\bm{A}\bm{h}_{i}$
$\geq\frac{1}{n}\sum_{i\in[n]}(1-s_{iy_{i}})\cdot\max_{\|\bm{A}\|\leq 1}\min_{i\in[n],c\neq y_{i}}(\bm{e}_{y_{i}}-\bm{e}_{c})^{T}\bm{A}\bm{h}_{i}.$
The upper bound follows because for $p$-norms with $p\geq 1$
$\|\nabla\mathcal{L}({\bm{W}})\|_{\star}\leq{\left\|\nabla\mathcal{L}({\bm{W}})\right\|_{\rm{sum}}}$
and we can use the bound for ${\left\|\nabla\mathcal{L}({\bm{W}})\right\|_{\rm{sum}}}$. ∎
A direct consequence of Lemma E.1 is the following.
Lemma E.2.
For any ${\bm{W}},{\bm{W}}_{0}\in\mathbb{R}^{k\times d}$, suppose that $\mathcal{L}({\bm{W}})$ is convex, we have
$|\mathcal{L}({\bm{W}})-\mathcal{L}({\bm{W}}_{0})|\leq 2B\lVert{\bm{W}}-{\bm{W}}_{0}\rVert.$
Proof.
We replace the term ${\left\|\nabla\mathcal{L}({\bm{W}}_{0})\right\|_{\rm{sum}}}{\left\|{\bm{W}}_{0}-{\bm{W}}\right\|_{\max}}$ in Lemma B.2 with $\lVert\nabla\mathcal{L}({\bm{W}}_{0})\rVert_{*}\lVert{\bm{W}}_{0}-{\bm{W}}\rVert$. The rest proof follows the same steps as Lemma B.2. ∎
Lemma E.3 generalizes Lemma A.3.
Lemma E.3.
For any $\bm{s}\in\Delta^{k-1}$ in the $k$-dimensional simplex, any index $c\in[k]$, and any ${\bm{v}}\in\mathbb{R}^{k}$ it holds:
${\bm{v}}^{\top}\left(\operatorname{diag}(\bm{s})-\bm{s}\bm{s}^{\top}\right){\bm{v}}\leq 4\,\|{\bm{v}}\|^{2}\,(1-s_{c})$
Proof.
By Cauchy-Schwartz,
${\bm{v}}^{\top}\left(\operatorname{diag}(\bm{s})-\bm{s}\bm{s}^{\top}\right){\bm{v}}$ $=\operatorname{vec}\big{(}\operatorname{diag}(\bm{s})-\bm{s}\bm{s}^{\top}\big{)}^{\top}\operatorname{vec}\big{(}{\bm{v}}{\bm{v}}^{\top}\big{)}$
$\leq\|\operatorname{vec}\big{(}\operatorname{diag}(\bm{s})-\bm{s}\bm{s}^{\top}\big{)}\|_{\star}\|\operatorname{vec}\big{(}{\bm{v}}{\bm{v}}^{\top}\big{)}\|$
$\leq\|{\bm{v}}\|^{2}\,\|\operatorname{vec}\big{(}\operatorname{diag}(\bm{s})-\bm{s}\bm{s}^{\top}\big{)}\|_{*}\,.$
But,
$\|\operatorname{vec}\big{(}\operatorname{diag}(\bm{s})-\bm{s}\bm{s}^{\top}\big{)}\|_{\star}\leq\|\operatorname{vec}\big{(}\operatorname{diag}(\bm{s})-\bm{s}\bm{s}^{\top}\big{)}\|_{1}$
and we can use Lemma A.3. ∎
When applying the above lemma to bound the Hessian term we also need to use the following:
$\|\bm{\Delta}\bm{h}\|\leq\|\bm{\Delta}\|\|\bm{h}\|_{\star},$ (29)
which is true because for $q=p/(1-p)$:
$\|\bm{\Delta}\bm{h}\|^{p}=\|\bm{\Delta}\bm{h}\|_{p}^{p}=\sum_{j}|\bm{e}_{j}^{\top}\bm{\Delta}\bm{h}|_{p}^{p}\leq\sum_{j}\|\bm{e}_{j}^{\top}\bm{\Delta}\|_{p}^{p}\|\bm{h}\|_{q}^{p}=\|\bm{h}\|_{q}^{p}\sum_{ij}|\bm{\Delta}[i,j]|^{p}=\|\bm{h}\|_{q}^{p}\|\bm{\Delta}\|_{p}^{p}$ (30)
Appendix F Other multiclass loss functions
F.1 Exponential Loss
The multiclass exponential loss is given as
$\mathcal{L}_{\rm{exp}}({\bm{W}}):=\frac{1}{n}\sum_{i\in[n]}\sum_{c\neq y_{i}}\exp\left(-(\bm{e}_{y_{i}}-\bm{e}_{c})^{\top}{\bm{W}}\bm{h}_{i}\right)\,.$
The gradient of $\mathcal{L}_{\exp}({\bm{W}})$ is
$\nabla\mathcal{L}_{\exp}({\bm{W}})=\frac{1}{n}\sum_{i\in[n]}\sum_{c\neq y_{i}}-\exp(-(\bm{e}_{y_{i}}-\bm{e}_{c})^{T}{\bm{W}}\bm{h}_{i})(\bm{e}_{y_{i}}-\bm{e}_{c})\bm{h}_{i}^{T}.$
Thus, for any matrix $\bm{A}\in\mathbb{R}^{k\times d}$, we have
$\langle{\bm{A}},-\nabla\mathcal{L}_{\exp}({\bm{W}})\rangle=\frac{1}{n}\sum_{i\in[n]}\sum_{c\neq y_{i}}\exp\left(-(\bm{e}_{y_{i}}-\bm{e}_{c})^{\top}{\bm{W}}\bm{h}_{i}\right)\cdot\left(\bm{e}_{y_{i}}-\bm{e}_{c}\right)^{\top}{\bm{A}}\bm{h}_{i}\,.$
This motivates us to define ${\mathcal{G}}({\bm{W}})$ as
${\mathcal{G}}_{\exp}({\bm{W}})=\frac{1}{n}\sum_{i\in[n]}\sum_{c\neq y_{i}}\exp\left(-(\bm{e}_{y_{i}}-\bm{e}_{c})^{\top}{\bm{W}}\bm{h}_{i}\right),$
from which we recognize that ${\mathcal{G}}_{\exp}({\bm{W}})=\mathcal{L}_{\exp}({\bm{W}})$ and the rest follows.
F.2 PairLogLoss
The PairLogLoss loss (Wang et al., 2021b) is given as
$\mathcal{L}_{\rm{pll}}({\bm{W}}):=\frac{1}{n}\sum_{i\in[n]}\sum_{c\neq y_{i}}\log\left(1+\exp\left(-(\bm{e}_{y_{i}}-\bm{e}_{c})^{\top}{\bm{W}}\bm{h}_{i}\right)\right)\,.$
Note that $\mathcal{L}=\frac{1}{n}\sum_{i\in[n]}\sum_{c\neq y_{i}}f\left((\bm{e}_{y_{i}}-\bm{e}_{c})^{\top}{\bm{W}}\bm{h}_{i}\right)$ where $f(t):=\log(1+e^{-t})$ denotes the logistic loss. Therefore, the Taylor expansion of PLL writes:
$\mathcal{L}_{\rm{pll}}({\bm{W}}+\bm{\Delta})$ $=\mathcal{L}({\bm{W}})+\frac{1}{n}\sum_{i\in[n]}\sum_{c\neq y_{i}}f^{\prime}\left((\bm{e}_{y_{i}}-\bm{e}_{c})^{\top}{\bm{W}}\bm{h}_{i}\right)\cdot(\bm{e}_{y_{i}}-\bm{e}_{c})^{\top}\bm{\Delta}\bm{h}_{i}$
$\qquad+\frac{1}{n}\sum_{i\in[n]}\sum_{c\neq y_{i}}f^{\prime\prime}\left((\bm{e}_{y_{i}}-\bm{e}_{c})^{\top}{\bm{W}}\bm{h}_{i}\right)\cdot\bm{h}_{i}^{\top}\bm{\Delta}^{\top}(\bm{e}_{y_{i}}-\bm{e}_{c})(\bm{e}_{y_{i}}-\bm{e}_{c})^{\top}\bm{\Delta}\bm{h}_{i}+o\left(\|\bm{\Delta}\|^{3}\right)\,.$ (31)
From the above, the gradient of the PLL loss is:
$\nabla\mathcal{L}_{\rm{pll}}({\bm{W}})$ $=\frac{1}{n}\sum_{i\in[n]}\sum_{c\neq y_{i}}f^{\prime}\left((\bm{e}_{y_{i}}-\bm{e}_{c})^{\top}{\bm{W}}\bm{h}_{i}\right)\cdot\left(\bm{e}_{y_{i}}-\bm{e}_{c}\right)\bm{h}_{i}^{\top}$
$=\frac{1}{n}\sum_{i\in[n]}\sum_{c\neq y_{i}}\frac{-\exp\left(-(\bm{e}_{y_{i}}-\bm{e}_{c})^{\top}{\bm{W}}\bm{h}_{i}\right)}{1+\exp\left(-(\bm{e}_{y_{i}}-\bm{e}_{c})^{\top}{\bm{W}}\bm{h}_{i}\right)}\left(\bm{e}_{y_{i}}-\bm{e}_{c}\right)\bm{h}_{i}^{\top}$ (32)
Thus, for any matrix ${\bm{A}}\in\mathbb{R}^{k\times d}$,
$\langle{\bm{A}},-\nabla\mathcal{L}_{\rm{pll}}({\bm{W}})\rangle=\frac{1}{n}\sum_{i\in[n]}\sum_{c\neq y_{i}}|f^{\prime}\left((\bm{e}_{y_{i}}-\bm{e}_{c})^{\top}{\bm{W}}\bm{h}_{i}\right)|\cdot\left(\bm{e}_{y_{i}}-\bm{e}_{c}\right)^{\top}{\bm{A}}\bm{h}_{i}\,.$ (33)
This motivates us to define
${\mathcal{G}}_{\rm{pll}}({\bm{W}})=\frac{1}{n}\sum_{i\in[n]}\sum_{c\neq y_{i}}\left|f^{\prime}\left(-(\bm{e}_{y_{i}}-\bm{e}_{c})^{\top}{\bm{W}}\bm{h}_{i}\right)\right|=\frac{1}{n}\sum_{i\in[n]}\sum_{c\neq y_{i}}\frac{\exp\left(-(\bm{e}_{y_{i}}-\bm{e}_{c})^{\top}{\bm{W}}\bm{h}_{i}\right)}{1+\exp\left(-(\bm{e}_{y_{i}}-\bm{e}_{c})^{\top}{\bm{W}}\bm{h}_{i}\right)}$ (34)
Lemma F.1 (Analogue of Lemma B.1 for PLL).
For any ${\bm{W}}$, the PairLogLoss (PLL) satisfies:
$2B\cdot{\mathcal{G}}_{\rm{pll}}({\bm{W}})\geq{\left\|\nabla\mathcal{L}_{\rm{pll}}({\bm{W}})\right\|_{\rm{sum}}}\geq\gamma\cdot{\mathcal{G}}_{\rm{pll}}({\bm{W}})\,.$
Proof.
The lower bound follows immediately from (33) and expressing ${\left\|\nabla\mathcal{L}_{\rm{pll}}({\bm{W}})\right\|_{\rm{sum}}}=\max_{{\left\|{\bm{A}}\right\|_{\max}}\leq 1}\langle{\bm{A}},-\nabla\mathcal{L}_{\rm{pll}}({\bm{W}})\rangle$. The lower bound follows from triangle inequality applied to (32):
${\left\|\nabla\mathcal{L}_{\rm{pll}}({\bm{W}})\right\|_{\rm{sum}}}\leq\frac{1}{n}\sum_{i\in[n]}\sum_{c\neq y_{i}}\left|f^{\prime}\left(-(\bm{e}_{y_{i}}-\bm{e}_{c})^{\top}{\bm{W}}\bm{h}_{i}\right)\right|\|\bm{e}_{y_{i}}-\bm{e}_{c}\|_{1}\|\bm{h}_{i}\|_{1}\leq 2B\cdot{\mathcal{G}}({\bm{W}})\,.$
For bounding with ${\mathcal{G}}({\bm{W}})$ the second-order term in the Taylor expansion of PLL, note the following. First, for all $i\in[n],c\neq y_{i}$:
$\bm{h}_{i}^{\top}\bm{\Delta}^{\top}(\bm{e}_{y_{i}}-\bm{e}_{c})(\bm{e}_{y_{i}}-\bm{e}_{c})^{\top}\bm{\Delta}\bm{h}_{i}$ $=\langle(\bm{e}_{y_{i}}-\bm{e}_{c})(\bm{e}_{y_{i}}-\bm{e}_{c})^{\top},\bm{\Delta}\bm{h}_{i}\bm{h}_{i}^{\top}\bm{\Delta}^{T}\rangle\leq{\left\|(\bm{e}_{y_{i}}-\bm{e}_{c})(\bm{e}_{y_{i}}-\bm{e}_{c})^{\top}\right\|_{\rm{sum}}}{\left\|\bm{\Delta}\bm{h}_{i}\bm{h}_{i}^{\top}\bm{\Delta}^{T}\right\|_{\max}}$
$\leq\|\bm{e}_{y_{i}}-\bm{e}_{c})\|_{1}^{2}\cdot\|\bm{\Delta}\bm{h}_{i}\|_{\infty}^{2}$
$\leq 4\cdot\left({\left\|\bm{\Delta}\right\|_{\max}}\right)^{2}\cdot\|\bm{h}_{i}\|_{1}^{2}\leq 4B^{2}\left({\left\|\bm{\Delta}\right\|_{\max}}\right)^{2}\,.$
Second, the (easy to check) property of logistic loss that $f^{\prime\prime}(t)\leq|f^{\prime}(t)|$. Putting these together:
$\frac{1}{n}\sum_{i\in[n]}\sum_{c\neq y_{i}}f^{\prime\prime}\left((\bm{e}_{y_{i}}-\bm{e}_{c})^{\top}{\bm{W}}\bm{h}_{i}\right)\cdot\bm{h}_{i}^{\top}\bm{\Delta}^{\top}(\bm{e}_{y_{i}}-\bm{e}_{c})(\bm{e}_{y_{i}}-\bm{e}_{c})^{\top}\bm{\Delta}\bm{h}_{i}\leq 4B^{2}\cdot{\mathcal{G}}({\bm{W}})\cdot\left({\left\|\bm{\Delta}\right\|_{\max}}\right)^{2}\,.$
Finally, we verify PLL satisfies Lemma B.3.
Lemma F.2 (Analogue of Lemma B.3 for PLL).
Let ${\bm{W}}\in\mathbb{R}^{k\times d}$, we have
  1. (i)
    $1\geq\frac{{\mathcal{G}}_{\rm{pll}}({\bm{W}})}{\mathcal{L}_{\rm{pll}}({\bm{W}})}\geq 1-\frac{n\mathcal{L}_{\rm{pll}}({\bm{W}})}{2}$
  2. (ii)
    Suppose that ${\bm{W}}$ satisfies $\mathcal{L}_{\rm{pll}}({\bm{W}})\leq\frac{\log 2}{n}$ or ${\mathcal{G}}_{\rm{pll}}({\bm{W}})\leq\frac{1}{2n}$, then $\mathcal{L}_{\rm{pll}}({\bm{W}})\leq 2{\mathcal{G}}_{\rm{pll}}({\bm{W}}).$
Proof.
(i) The upper bound follows by the well-known self-boundedness property of the logistic loss, namely $|f^{\prime}(t)|\leq f(t)$
To prove the upper bound, it suffices to prove for for $x>0$:
$\frac{x}{1+x}\geq\log(1+x)-\frac{1}{2}\log^{2}(1+x).$ (35)
The general case follows by summing over $x_{ic}=\exp\left(-(\bm{e}_{y_{i}}-\bm{e}_{c})^{\top}{\bm{W}}\bm{h}_{i}\right),i\in[n],c\neq y_{i}$ since then we have
${\mathcal{G}}({\bm{W}})=\sum_{i\in[n]}\sum_{c\neq y_{i}}\frac{x_{ic}}{1+x_{ic}}$ $\geq\sum_{i\in[n]}\sum_{c\neq y_{i}}\log(1+x_{ic})-\frac{1}{2}\sum_{i\in[n]}\sum_{c\neq y_{i}}\log^{2}(1+x_{ic})$
$\geq\sum_{i\in[n]}\sum_{c\neq y_{i}}\log(1+x_{ic})-\frac{1}{2}\left(\sum_{i\in[n]}\sum_{c\neq y_{i}}\log(1+x_{ic})\right)^{2}\,,$
where the last line used $\log(1+x_{ic})\geq 0$. For (15), let $a=\log(1+x)>0$. The inequality becomes $e^{-a}\leq 1-a+a^{2}/2$, which holds for $a>0$ by the second-order Taylor expansion of $e^{-a}$ around 00.
(ii) Denote $\mathcal{L}\coloneqq\mathcal{L}_{pll}$ and ${\mathcal{G}}\coloneqq{\mathcal{G}}_{pll}$. Given $\mathcal{L}\leq\frac{\log(2)}{n}\leq\frac{1}{n}$, we have $1-\frac{n\mathcal{L}}{2}\geq\frac{1}{2}$, then the first part follows from (i). For the second part, denote $l_{ic}:=(\bm{e}_{y_{i}}-\bm{e}_{c})^{\top}{\bm{W}}\bm{h}_{i},i\in[n],c\neq y_{i}$. For $\mathcal{L}\leq 2{\mathcal{G}}$ to hold, it is sufficient to show that $\log(1+e^{-l_{ic}})\leq 2\frac{e^{-l_{ic}}}{1+e^{-l_{ic}}}$ for all $i\in[n],c\neq y_{i}$. This holds true when $l_{ic}\geq-1.366$, which is clearly satisfied given the assumption ${\mathcal{G}}\leq\frac{1}{2n}$ implying $l_{ic}\geq 0$. ∎
Lemma F.3 (Analogue of Lemma B.5 for PLL).
For any $\psi\in[0,1]$, we have the following:
$\frac{{\mathcal{G}}_{pll}({\bm{W}}+\psi\triangle{\bm{W}})}{{\mathcal{G}}_{pll}({\bm{W}})}\leq e^{2B\psi{\left\|\triangle{\bm{W}}\right\|_{\max}}}+2$
Proof.
For logistic loss $f(z)=\log(1+e^{-z})$, for any $z_{1},z_{2}\in\mathbb{R}$, we have the following
$\bigm{|}\frac{f^{\prime}(z_{1})}{f^{\prime}(z_{2})}\bigm{|}=\bigm{|}\frac{1+e^{z_{2}}}{1+e^{z_{1}}}\bigm{|}$ $=\bigm{|}\frac{1+e^{z_{2}}-e^{z_{1}}+e^{z_{1}}}{1+e^{z_{1}}}\bigm{|}$
$=\bigm{|}\frac{e^{z_{2}}-e^{z_{1}}}{1+e^{z_{1}}}+1\bigm{|}\leq\bigm{|}\frac{e^{z_{2}}-e^{z_{1}}}{1+e^{z_{1}}}\bigm{|}+1$
$\leq\bigm{|}e^{z_{2}-z_{1}}-1\bigm{|}+1$
$\leq e^{|z_{2}-z_{1}|}+2.$
Denote $x_{ic}^{{\bm{W}}}:=(\bm{e}_{y_{i}}-\bm{e}_{c})^{T}{\bm{W}}\bm{h}_{i}$ and $x_{ic}^{{\bm{W}}^{\prime}}:=(\bm{e}_{y_{i}}-\bm{e}_{c})^{T}({\bm{W}}+\psi\bm{\Delta}{\bm{W}})\bm{h}_{i}$, then we have for $i\in[n]$, $c\neq y_{i}$
$\frac{f^{\prime}(x_{ic}^{{\bm{W}}^{\prime}})}{f^{\prime}(x_{ic}^{{\bm{W}}})}=|\frac{f^{\prime}(x_{ic}^{{\bm{W}}^{\prime}})}{f^{\prime}(x_{ic}^{{\bm{W}}})}|\leq e^{|x_{ic}^{{\bm{W}}}-x_{ic}^{{\bm{W}}^{\prime}}|}+2$ $=e^{\psi|(\bm{e}_{c}-\bm{e}_{y_{i}})^{T}\bm{\Delta}{\bm{W}}\bm{h}_{i}|}+2=e^{\psi|\langle\bm{\Delta}{\bm{W}},(\bm{e}_{c}-\bm{e}_{y_{i}})\bm{h}_{i}^{T}\rangle|}+2$
$\leq e^{\psi\lVert\bm{\Delta}{\bm{W}}\rVert_{\max}{\left\|(\bm{e}_{c}-\bm{e}_{y_{i}})\bm{h}_{i}^{T}\right\|_{\rm{sum}}}}+2$
$=e^{\psi\lVert\bm{\Delta}{\bm{W}}\rVert_{\max}{\left\|\bm{e}_{c}-\bm{e}_{y_{i}}\right\|_{\rm{sum}}}{\left\|\bm{h}_{i}\right\|_{\rm{sum}}}}+2$
$\leq e^{2B\psi{\left\|\bm{\Delta}{\bm{W}}\right\|_{\max}}}+2.$
This leads to $\sum_{i\in[n]}\sum_{c\neq y_{i}}f^{\prime}(x_{ic}^{{\bm{W}}^{\prime}})\leq(e^{2B\psi{\left\|\bm{\Delta}{\bm{W}}\right\|_{\max}}}+2)\sum_{i\in[n]}\sum_{c\neq y_{i}}f^{\prime}(x_{ic}^{{\bm{W}}})$. Rearrange and using the definition of ${\mathcal{G}}_{pll}({\bm{W}})$, we obtain the desired. ∎
Lemma F.4 (Analogue of Lemma B.4 for PLL).
Suppose that there exists ${\bm{W}}\in\mathbb{R}^{k\times d}$ such that $\mathcal{L}_{pll}({\bm{W}})\leq\frac{\log 2}{n}$, then we have
$(\bm{e}_{y_{i}}-\bm{e}_{c})^{T}{\bm{W}}\bm{h}_{i}\geq 0,\quad\text{for all $i\in[n]$ and for all $c\in[k]$ such that $c\neq y_{i}$}.$ (36)
Proof.
Denote $x_{ic}=(\bm{e}_{y_{i}}-\bm{e}_{c})^{T}{\bm{W}}\bm{h}_{i}$. Then, by the assumption, we have for any $i\in[n],c\neq y_{i}$
$\log(1+e^{-x_{ic}})\leq\sum_{i\in[n]}\sum_{c\neq y_{i}}\log(1+e^{-x_{ic}})\leq\log(2).$
This implies that $x_{ic}\geq 0$ for all $i\in[n],c\neq y_{i}$. ∎
Lemma F.5 (Analogue of Lemma B.2 for PLL).
For any ${\bm{W}},{\bm{W}}_{0}\in\mathbb{R}^{k\times d}$, suppose that $\mathcal{L}({\bm{W}})$ is convex, we have
$|\mathcal{L}_{pll}({\bm{W}})-\mathcal{L}_{pll}({\bm{W}}_{0})|\leq 2B{\left\|{\bm{W}}-{\bm{W}}_{0}\right\|_{\max}}.$
Proof.
This lemma is a direct consequence of Lemma F.1 and be proved in the same way as Lemma B.2. ∎
Thus, we have proved all the Lemmas for ${\mathcal{G}}_{pll}({\bm{W}})$ and its relationships to $\mathcal{L}_{pll}({\bm{W}})$ in analogous to those in section B. The proof of SignGD ((4)) or NSD ((5)) with PairLogLoss follow the same steps as with cross-entropy loss given in section C.